### v1

# Generalized Recurrence Criteria for Classes of Open Quantum Walks

Abstract
In this paper, we study the recurrence of open quantum walks (OQWs) induced by finite-dimensional coins $(L,B,R)$. The focus is on homogeneous OQWs with a set of vertices $\mathbb{Z}$, the set of integers. We present three distinct recurrence criteria, each adapted to different types of coins. The first criterion was developed for a class of Lazy OQWs in any finite dimension, where the presented criterion is associated with an auxiliary map and its only invariant state, resulting in the first recurrence criterion for Lazy OQWs. The second one is restricted to Lazy OQWs of dimension 2, where we provide a complete characterization of the recurrence for this lower dimension. Finally, we present a general criterion for finite-dimensional coins in the non-lazy case $(B=0)$, which generalizes many of the previously known results. This new criterion holds for irreducible and reducible OQWs through a decomposition of the Hilbert space where our quantum states act.
keywords:
Open Quantum Walks, Recurrence, Lazy walks, Coin, General Criterion
1 Introduction
Open Quantum Walks (OQWs) represent a natural extension of classical Markov chains through open quantum systems. Originally proposed by [2], OQWs are models that capture the behavior of open quantum systems, in which the measurement of the particle position after each time step leads to a classical Markov chain when it is joined with the space of particle states.
In this context, we consider homogeneous OQWs defined on the set of integers, that is, the graph of this type of random walk is the set of vertices $V=\mathbb{Z}$, where the linear operators $R$ and $L$ play the roles of transitions between the vertices of the quantum walk, which are constant and each jump only occurs from one vertex to its nearest neighbors. Homogeneous OQWs in the Lattice are explored and detailed in [6].
Recurrence is an important statistic of any random walk, emerging as a central question of interest in the quantum context, and can be directly connected with the classical recurrence criteria for random walks in $\mathbb{Z}$. For example, for homogeneous classical random walks on $\mathbb{Z}$, it is well known that a walk is recurrent if and only if the transition parameters $r$ (rate of jumping to right) and $l$ (rate of jumping to left) are equal (or $r=1/2=l$ in the case $b=0$) [11]. This classical random walk has its graph illustrated in Figure 1.
[…]
Figure 1: Classical Lazy Random Walk on $\mathbb{Z}$.
With this description, it is natural to seek criteria to determine under what conditions a homogeneous OQW is recurrent given $R$ and $L$. This problem has been addressed in the literature with the development of specific criteria, which are essential to understanding the dynamics and long-term behavior of these systems.
A Criterion for PQ Matrices can be found in [10]. Also, in that work the authors analyze the positive recurrence of this type of walk. From the analysis of the spectrum of $L^{*}L$ and $R^{*}R,$ it was showed in [7] that $R$ and $L$ having only singular values equal to $1/\sqrt{2}$ is a sufficient condition to guarantee that those OQWs are recurrent. More recently, it was obtained general criteria for recurrence of non-Lazy OQWs in dimension 2 and for irreducible finite-dimensional OQWs [9].
In this article, we present a general criterion for the recurrence of homogeneous OQWs, based on the transition operators $R$ and $L$ when acting in finite-dimensional vector spaces, that is, we allow the OQW to be reducible in any finite dimension. Furthermore, we present the first recurrence criterion for lazy OQWs, using results obtained from recent studies on Central Limit Theorems for this kind of OQWs.
A first Central Limit Theorem for OQWs was presented by [1]. Here we will use more recent results. These works presented a Central Limit Theorem for Lazy OQWs [8] and one for any homogeneous OQW in Lattice when the degrees of freedom are finite [5].
This work is structured as follows. In Section 2 we define the open quantum walks induced by a coin and their transition probabilities. Section 3 is devoted to presenting a recurrence criterion for a class of Lazy OQWs, and one restricted to Lazy OQWs of dimension 2. In Section 4 we focus on the case where the coin is non-lazy and we present a general recurrence criterion for finite-dimensional coins. Section 5 illustrates the results with comprehensive examples.
2 General Settings
Let us consider a set of vertices $V$, which might be finite or countable infinite. Let $\mathcal{K}=\mathbb{C}^{V}$ be the state space of a quantum system with as many degrees of freedom as the number of vertices. In the case where $V$ is infinite countable, we assume $\mathcal{K}$ to be any separable Hilbert space with an orthonormal basis indexed by $V$, denote by $\left\{\lvert i\rangle\right\}_{i\in V}$. To emphasize the position of the particle of our quantum walk, we consider a quantum system described by the separable complex Hilbert space
$\mathcal{H}=\bigoplus_{i\in V}\mathfrak{h}_{i},$
where the $\mathfrak{h}_{i}$ are separable Hilbert spaces. The label $i\in V$ represents the position of the particle and, when the particle is located at the vertex $i\in V$ , its internal state is encoded in the space $\mathfrak{h}_{i},$ as it will be detailed below.
We consider the set of density operators over $\mathcal{H},$
$\mathcal{D}(\mathcal{H})=\{\rho\in\mathcal{H}:\rho\geq 0,\mathrm{Tr}(\rho)=1\},$
and introduce
$\mathcal{S}(\mathcal{H}\otimes\mathcal{K})=\left\{\tau\in\mathcal{I}(\mathcal{H}\otimes\mathcal{K}):\textstyle\tau=\displaystyle\sum_{j\in\mathbb{Z}}\rho(j)\otimes\lvert j\rangle\langle j\rvert\right\},$
the set of density operators acting on $\mathcal{H}\otimes\mathcal{K}.$ Fixed a set of vertices $V,$ an open quantum random walk (OQW) on $V$ is a completely positive map
$\Phi:\mathcal{H}\otimes\mathcal{K}\mapsto\mathcal{H}\otimes\mathcal{K},$
given by
$\Phi(\tau)=\sum_{i\in V}\Bigg{(}\sum_{j\in V}B_{i}^{j}\rho_{j}B_{i}^{j^{*}}\Bigg{)}\otimes|i\rangle\langle i|,\;\;\;\tau\in\mathcal{S}(\mathcal{H}\otimes\mathcal{K}),$
where $B_{i}^{j}$ are bounded operators on $\mathcal{H}$ corresponding to the transition effect from site $j$ to $i$, and
$\sum_{i\in V}B_{i}^{j*}B_{i}^{j}=I,\;\;\;j\in V,$
which guarantees that $\mathrm{Tr}(\Phi(\tau))=\mathrm{Tr}(\tau)$, for every $\tau\in\mathcal{D}(\mathcal{H})$, meaning that $\Phi$ is trace-preserving, thus $\Phi$ preserves the set $\mathcal{S}(\mathcal{H}\otimes\mathcal{K}),$ i.e., $\Phi(\tau)\in\mathcal{S}(\mathcal{H}\otimes\mathcal{K})\;\forall\tau\in\mathcal{S}(\mathcal{H}\otimes\mathcal{K}).$
In this work, we focus on the OQW model presented in [2], where the set of vertices is $V=\mathbb{Z}.$ For simplicity, we will say integer line, or simply the line, to refer to the set $\mathbb{Z}$ of integer numbers. An OQW on $\mathbb{Z}$ concerns the evolution of density operators of the form
$\tau=\sum_{j\in\mathbb{Z}}\rho(j)\otimes\lvert j\rangle\langle j\rvert\;\in\;\mathcal{H}\otimes\mathbb{C}^{\mathbb{Z}},$
where the $V=\mathbb{Z}$ represents the position of the walker and $\mathfrak{h}_{i}$ is the Hilbert space describing the internal degrees of freedom for site $\lvert i\rangle$, which we call the coin space and will be denoted by $\mathfrak{h}_{i}=\mathbb{C}^{k}$. Finally, we consider homogeneous lazy open quantum walks: at each step either the particle on site $\lvert j\rangle$ moves exactly one position to some of its two nearest-neighbours, in the left ($\lvert j-1\rangle$) or in the right ($\lvert j+1\rangle$), or it remains at same site ($\lvert j\rangle$), and the coin is constant, i.e., the same transition rule is applied at each site. Therefore, the linear operator $\Phi$ can be written by
$\Phi(\tau)=\sum_{i\in\mathbb{Z}}\sum_{j=i-1}^{i+1}M_{i}^{j}\tau(M_{i}^{j})^{*},$
where
$M_{i}^{i-1}=L\otimes\lvert i-1\rangle\langle i\rvert,\quad M_{i}^{i}=B\otimes\lvert i\rangle\langle i\rvert,\quad M_{i}^{i+1}=R\otimes\lvert i+1\rangle\langle i\rvert.$
The linear operators $L,B,R\in\mathbb{M}_{k}({\mathbb{C}})$ are matrices such that $L^{*}L+B^{*}B+R^{*}R=I_{k},$ and simplified by
$\Phi(\tau)=\sum_{i\in\mathbb{Z}}\Big{(}L\rho(i+1)L^{*}+B\rho(i)B^{*}+R\rho(i-1)R^{*}\Big{)}\otimes|i\rangle\langle i|,\;\;\;\tau\in\mathcal{S}(\mathcal{H}\otimes\mathcal{K}).$ (1)
We say that the triple of matrices $(L,B,R)$ is a coin if eq. (1) holds. Therefore, we can assume that $\mathfrak{h}_{i}=\mathfrak{h}\;\forall i\in\mathbb{Z},$ and if $\mathfrak{h}=\mathbb{C}^{d}$ for some $d$, the dimension of the coin is said to be $d$. In this article, we consider only the case $d<\infty.$
If $B\neq 0,$ the quantum walk exhibits a lazy behavior, meaning that in the context of quantum walks there is a non-zero probability for the walker to stay in the same vertex after a step. We refer to the OQW above as an OQW induced by the coin $(L,B,R)$ in which the walk is inherently lazy. This lazy behavior introduces distinct properties compared to its non-lazy counterpart. Notably, the presence of the matrix $B$ influences the walk’s propensity to remain in its current vertex, thereby affecting the overall dynamics. If $B=0,$ then $\Phi$ will be called an OQW induced by a coin $(L,R)$ or an non-lazy OQW induced by a coin $(L,R).$
The dynamic described above is represented by the graph in Figure 2.
[…]
Figure 2: Lazy Quantum Walk on $\mathbb{Z}$.
The OQW induced by a coin $(L,B,R)$ is the homogeneous, nearest-neighbor OQW on the integer line with the left and right transitions given by $L$ and $R$, respectively, and $B$ represents the rate of no jump at some instant. In this sense, the quantum trajectory of this kind of OQW on the line, induced by a coin $(L,B,R)$ and starting from a state $\tau$ of the form $\tau=\sum_{i\in\mathbb{Z}}\rho(i)\otimes|i\rangle\langle i|,$ is any path generated by the Markov chain $(X_{n},\tau_{n})_{n\geq 0}$, where $X_{n}$ denotes the position of the particle at time $n$ and $\tau_{n}$ its internal degree. The transition probabilities are given by
$\begin{split}\mathbb{P}\left(\,(X_{n+1},\tau_{n+1})=\left(i-1,\frac{L\sigma L^{*}}{\mathrm{Tr}(L\sigma L^{*})}\right)\;\Bigg{|}\;(x_{n},\rho_{n})=(i,\sigma)\,\right)&=\mathrm{Tr}(L\sigma L^{*}),\\ \mathbb{P}\left(\,(X_{n+1},\tau_{n+1})=\left(i,\frac{B\sigma B^{*}}{\mathrm{Tr}(B\sigma B^{*})}\right)\;\Bigg{|}\;(x_{n},\rho_{n})=(i,\sigma)\,\right)&=\mathrm{Tr}(B\sigma B^{*}),\\ \mathbb{P}\left(\,(X_{n+1},\tau_{n+1})=\left(i+1,\frac{R\sigma R^{*}}{\mathrm{Tr}(R\sigma R^{*})}\right)\;\Bigg{|}\;(x_{n},\rho_{n})=(i,\sigma)\,\right)&=\mathrm{Tr}(R\sigma R^{*}).\end{split}$
for every $i\in\mathbb{Z}$, $\sigma\in\mathcal{S}(\mathfrak{h}_{i})$, and initial law
$\mathbb{P}\left(\,(X_{0},\rho_{0})=\left(i,\frac{\rho(i)}{\mathrm{Tr}\,\rho(i)}\right)\;\right)=\mathrm{Tr}(\rho(i)),$
and all other transition probabilities are null.
3 Recurrence criteria for Lazy Open Quantum Walks
3.1 Faithfulness of Density Operators
Density operators play a fundamental role in open quantum random walks, providing a statistical description of this type of quantum system. Faithful and non-faithful density operators are two categories that arise in the study of quantum states, being characterized by their positive definiteness and positive-semidefiniteness, respectively.
Faithful density operators are defined as being positive definite. This means that the eigenvalues of a faithful density operator are strictly positive and it is invertible. In essence, faithfulness ensures a faithful representation of the quantum system, capturing the full richness of its physical properties.
On the other hand, non-faithful density operators are positive-semidefinite, thus they have $\lambda=0$ as one of their eigenvalues. Thereby, that there may be ambiguity in the quantum state description, as multiple states can correspond to the same non-faithful density operator. The positivity-semidefiniteness criterion relaxes the strictness of positive definiteness, permitting a broader range of quantum state representations.
An important subset of the set of non-faithful density operators is composed of the pure density operators. A pure density is a density $\rho$ that is a projection of rank 1, i.e., there exists some vector $\lvert u\rangle\in\mathbb{C}^{k}$ such that $\rho=\lvert u\rangle\langle u\rvert$.
Let us consider an OQW starting the walk on vertex $\lvert i\rangle$ with initial density operator $\rho.$ For evolution on time $n=0,1,2,\ldots,$ we define
$p_{n}(j)=\mathbb{P}(\mbox{the quantum particle, at time }n,\mbox{ is in site}\;\lvert j\rangle)=\mathrm{Tr}(\rho_{n}(j)),$
where
$\Phi^{n}(\rho\otimes\lvert i\rangle\langle i\rvert)=\sum_{s\in V}\rho_{n}(s)\otimes\lvert s\rangle\langle s\rvert.$
With this notation,
$p_{ji;\rho}(n)=p_{n}(\rho\otimes|i\rangle\to|j\rangle)=\mathrm{\mathrm{Tr}}(\rho_{n}(j)\otimes\lvert j\rangle\langle j\rvert)=\mathrm{\mathrm{Tr}}\left(\Phi^{n}(\rho\otimes\lvert i\rangle\langle i\rvert)(I\otimes\lvert j\rangle\langle j\rvert)\right)$
denote the probability of being at site $j$ at step $n,$ given that we started at site $i,$ with initial density $\rho$ concentrated at $i.$ Therefore, the dynamics starts with a density operator $\rho$ concentrated at some vertex $\lvert i\rangle,$ takes the evolution up to time ($n$) through the linear operator $\Phi^{n},$ producing a new density operator
$\rho_{n}=\sum_{s}\rho_{n}(s)\otimes\lvert s\rangle\langle s\rvert=\Phi^{n}(\rho\otimes\lvert i\rangle\langle i\rvert),\quad\mathrm{Tr}\left(\sum_{s}\rho_{n}(s)\right)=1,$
and then we project $\rho_{n}$ onto the subspace generated by vertex $\lvert j\rangle.$ Note that
$\Phi^{n}(\rho\otimes\lvert i\rangle\langle i\rvert)(I\otimes\lvert j\rangle\langle j\rvert)=\rho_{n}(j)\otimes\lvert j\rangle\langle j\rvert,$
which represents the data concentrated at vertex $\lvert j\rangle$ at time $n.$
Fixed some $i\in V,$ we denote the set of density operators acting on $\mathfrak{h}_{i}$ by
$\mathcal{S}(\mathfrak{h}_{i})=\left\{\rho:\mathfrak{h}_{i}\rightarrow\mathfrak{h}_{i}:\;\mathrm{Tr}(\rho)=1,\;\rho=\rho^{*};\;\rho\geq 0\right\}.$
Consider an OQW on some set of vertices $V$, let $\lvert i\rangle\in V$ and $\rho\in\mathcal{S}(\mathfrak{h}_{i})$. We say that a vertex $\lvert i\rangle$ is
  1. 1.
    $\rho$-recurrent if
    $\sum_{n=0}^{\infty}p_{ii;\rho}(n)=\infty.$
    Otherwise, $i$ is said to be $\rho$-transient;
  2. 2.
    recurrent, if $i$ is $\rho-$recurrent for all $\rho\in\mathcal{S}(\mathfrak{h}_{i});$
  3. 3.
    transient, if $i$ is $\rho-$transient for all $\rho\in\rho\in\mathcal{S}(\mathfrak{h}_{i}).$
Concerning all vertices, an OQW is called: recurrent if every vertex is recurrent; transient if every vertex is transient.
The following Proposition describes $\rho$-recurrence in a concentrated vertex $\lvert i\rangle\in V$ when assumed that $\lvert i\rangle$ is recurrent for some density and $\mathfrak{h}_{i}$ has a finite internal degree of freedom. An immediate consequence will also be given for the case where $\lvert i\rangle$ is $\sigma$-transient for some non-faithful density.
Theorem 3.1.
Consider an OQW on some set of vertices $V,$ let $i\in V$ and assume that $\mathfrak{h}_{i}$ is finite-dimensional. If $\lvert i\rangle$ is $\tau$-recurrent for some $\tau\in\mathcal{S}(\mathfrak{h}_{i}),$ then
  1. 1.
    vertex $\lvert i\rangle$ is $\rho$-recurrent for all faithful $\rho\in\mathcal{S}(\mathfrak{h}_{i})$;
  2. 2.
    there exists a non-faithful density operator $\tilde{\rho}$ such that $\lvert i\rangle$ is $\tilde{\rho}$-recurrent;
  3. 3.
    there exists a pure density operator $\lvert w\rangle\langle w\rvert$ such that the OQW is $\lvert w\rangle\langle w\rvert$-recurrent, where $\lvert w\rangle$ is an unitary vector of $\mathfrak{h}_{i}.$
Proof.
1. Let $\rho\in\mathcal{S}(\mathfrak{h}_{i})$ be faithful. By the compactness of $\mathcal{S}(\mathfrak{h}_{i}),$ there exists $c>0$ such that $\rho>c\tau,$ thus
$\begin{split}\sum_{n=0}^{\infty}p_{ii;\rho}(n)=&\sum_{n=0}^{\infty}\textmd{Tr}\left[\Phi^{n}\left(\rho\otimes\lvert i\rangle\langle i\rvert\right)\right(I\otimes\lvert i\rangle\langle i\rvert)]\\ >&c\sum_{n=0}^{\infty}\textmd{Tr}\left[\Phi^{n}\left(\tau\otimes\lvert i\rangle\langle i\rvert\right)(I\otimes\lvert i\rangle\langle i\rvert)\right]\\ =&c\sum_{n=0}^{\infty}p_{ii;\tau}(n)\\ =&\infty.\end{split}$
2. Let $\rho\in\mathcal{S}(\mathfrak{h}_{i})$ be non-faithful. By the Spectral Theorem, $\rho$ can be written as
$\rho=\sum_{x=1}^{n}\lambda_{x}\lvert x\rangle\langle x\rvert,$ (2)
where the vectors $\lvert x\rangle$ are the eigenvectors of $\rho$ with corresponding eigenvalues $\lambda_{x}$. Since $\rho$ is non-faithful, there is at least one null eigenvalue and the remainder of eigenvalues are non-negative summing $1$ Thus, (2) can be rewritten as
$\rho=\sum_{x\in S}\lambda_{x}\lvert x\rangle\langle x\rvert,\quad S\varsubsetneq\{1,\ldots,n\}.$
Take a sequence of positive numbers $(\alpha_{r})_{r\in R},$ where $R:=\{1,\ldots,n\}/S\neq\emptyset,$ whose sum is $1$
Defining
$\rho_{X}=\sum_{x\in S}\frac{\lambda_{x}}{2}\lvert x\rangle\langle x\rvert+\sum_{x\in R}\frac{\alpha_{x}}{2}\lvert x\rangle\langle x\rvert=\sum_{x=1}^{n}\frac{\tilde{\alpha}_{x}}{2}\lvert x\rangle\langle x\rvert,\quad\tilde{\alpha}_{x}=\begin{cases}\lambda_{x},&\mbox{if }x\in S\\ \alpha_{x},&\mbox{if }x\in R\end{cases},$
we get by the first item that $\lvert i\rangle$ is $\rho_{X}$-recurrent, since $\rho_{X}$ is faithful.
Now, define
$\rho_{Y}=\sum_{x\in R}\alpha_{x}\lvert x\rangle\langle x\rvert,$
which is a non-faithful density operator. Since $2\rho_{X}=\rho+\rho_{Y},$ we obtain by linearity the equality
$\sum_{n=0}^{\infty}p_{ii;\rho}(n)+\sum_{n=0}^{\infty}p_{ii;\rho_{Y}}(n)=\sum_{n=0}^{\infty}\left(p_{ii;\rho}(n)+p_{ii;\rho_{Y}}(n)\right)=2\sum_{n=0}^{\infty}p_{ii;\rho_{X}}(n).$
The sum on the right hand diverges once $\lvert i\rangle$ is $\rho_{X}$-recurrent, implying that at least one of the sums on the left-hand side diverges. Therefore, $\lvert i\rangle$ is $\rho$-recurrent or $\rho_{Y}$-recurrent, both being non-faithful.
3. Let $\tilde{\rho}$ be one non-faithful density given by item (2), then we have
$\tilde{\rho}=\sum_{j}\lambda_{j}\lvert e_{j}\rangle\langle e_{j}\rvert,$
where each $\lambda_{j}\geq 0$ is an eigenvalue of $\tilde{\rho}$ with corresponding eigenvector $\lvert e_{j}\rangle.$ If $\tilde{\rho}$ has only one eigenvector, then we are done. If $\tilde{\rho}$ has at least two eigenvectors, then
$\infty=\sum_{n=0}^{\infty}\textmd{Tr}\left[\Phi^{n}\left(\rho\otimes\lvert i\rangle\langle i\rvert\right)\right(I\otimes\lvert i\rangle\langle i\rvert)]=\sum_{j,\lambda_{j}\neq 0}\left(\lambda_{j}\sum_{n=0}^{\infty}\textmd{Tr}\left[\Phi^{n}\left(\lvert e_{j}\rangle\langle e_{j}\rvert\otimes\lvert i\rangle\langle i\rvert\right)\right(I\otimes\lvert i\rangle\langle i\rvert)]\right),$
where at least one of the members of the sum must be infinity, and some $\lambda_{j}$ is positive. This completes the proof.
$\square$
By a contra-position argument, we obtain the following direct consequence of Theorem 3.1.
Corollary 3.2.
Let $\Phi$ be an OQW on some set of vertices $V$ and $i\in V,$ and $\dim(\mathfrak{h}_{i})<\infty.$ If $\lvert i\rangle$ is $\tau$-transient for some faithful $\tau\in\mathcal{S}(\mathfrak{h}_{i}),$ then $\lvert i\rangle$ is transient.
Another direct consequence of Theorem 3.1 is related to the homogeneous case.
Corollary 3.3.
Let $\Phi$ be a homogeneous OQW on $\mathbb{Z}$ with $\dim(\mathfrak{h})=d<\infty.$ If $\Phi$ is $\tau$-recurrent for some $\tau\in\mathcal{S}(\mathfrak{h}),$ then
  1. 1.
    $\Phi$ is $\rho$-recurrent for all faithful $\rho\in\mathcal{S}(\mathfrak{h})$;
  2. 2.
    there exists a non-faithful density operator $\tilde{\rho}$ such that $\Phi$ is $\tilde{\rho}$-recurrent;
  3. 3.
    there exists a pure density operator $\lvert w\rangle\langle w\rvert$ such that $\Phi$ is $\lvert w\rangle\langle w\rvert$-recurrent, where $\lvert w\rangle$ is an unitary vector of $\mathbb{C}^{d}.$
3.2 Enclosures and Invariant States
Firstly, a subspace $\mathcal{Y}$ of $\mathfrak{h}$ is called an enclosure for a quantum channel $\mathcal{L}$ if for every positive $\sigma\in L^{1}(\mathfrak{h}),$ we have
$\mathrm{supp}(\sigma)\subset\mathcal{Y}\Rightarrow\mathrm{supp}\left(\mathcal{L}(\sigma)\right)\subset\mathcal{Y}.$
If an enclosure contains no other non-trivial enclosures, then it is called a minimal enclosure.
A minimal enclosure with finite dimension is the support of a unique invariant state of the channel, that is, a density operator $\sigma$ such that $\mathcal{L}(\sigma)=\sigma$. For more on enclosures of OQWs, we refer the reader to [5, 6, 3].
Following this property and using the homogeneity of OQWs induced by a coin, we split $\mathcal{S}(\mathfrak{h})$ into two subsets $\mathcal{S}(\mathcal{Y})$ and $\mathcal{S}(\mathcal{X})$ in such a way that the walk is transient for $\mathcal{S}(\mathcal{Y})$ and recurrent for $\mathcal{S}(\mathcal{X}),$ where $\mathcal{Y}$ is a enclosure of $\mathcal{L}:$
Proposition 3.4.
Consider an OQW induced by a coin $(L,B,R)$ of dimension $d<\infty$. We are in one (and only one) of the following situations:
  • $(L,B,R)$ is recurrent;
  • $(L,B,R)$ is transient;
  • $(L,B,R)$ is recurrent with respect to all densities except for a convex set $\mathcal{S}(\mathcal{Y}).$ In this case, $\mathcal{Y}$ is an enclosure of $\mathcal{L},$
    $\mathfrak{h}=\mathcal{Y}\oplus\mathcal{X}$
    for some subspace $\mathcal{X}$ of $\mathfrak{h},$ with $\mathcal{Y},\mathcal{X}\neq\{0\}.$
Proof.
Let us suppose that $(L,B,R)$ is neither recurrent nor transient. By Proposition 3.1, there exist unitary vectors $\lvert u\rangle,\lvert v\rangle\in\mathfrak{h}$ such that the walk is $\lvert u\rangle\langle u\rvert$-recurrent and $\lvert v\rangle\langle v\rvert$-transient. Let $\mathcal{Y}$ be the subset of $\mathfrak{h}$ such that $\mathcal{S}(\mathcal{Y})$ is the set of densities of $\mathcal{B}(\mathfrak{h})$ in which $(L,B,R)$ is transient, thus we pick $\rho\in\mathcal{S}(\mathcal{Y})$ and we can affirm that $\rho_{n}\in\mathcal{S}(\mathcal{Y}),$ where $(X_{n},\rho_{n})_{n}$ are the quantum trajectories of the OQW. Indeed, if we had $\rho_{n}\in\mathcal{S}(\mathfrak{h})\setminus\mathcal{S}(\mathcal{Y})$ for some $n,$ then we could refresh the walk and the vertex $\lvert X_{n}\rangle$ would be $\rho_{n}$-recurrent, contradicting the hypothesis of homogeneity, thus we must have that $\mathcal{Y}$ is an enclosure of $\mathcal{L}.$ Therefore, take $\mathcal{X}=\mathcal{Y}^{\perp}$ and the trichotomy holds.
$\square$
In our subsequent discussion, we consider a coin $(L,B,R)$ and its auxiliary map $\mathcal{L}:\mathcal{B}(\mathfrak{h})\rightarrow\mathcal{B}(\mathfrak{h})$ given by
$\mathcal{L}(\rho)=L\rho L^{*}+B\rho B^{*}+R\rho R^{*},\quad\rho\in\mathcal{B}(\mathfrak{h}).$
Note that $\mathcal{L}$ is a quantum channel, thus it has at least one invariant state, that is, there exists some density operator $\rho_{\infty}\in\mathcal{B}(\mathfrak{h})$ satisfying $\mathcal{L}(\rho_{\infty})=\rho_{\infty}.$ When the invariant state of $\mathcal{L}$ is unique, we will say that the map $\mathcal{L}$ is ergodic.
We will use the following result obtained straightforwardly from [8].
Theorem 3.5.
Consider a Lazy OQW induced by a coin $(L,B,R)$ of dimension $d<\infty.$ If the auxiliary map $\mathcal{L}$ admits only one invariant state $\rho_{\infty},$ then
$\frac{X_{n}}{n}\stackrel{{\scriptstyle n\rightarrow\infty}}{{\longrightarrow}}m,\qquad\text{where }\qquad m=\mathrm{Tr}\left(L\rho_{\infty}L^{*}\right)-\mathrm{Tr}\left(R\rho_{\infty}R^{*}\right),$
in probability.
Proof.
This result can be obtained by using a Doob decomposition as described in [8, Equation 27] and [1, Theorem 5.2] for a Martingale $M_{n}$ satisfying
$X_{n}-n.m=X_{0}+M_{n}+\beta_{n},$
where $M_{n}/n\rightarrow 0$ and $\beta_{n}\rightarrow 0$ when $n\rightarrow 0.$
$\square$
Theorem 3.6 (Chung-Fuchs Theorem for OQWs with ergodic auxiliary map).
Consider an OQW induced by a coin $(L,B,R)$ such that $\mathcal{Y}$ is an enclosure Suppose that the initial density operator is $\rho\in\mathcal{S}(\mathcal{Y})$ with quantum trajectories $(X_{n},\rho_{n})_{n\geq 0}$ and $X_{n}/n\rightarrow 0$ in probability, then $(L,B,R)$ is $\sigma$-recurrent for some $\sigma\in\mathcal{S}(\mathcal{Y}).$
Proof.
The proof is analogous to [9, Theorem 7] and we will proceed by remarking on how it can be generalized. Firstly, the maximum appearing in [9, Lemma 6] can be taken in $\mathcal{D}=\mathcal{S}(\mathcal{Y}).$ Indeed, if the initial state is $\rho\in\mathcal{S}(\mathcal{Y}),$ then $\rho_{n}\in\mathcal{S}(\mathcal{Y})$ for every $n,$ thereby the maximum taken in the proof of [9, Lemma 6] can be reduced to $\mathcal{S}(\mathcal{Y}).$ Therefore, for the same reason, the density operator considered in [9, Theorem 7] belongs to $\mathcal{S}(\mathcal{Y}).$
$\square$
Theorem 3.7 (Recurrence Criteria).
Let us consider an OQW on $\mathbb{Z}$ induced by a coin $(L,B,R)$ of dimension $d$ such that the auxiliary map $\mathcal{L}$ is ergodic. Then $\mathcal{L}$ has a unique invariant state $\rho_{\infty}$ and
  • $\textmd{Tr}\left(L\rho_{\infty}L^{*}\right)=\textmd{Tr}\left(R\rho_{\infty}R^{*}\right)\Rightarrow$ the walk is recurrent;
  • $\textmd{Tr}\left(L\rho_{\infty}L^{*}\right)\neq\textmd{Tr}\left(R\rho_{\infty}R^{*}\right)\Rightarrow$ the walk is transient.
Proof.
We suppose that the walk is neither recurrent nor transient. In this case, the walk is recurrent for some density and transient for all densities in some $\mathcal{S}(\mathcal{Y}),$ where $\mathcal{Y}$ is an enclosure (Proposition 3.4). Moreover, the unique invariant state $\rho_{\infty}$ of $\mathcal{L}$ must belong to $\mathcal{S}(\mathcal{Y}).$
If $\;\textmd{Tr}\left(L\rho_{\infty}L^{*}\right)=\textmd{Tr}\left(R\rho_{\infty}R^{*}\right),$ then $X_{n}/n\rightarrow 0$ by Theorem 3.5, thus $(L,B,R)$ is $\rho$-recurrent for some $\rho\in\mathcal{S}(\mathcal{Y})$ by Theorem 3.6, which is a contradiction.
[…]
Therefore, if $\textmd{Tr}\left(L\rho_{\infty}L^{*}\right)=\textmd{Tr}\left(R\rho_{\infty}R^{*}\right),$ then the walk is $\rho$-recurrent for some density and thus it is recurrent. If $\textmd{Tr}\left(L\rho_{\infty}L^{*}\right)\neq\textmd{Tr}\left(R\rho_{\infty}R^{*}\right),$ then the walks is transient for every density.
$\square$
We remark that the operator $B$ does influence in the recurrence of $(L,B,R).$ Indeed, even it does not appear on the bullet points that describe the recurrence criterion, the invariant state $\rho_{\infty}$ may change when we pick distinct $B$’s. See example 5.3 below.
3.3 General Recurrence Criteria for OQWs induced by a coin $(L,B,R)$ of dimension 2
The following Proposition is a generalisation of [9, Proposition 12, item (3)]. We can prove it in an analogous demonstration given by [9], however, we would like to do it in terms of the non-ergodicity of the auxiliary map induced by this coin.
Proposition 3.8.
Consider an OQW induced by a coin $(L,B,R)$ of dimension 2 such that $L,B,R$ share two unitary linearly independent eigenvectors. Then for any $\lvert u\rangle$ orthogonal to one of these two eigenvectors, we have that $\lvert u\rangle$ is also a common eigenvector of $L,B,R.$
Proof.
Suppose $L,B,R$ share two unitary and linearly independent unitary eigenvectors $\lvert v\rangle,\lvert w\rangle,$ then $\tau_{1}=\lvert v\rangle\langle v\rvert$ and $\tau_{2}=\lvert w\rangle\langle w\rvert$ are distinct invariant states of the auxiliary map $\mathcal{L},$ thus $\mathcal{L}$ is not ergodic. This means that there are two minimal enclosures $\mathcal{Y}_{1},\mathcal{Y}_{2},$ and thus $\mathcal{Y}_{1}\perp\mathcal{Y}_{2}.$ The dimension of the coin gives that $\mathcal{Y}_{1}$ and $\mathcal{Y}_{2}$ are one-dimensional and therefore there exist $\lvert u_{1}\rangle,\lvert u_{2}\rangle\in\mathbb{C}^{2}$ such that $\mathcal{Y}_{1}=\mbox{span}(\lvert u_{1}\rangle)$ and $\mathcal{Y}_{2}=\mbox{span}(\lvert u_{2}\rangle).$ Thereby $\lvert u_{1}\rangle$ and $\lvert u_{2}\rangle$ are common orthonormal eigenvectors of $L,B,R.$
$\square$
Theorem 3.9.
Consider an OQW on $\mathbb{Z}$ induced by a coin $(L,B,R)$ of dimension 2.
(1) If $L,B,R$ have at most one common eigenvector, let $\rho_{\infty}$ be the unique invariant density of the auxiliary map. Then, we have
$(1.1)\quad\textmd{Tr}(L\rho_{\infty}L^{*})\neq\textmd{Tr}(R\rho_{\infty}R^{*})\Rightarrow(L,B,R)\mbox{ is transient},$
$(1.2)\quad\textmd{Tr}(L\rho_{\infty}L^{*})=\textmd{Tr}(R\rho_{\infty}R^{*})\Rightarrow(L,B,R)\mbox{ is recurrent}.$
(2) If $L,B,R$ have two linearly independent eigenvectors in common, $\lvert u_{1}\rangle,\lvert u_{2}\rangle$ unitaries, then they satisfy $\lvert u_{1}\rangle\bot\lvert u_{2}\rangle,$ we can put
$L\lvert u_{j}\rangle=l_{j}\lvert u_{j}\rangle,\qquad B\lvert u_{j}\rangle=b_{j}\lvert u_{j}\rangle,\qquad R\lvert u_{j}\rangle=r_{j}\lvert u_{j}\rangle,\qquad j=1,2,$
and
$(2.1)\quad|l_{1}|=|r_{1}|\mbox{ and }|l_{2}|=|r_{2}|$$\Rightarrow$ $(L,B,R)\mbox{ is recurrent,}$,
$(2.2)\quad|l_{1}|\neq|r_{1}|\mbox{ and }|l_{2}|\neq|r_{2}|$$\Rightarrow$ $(L,B,R)\mbox{ is transient,}$,
$(2.3)\quad|l_{i}|=|r_{i}|\mbox{ and }|l_{j}|\neq|r_{j}|$$\Rightarrow$ $(L,B,R)\mbox{ is transient with respect to }\sigma_{j}=\lvert u_{j}\rangle\langle u_{j}\rvert\mbox{ and it is}$
recurrent with respect to all densities but $\mbox{recurrent with respect to all densities but }\sigma_{j}.$
Proof.
If the elements of the coin do not share at least two orthogonal eigenvectors, then the auxiliary map is ergodic, thus item (1) is a particular case of Theorem 3.7.
Now we suppose that $L,B,R$ share two linearly independent eigenvectors, then we can apply Proposition 3.8 to obtain $\lvert u_{1}\rangle,\lvert u_{2}\rangle$ common orthonormal eigenvectors of $L,B,R.$ We assume that $|l_{1}|=|r_{1}|$ and $|l_{2}|=|r_{2}|,$ then for $\rho=\lvert u\rangle\langle u\rvert,$ where $\lvert u\rangle$ is some common eigenvector of $L,B,R,$ we have $p_{j,j+1;\rho}(n)=p_{j,j-1;\rho}(n)=w$ for all $j\in\mathbb{Z}$ and all $n\geq 0$ for some constant $w$ and therefore, in this case, we have a classical homogeneous walk which is recurrent. Since it is valid for any common eigenvector, the walk is recurrent by Proposition 3.1.
Analogously, if $|l_{1}|\neq|r_{1}|$ and $|l_{2}|\neq|r_{2}|,$ then for $\rho_{1}=\lvert u_{1}\rangle\langle u_{1}\rvert,\;\rho_{2}=\lvert u_{2}\rangle\langle u_{2}\rvert$ where $\lvert u_{1}\rangle,\lvert u_{2}\rangle$ are orthonormal common eigenvectors of $L,B,R,$ we have $p_{j,j+1;\rho}(n)<p_{j,j-1;\rho}(n)$ (or $>$) for all $j\in\mathbb{Z}$ and all $n\geq 0$ and therefore in this case we have a classical homogeneous walk which is transient for both $\rho_{1},\rho_{2}$. Since the walk is transient for two distinct eigenvectors, the walk is transient by Proposition 3.1.
For the last case, $|l_{i}|=|r_{i}|$ and $|l_{j}|\neq|r_{j}|,$ $(L,B,R)$ is $\rho_{i}$-recurrent and $\rho_{j}$-transient, where $\rho_{i}=\lvert u_{i}\rangle\langle u_{i}\rvert$ and $\rho_{j}=\lvert u_{j}\rangle\langle u_{j}\rvert.$ This situation is a particular case of the last item in Proposition 3.4.
$\square$
4 A General Recurrence Criteria for Open Quantum Walks
Consider a coin $(L,R)$ of dimension $n<\infty.$ By [5], we can obtain a decomposition of $\mathfrak{h}$ in the following way:
$\mathfrak{h}=\mathfrak{R}\oplus\mathfrak{T},$
where
$\mathfrak{R}=\sup\{\mbox{supp}(\omega)\;|\;\omega\mbox{ is an invariant state for }\mathcal{L}\}.$
Since $\mathfrak{h}$ is finite-dimensional, any minimal enclosure is included in $\mathfrak{R},$ and is the support of a unique extremal invariant state. Moreover, we have a unique decomposition of $\mathfrak{R}$ of the form
$\mathfrak{R}=\bigoplus_{\alpha\in A}\mathfrak{X}_{\alpha},$
where $\{\mathfrak{X}_{\alpha}\}_{\alpha}\in A$ is a finite set of mutually orthogonal enclosures and every $\mathfrak{X}_{\alpha}$ is minimal in the set of enclosures satisfying
[…]
Further, we can write
$\mathfrak{R}=\bigoplus_{\alpha\in A}\mathfrak{X}_{\alpha}=\bigoplus_{\alpha\in A}\bigoplus_{\beta\in I_{\alpha}}\mathfrak{V}_{\alpha,\beta},$
where every $\mathfrak{X}_{\alpha}$ either is a minimal enclosure or assumes a decomposition (not unique) as the sum of mutually orthogonal isomorphic minimal enclosures $\{\mathfrak{V}_{\alpha,\beta},\;\beta\in I_{\alpha}\}$. Given a minimal enclosure $\mathfrak{V}_{\alpha},$ let $\tau_{\alpha}$ be the unique invariant state with support on $\mathfrak{V}_{\alpha}.$ Introducing the parameter
$m_{\alpha}=\mathrm{Tr}\left(L\tau_{\alpha}L^{*}\right)-\mathrm{Tr}\left(R\tau_{\alpha}R^{*}\right),$
we have by [5, Lema 4.2] that $m_{\alpha}$ is independent of the particular minimal enclosure $\mathfrak{V}_{\alpha}$ in $\mathfrak{X}_{\alpha}.$ Therefore, we can combine Proposition 3.4, Theorem 3.7 and the discussion above to state the property
Proposition 4.1.
Consider a coin $(L,R).$ We have a decomposition
$\mathfrak{h}=\left(\bigoplus_{\alpha\in A}\mathcal{Y}_{\alpha}\right)\oplus\mathcal{X},$ (3)
where $A$ is a finite set, $\mathcal{Y}_{\alpha}$ is a subspace of $\mathfrak{h}$ being the support of exactly one invariant state $\tau_{\alpha}$ for $\mathcal{L},$ $\mathcal{X}$ contains no enclosure, and
  • $\textmd{Tr}\left(L\tau_{\alpha}L^{*}\right)=\dfrac{1}{2}\Rightarrow$ the walk is $\tau_{\alpha}$-recurrent for all $\tau\in\mathcal{S}(\mathcal{Y}_{\alpha})$;
  • $\textmd{Tr}\left(L\tau_{\alpha}L^{*}\right)\neq\dfrac{1}{2}\Rightarrow$ the walk is $\tau_{\alpha}$-transient for all $\tau\in\mathcal{S}(\mathcal{Y}_{\alpha}).$
Once the decomposition on Equation (3) is obtained, we can say that $\mathcal{Y}_{\alpha}$ is either recurrent or transient.
The following result is due to [5] and will be crucial to our criteria. To state it, consider a coin $(L,R),$ and let $\mathcal{Y}$ be an enclosure for $\mathcal{L}.$ Denoting $p_{\mathcal{Y}}$ the orthogonal projection onto $\mathcal{Y},$ we can define the associated absorption operator (see [4]):
$A(\mathcal{Y}):=\lim_{n\rightarrow\infty}\mathcal{L}^{*n}(p_{\mathcal{Y}}).$
Corollary 4.2.
Consider a coin $(L,R)$ with the decomposition of $\mathfrak{h}$ given on Equation (3), and let $a_{\alpha}(\rho)=\mathbb{E}_{\rho}\left[\mathrm{Tr}\left(A(\mathcal{Y}_{\alpha})\rho_{0}\right)\right].$ Then
$\frac{X_{n}-X_{0}}{n}\stackrel{{\scriptstyle n\rightarrow\infty}}{{\longrightarrow}}\sum_{\alpha\in A}a_{\alpha}(\rho)\delta_{m_{\alpha}},$ (4)
where $\delta_{m_{\alpha}}$ is the Dirac measure concentrated in $m_{\alpha}.$
Let us start an OQW induced by a finite-dimensional coin $(L,R)$ and we assume that, in the $n$-th step, the density operator is $\rho_{n}.$ Note that each $\mathcal{Y}_{\alpha}$ on the Proposition is an enclosure, and any $\rho_{n}\in\mathcal{S}(\mathfrak{h})$ can be written in the form
$\rho_{n}=\sum_{\alpha\in A}a_{\alpha}^{(n)}\tau_{\alpha}^{(n)}+\tau^{(n)},$
where $\tau_{\alpha}^{(n)}$ is a positive operator with support on $\mathcal{Y}_{\alpha}$ and $\tau^{(n)}$ is either null or a positive operator with support on $\mathcal{X}$ as in Equation (3). There are three possibilities concerning recurrence when $\tau^{(0)}$ is non-null and its support is not disjoint from $\mathcal{X}.$
1. $(L,R)$ is recurrent for $\mathcal{Y}_{\alpha}$ for some $\alpha$ and supp$\left(\mathcal{L}(\rho)\right)\cap\mathcal{Y}_{\alpha}\neq\emptyset$: in this case, the walk is $\rho$-recurrent. Indeed, concerning the quantum trajectories $(\rho_{n})_{n}$, we could refresh the walk on some instant $n$ at some vertex $\lvert X_{n}\rangle,$ which is $\rho_{n}$-recurrent, since supp$\left(\rho_{n}\right)\cap\mathcal{Y}_{\alpha}\neq\emptyset$ for a sufficient large $n$. If there was some density $\sigma$ with supp$(\sigma)\cap\mathcal{X}$ and $\sigma$-transient, there would be some enclosure contained on $\mathcal{X},$ which is a contradiction by Proposition 3.4. The conclusion is held by the homogeneity of the coin.
2. $(L,R)$ is transient for all $\mathcal{Y}_{\alpha}$ satisfying supp$\left(\mathcal{L}(\rho)\right)\cap\mathcal{Y}_{\alpha}\neq\emptyset$: the walk is $\sigma$-transient for all $\sigma\in\mathcal{S}(\mathcal{X})$. Indeed, in this case, starting the walk on vertex $X_{0}=0,$ $X_{n}/n$ converges (when $n\rightarrow\infty$) to
$m=\sum_{\alpha\in A}a_{\alpha}(\rho)\delta_{m_{\alpha}}=\sum_{\alpha\in A}a_{\alpha}(\rho),$
as given by Equation (4), where the second equality follows from $m_{\alpha}\neq 0,$ since $(L,R)$ is $\mathcal{Y}_{\alpha}$ transient. Thus, $X_{n}$ converges either to $\infty$- ∞ or $\infty$,∞ , in probability, imposing that the walk is $\rho$-transient and thus $(L,R)$ is transient on $\mathcal{S}(\mathcal{X}).$
Therefore, we have shown the following recurrence criterion:
Theorem 4.3 (A Generalized Criteria).
Consider a coin $(L,R)$ of dimension $d<\infty.$ Consider the decomposition of $\mathfrak{h}$ as given in Equation (3), and $\tau_{1},\ldots,\tau_{m}$ the respectively invariant states of $\mathcal{L}$ in $\mathcal{Y}_{\alpha},\alpha=1,\ldots,m.$
$(1)\quad\mathrm{Tr}(L\tau_{j}L^{*})\neq\dfrac{1}{2}\;\forall j\in\{1,\ldots m\}\Rightarrow(L,R)\mbox{ is transient;}$;
$(2)\quad\mathrm{Tr}(L\tau_{j_{k}}L^{*})=\dfrac{1}{2}\;$ for a set of indices $j_{k}\in J\subset\{1,\ldots,m\}$ and $\mathrm{Tr}(L\tau_{j_{l}}L^{*})\neq\dfrac{1}{2}\;$ for the remaining indices $j_{l}\in J^{C},$ and $\sigma$ is a density such that supp$(\sigma)\cap\mathcal{X}\neq\emptyset.$
(2.1) supp$\left(\mathcal{L}(\sigma)\right)\cap\mathcal{Y}_{\alpha}\neq\emptyset$ for some $\alpha\in J,$ then (L,R) is
$\begin{cases}\rho\text{-transient for }\rho\in\mathcal{D}_{1}:=\mathcal{S}(\oplus_{k\in J^{C}}\mathcal{Y}_{k})\\ \rho\text{-recurrent for }\rho\in\mathcal{D}_{1}^{C}.\end{cases}$
(2.2) supp$\left(\mathcal{L}(\rho)\right)\cap\mathcal{Y}_{\alpha}=\emptyset$ for all $\alpha\in J,$ then (L,R) is
$\begin{cases}\rho\text{-transient for }\rho\in\mathcal{D}_{2}:=\mathcal{S}(\oplus_{k\in J^{C}}\mathcal{Y}_{k}\oplus\mathcal{X})\\ \rho\text{-recurrent for }\mathcal{D}_{2}^{C}\end{cases}$
$(3)\quad\mathrm{Tr}(L\tau_{\alpha}L^{*})=\dfrac{1}{2}\;$ for every $\alpha\Rightarrow$ the walk is recurrent.
5 Examples
Example 5.1.
Our first examples will consider lazy OQWs of dimension 2.
  1. 1.
    Let $m,n>0$ and assume that $b:=1-m^{2}-n^{2}\in(0,1).$ Then the coin $(L,B,R)$ induced by
    $L=\begin{bmatrix}0&0\\ 0&-n\end{bmatrix},\qquad B=\begin{bmatrix}-\sqrt{b}&m\\ m&\sqrt{b}\end{bmatrix},\qquad R=\begin{bmatrix}n&0\\ 0&0\end{bmatrix}$
    is well-defined.
    The only invariant state for the auxiliary map is $\rho_{\infty}=I/2.$ Therefore,
    $\mathrm{Tr}(L\rho_{\infty}L^{*})=\frac{n^{2}}{2}=\mathrm{Tr}(R\rho_{\infty}R^{*})\Rightarrow\;(L,B,R)\mbox{ is recurrent,}$,
    by Item (1) of Theorem 3.9.
    It is worth noting that if we assume $m=0,$ then the walk is transient by Item (2) of Theorem 3.9. More explicitly, in this case, for the item (2.2), $|l_{1}|=0,\;|l_{2}|=n,\;|r_{1}|=n,\;|r_{2}|=0.$
    $\Diamond$
  2. 2.
    Let $x_{1},x_{2}\in\mathbb{C},$
    $|x_{1}|\leq\sqrt{\frac{2}{3}},\quad|x_{2}|\leq\sqrt{\frac{3}{4}},$
    then consider the coin $(L,B,R),$ where
    $L=\begin{bmatrix}\frac{1}{\sqrt{3}}&0\\ 0&\frac{1}{2}\end{bmatrix},\qquad B=\begin{bmatrix}\sqrt{\frac{2}{3}-x_{1}^{2}}&0\\ 0&\sqrt{\frac{3}{4}-x_{2}^{2}}\end{bmatrix},\qquad R=\begin{bmatrix}x_{1}&0\\ 0&x_{2}\end{bmatrix}.$
    We are in Item (2) of Theorem 3.9, thus we conclude that
    $(1)\quad|x_{1}|=\frac{1}{\sqrt{3}}\mbox{ and }|x_{2}|=\frac{1}{2}$$\Rightarrow$ $(L,B,R)\mbox{ is recurrent,}$,
    $(2)\quad|x_{1}|\neq\frac{1}{\sqrt{3}}\mbox{ and }|x_{2}|\neq\frac{1}{2}$$\Rightarrow$ $(L,B,R)\mbox{ is transient,}$,
    $(3)\quad|x_{1}|=\frac{1}{\sqrt{3}}\mbox{ and }|x_{2}|\neq\frac{1}{2}$$\Rightarrow$ $(L,B,R)\mbox{ is transient with respect to }\sigma=\begin{bmatrix}0&0\\ 0&1\end{bmatrix}\mbox{ and it is}$
    recurrent with respect to all densities but $\mbox{recurrent with respect to all densities but }\sigma,$
    $(4)\quad|x_{1}|\neq\frac{1}{\sqrt{3}}\mbox{ and }|x_{2}|=\frac{1}{2}$$\Rightarrow$ $(L,B,R)\mbox{ is transient with respect to }\sigma=\begin{bmatrix}1&0\\ 0&0\end{bmatrix}\mbox{ and it is}$
    recurrent with respect to all densities but $\mbox{recurrent with respect to all densities but }\sigma.$
$\Diamond$
Example 5.2.
Consider the Lazy OQW induced by the following coin of dimension 3: $(L,B,R),$ where
$L=\frac{1}{8}\begin{bmatrix}2(1+\sqrt{2})&0&2(1-\sqrt{2})\\ 0&\sqrt{31}&0\\ 2(1-\sqrt{2})&0&2(1+\sqrt{2})\end{bmatrix},\quad B=\frac{1}{16}\begin{bmatrix}\sqrt{30}&2i&\sqrt{30}\\ 2&0&2\\ \sqrt{30}&-2i&\sqrt{30}\end{bmatrix},\quad R=\frac{1}{8}\begin{bmatrix}2(1-\sqrt{2})&0&2(1+\sqrt{2})\\ 0&\sqrt{31}&0\\ 2(1+\sqrt{2})&0&2(1-\sqrt{2})\end{bmatrix}.$
There is a unique invariant state $\rho_{\infty}$ given by
$\rho_{\infty}=\frac{1}{2}\begin{bmatrix}1&0&-1\\ 0&0&0\\ -1&0&1\end{bmatrix},$
and $\mathrm{Tr}\left(L\rho_{\infty}L^{*}\right)=1/2=\mathrm{Tr}\left(R\rho_{\infty}R^{*}\right).$ Therefore, this coin is recurrent by Theorem 3.7.
$\Diamond$
Since the invariant state also depends on the choice of $B,$ it is natural to ask if for a given transient coin $(L,B,R),$ there is another $B^{\prime}$ that makes $(L,B^{\prime},R)$ to be recurrent.
Example 5.3.
Firstly, let
$L=\frac{1}{7}\begin{bmatrix}1&0\\ 0&2\end{bmatrix},\qquad B=\frac{1}{7}\begin{bmatrix}-1&1\\ 2&0\end{bmatrix},\qquad R=\frac{1}{7}\begin{bmatrix}1&1\\ 0&1\end{bmatrix}.$
Then
$\rho_{\infty}=\frac{1}{3}\begin{bmatrix}1&0\\ 0&2\end{bmatrix},$
$\mathrm{Tr}\left(L\rho_{\infty}L^{*}\right)=3/7$ and $\mathrm{Tr}\left(R\rho_{\infty}R^{*}\right)=5/21.$ The coin $(L,B,R)$ is transient. However, if we take
$B^{\prime}=\frac{1}{\sqrt{35}}\begin{bmatrix}5&-1\\ 0&2\end{bmatrix},$
then the invariant state is
$\rho^{\prime}_{\infty}=\begin{bmatrix}1&0\\ 0&0\end{bmatrix},$
and $\mathrm{Tr}\left(L\rho_{\infty}L^{*}\right)=1/\sqrt{7}=\mathrm{Tr}\left(R\rho_{\infty}R^{*}\right).$ Thus the coin $(L,B^{\prime},R)$ is recurrent.
$\Diamond$
We remark that the choices of the coins above were taken in such a way that $L,B^{\prime}$ and $R$ had a common invariant eigenvector, namely $[1,0]^{T},$ which was not an eigenvector of $B.$ This is due to the hard job of, given non-trivial $R$ and $L,$ to find $B^{\prime}$ in such a way that $(L,B^{\prime},R)$ is recurrent. Attempt were conduct using mathematic software with many $L$’s and $R$’s having simple entries, but even though the entries of $B$’s that made $(L,B,R)$ to be transient were very long to write in this paper.
Example 5.4.
Let us consider the coin given in [5, Section 6.2]:
$R=\begin{bmatrix}\sqrt{\frac{3}{8}}&0&0&0\\ -\sqrt{\frac{p_{1}}{2}}&\frac{1}{\sqrt{2}}&0&0\\ -\sqrt{\frac{p_{2}}{2}}&0&\frac{1}{\sqrt{2}}&0\\ \sqrt{\frac{2p_{3}}{3}}&0&0&\frac{1}{\sqrt{3}}\end{bmatrix},\qquad L=\begin{bmatrix}\frac{1}{2\sqrt{2}}&0&0&0\\ \sqrt{\frac{p_{1}}{2}}&\frac{1}{\sqrt{2}}&0&0\\ \sqrt{\frac{p_{2}}{2}}&0&\frac{1}{\sqrt{2}}&0\\ -\sqrt{\frac{p_{3}}{3}}&0&0&\frac{\sqrt{2}}{\sqrt{3}}\end{bmatrix},$
where $p_{1},\;p_{2},p_{3}\geq 0$ and $p_{1}/2+p_{2}/2+p_{3}/2=1/2.$
The invariant states of the auxiliary map are of the form $\rho_{\infty}=\sigma_{1}+\sigma_{2},$ where
$\sigma_{1}=\omega_{1}\begin{bmatrix}0&0&0&0\\ 0&a&b&0\\ 0&b^{*}&1-a&0\\ 0&0&0&0\end{bmatrix},\quad\sigma_{2}=\omega_{2}\begin{bmatrix}0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&0&1\end{bmatrix},\quad\omega_{i},\omega_{j}\geq 0,\;\;\omega_{1}+\omega_{2}=1.$
We have
$m_{1}=\mathrm{Tr}(L\sigma_{1}L^{*})=\frac{1}{2},\qquad m_{2}=\mathrm{Tr}(L\sigma_{2}L^{*})=\frac{2}{3}\neq\frac{1}{2}.$
By item (2) of Theorem 4.3, the walk is transient for
$\rho=\begin{bmatrix}0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&0&1\end{bmatrix}$
and recurrent for all the remaining density operators on $\mathbb{C}^{4}$, for any appropriate choices of $p_{k},\;k=1,2,3.$
Example 5.5.
Consider an OQW induced by the coin $(L,R),$ where
$L=\begin{bmatrix}\frac{\sqrt{5}}{5}&0&-\frac{\sqrt{5}}{5}\\ 0&\frac{2\sqrt{5}}{5}&\frac{\sqrt{5}}{10}\\ 0&0&\frac{1}{2}\end{bmatrix},\qquad R=\begin{bmatrix}\frac{2\sqrt{5}}{5}&0&\frac{\sqrt{5}}{10}\\ 0&\frac{\sqrt{5}}{5}&-\frac{\sqrt{5}}{5}\\ 0&0&\frac{1}{2}\end{bmatrix}.$
The invariant states of the auxiliary map are of the form $\rho_{\infty}=\sigma_{1}+\sigma_{2},$ where
$\sigma_{1}=\omega_{1}\begin{bmatrix}1&0&0\\ 0&0&0\\ 0&0&0\end{bmatrix},\quad\sigma_{2}=\omega_{2}\begin{bmatrix}0&0&0\\ 0&1&0\\ 0&0&0\end{bmatrix},\quad\omega_{i},\omega_{j}\geq 0,\;\;\omega_{1}+\omega_{2}=1.$
We have
$m_{1}=\mathrm{Tr}(L\sigma_{1}L^{*})=\frac{\sqrt{5}}{5}\neq\frac{1}{2},\qquad m_{2}=\mathrm{Tr}(L\sigma_{2}L^{*})=\frac{2\sqrt{5}}{5}\neq\frac{1}{2}.$
By item (1) of Theorem 4.3, the walk is transient.
Example 5.6.
Let us consider the coin $(L,R)$ of dimension 4 given by
$R=\begin{bmatrix}\frac{\sqrt{2}}{2}&-\frac{\sqrt{5}}{4}&0&\frac{1}{4}\\ 0&\frac{\sqrt{2}}{4}&0&0\\ 0&0&\frac{\sqrt{2}}{2}&0\\ 0&0&0&\frac{\sqrt{6}}{4}\end{bmatrix},\qquad L=\begin{bmatrix}\frac{\sqrt{2}}{2}&\frac{\sqrt{5}}{4}&0&-\frac{1}{4}\\ 0&\frac{1}{2}&0&\frac{\sqrt{5}}{4}\\ 0&0&-\frac{\sqrt{2}}{2}&0\\ 0&0&0&\frac{\sqrt{3}}{4}\end{bmatrix},$
The invariant states of the auxiliary map are of the form $\rho_{\infty}=\sigma_{1}+\sigma_{2},$ where
$\sigma_{1}=\omega_{1}\begin{bmatrix}1&0&0&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\end{bmatrix},\quad\sigma_{2}=\omega_{2}\begin{bmatrix}0&0&0&0\\ 0&0&0&0\\ 0&0&1&0\\ 0&0&0&0\end{bmatrix},\quad\omega_{i},\omega_{j}\geq 0,\;\;\omega_{1}+\omega_{2}=1.$
We have
$m_{1}=\mathrm{Tr}(L\sigma_{1}L^{*})=\frac{1}{2},\qquad m_{2}=\mathrm{Tr}(L\sigma_{2}L^{*})=\frac{1}{2}.$
By item (3) of Theorem 4.3, the walk is recurrent.
6 Conclusion
We present new general recurrence criteria for classes of homogeneous OQWs, whose dynamics are governed by Kraus operators that are induced by a finite-dimensional coin.
In the context of these OQWs, we presented three distinct criteria for analyzing recurrence. Firstly, Theorem 3.7 introduces a criterion for ergodic Lazy coins, where the recurrence is determined by the operators $L$ and $R$ and by the only invariant state $\rho_{\infty}$ associated with the corresponding auxiliary map.
Next, Theorem 3.9 focuses on Lazy OQWs of dimension 2, offering a complete and easy-to-apply criterion. The simplicity of this criterion comes from the fact that it depends on the knowledge of the eigenvectors of the operators $L$, $B$, and $R$, allowing a practical analysis of the recurrence of these processes that are not necessarily ergodic.
Finally, Theorem 4.3 extends the analysis to non-lazy coins, using a decomposition of the vector space where the operators $L$, $B$, and $R$ act. By checking recurrence in vector subspaces that contain unique invariant states, this general criterion is extended to any initial density, significantly expanding the scope of recurrence analysis for OQWs in more complex configurations, since it is valid for any finite dimensional coin.
[…]
Acknowledgements This work was supported by Universidade Federal do Pampa (UNIPAMPA).
[…]
Conflict of interest The author declare that there is no conflict of interest.
[…]
Data Availability The datasets generated during the current study are available from the corresponding author on reasonable request.