### v2

# Recurrence Criteria for Reducible Homogeneous Open Quantum Walks on the Line

Abstract
In this paper, we study the recurrence of homogeneous Open Quantum Walks (HOQWs) induced by finite-dimensional coins $(L,B,R)$. We present three distinct recurrence criteria, each adapted to different types of coins. The first criterion applies to coins whose auxiliary map has a unique invariant state, resulting in the first recurrence criterion for Lazy HOQWs. The second one applies to Lazy HOQWs of dimension 2, where we provide a complete characterization of the recurrence for this low-dimensional case. Finally, we present a general criterion for finite-dimensional coins in the non-lazy case $(B=0)$, which generalizes many of the previously known results. This new criterion holds for irreducible and reducible HOQWs.
keywords:
Homogeneous Open Quantum Walks, Recurrence, Lazy Walks, Coin, Reducibility
1 Introduction
Open Quantum Walks (OQWs) naturally extend classical Markov chains within the framework of open quantum systems. Originally proposed by [2], OQWs describe open quantum systems in which the repeated measurement of the particle’s position, together with the space of particle states, gives rise to a Markov chain.
In this context, we consider homogeneous OQWs (HOQWs) defined on the set of integers, that is, the graph of this type of random walk is the set of vertices $V=\mathbb{Z}$, where 3 linear operators $L,B,R$ play the roles of transitions between the vertices of the quantum walk, which are constant and each jump only occurs from one vertex to its nearest neighbors, or to itself. HOQWs on more general lattices are explored and detailed in [7].
Recurrence is an important probabilistic property of any random walk, often characterized by the expected number of returns to the initial vertex, emerging as a central question of interest in the quantum context, and can be directly connected with the classical recurrence criteria for random walks in $\mathbb{Z}$. For example, consider the classical homogeneous random walk on $\mathbb{Z}$ with graph given by Figure 1. It is well known that this lazy random walk with $r,l\geq 0,\;b>0,\;l+b+r=1,$ is recurrent if and only if the transition parameters $r$ (rate of jumping to right) and $l$ (rate of jumping to left) are equal. Also, if $b=0,$ then the walk is non-lazy and it is recurrent if and only if $r=1/2=l$ [14].
[…]
Figure 1: Classical Homogeneous Lazy Random Walk on $\mathbb{Z}$.
In the context of discrete-time Markov chains, a state is called recurrent if, once the chain enters that state, the probability of returning to it is 1. This is equivalent to saying that the average number of returns to that state is infinite, indicating that the chain will revisit the state indefinitely [14].
With this description, it is natural to seek criteria to determine under what conditions a HOQW is recurrent given $L,B,R.$ This problem has been addressed in the literature with the development of specific criteria, which are essential to understanding the dynamics and long-term behavior of these systems. See for instance [8, 13, 12]. These works define recurrence in terms of probability of return, where a state is said to be recurrent if the probability of the walk returning to it is equal to 1. However, it has been shown that for OQWs, the two definitions of recurrence based on returning with probability 1 is not equivalent to the one based on the average number of returns being infinite [3]. This property highlights a distinct characteristic of OQWs compared to classical Markov chains.
Once this difference is understood, we discuss the recurrence criteria for OQWs, adopting the definition proposed in [3]. Specifically, we analyze the recurrence of OQWs based on the average number of returns.
Furthermore, recurrence criteria concerning the mean number of returns for higher-order HOQWs have been presented in [11]. The authors of that work extend the analysis of recurrence to a more general class of open quantum walks, such as irreducible coins.
In this article, we present more general criteria for the recurrence of lazy and non-lazy HOQWs, based on the transition operators $L,B,R$ when acting on finite-dimensional vector spaces, that is, we allow the HOQW to be reducible. Furthermore, we present the first recurrence criterion for lazy HOQWs, using results obtained from recent studies on Central Limit Theorems for this kind of HOQWs.
A first Central Limit Theorem (CLT) for Open Quantum Walks (OQWs) was established in [1]. In this work, we rely on more recent developments, including the CLT for Lazy OQWs presented in [10], as well as the general CLT for homogeneous OQWs (HOQWs) on lattices with finite internal degrees of freedom given in [6].
This work is structured as follows. In Section 2 we define the open quantum walks induced by a coin and their transition probabilities. Section 3 is devoted to presenting a recurrence criterion for a class of Lazy HOQWs, and one restricted to Lazy HOQWs of dimension 2. In Section 4 we focus on the case where the coin is non-lazy and we present a general recurrence criterion for finite-dimensional coins. Section 5 illustrates the results with comprehensive examples.
2 General Settings
Let us consider the set of vertices $\mathbb{Z},$ and $\mathcal{K}=\mathbb{C}^{\mathbb{Z}}$ be the state space of a quantum system with as many degrees of freedom as the number of vertices. To emphasize the position of the particle of our quantum walk, we consider a quantum system described by the separable complex Hilbert space
$\mathcal{H}=\bigoplus_{i\in\mathbb{Z}}\mathfrak{h}_{i},$
where the $\mathfrak{h}_{i}$ are separable Hilbert spaces. The label $i\in\mathbb{Z}$ represents the position of the particle and, when the particle is located at the vertex $i\in\mathbb{Z}$, its internal state is encoded in the space $\mathfrak{h}_{i},$ as it will be detailed below.
We consider the set of density operators over a finite Hilbert space $\mathcal{W},$
$\mathcal{D}(\mathcal{W})=\{\rho:\mathcal{W}\rightarrow\mathcal{W}:\;\rho\geq 0,\;\mathrm{Tr}(\rho)=1\},$
and introduce
$\mathcal{S}(\mathcal{H}\otimes\mathcal{K})=\left\{\tau\in\mathcal{I}(\mathcal{H}\otimes\mathcal{K}):\textstyle\tau=\displaystyle\sum_{j\in\mathbb{Z}}\rho(j)\otimes\lvert j\rangle\langle j\rvert,\quad\rho(j)\geq 0,\quad\sum_{j\in\mathbb{Z}}\mathrm{Tr}\left(\rho(j)\right)=1\right\},$
the set of density operators acting on $\mathcal{H}\otimes\mathcal{K}.$ We are interested in HOQWs of the form
$\Phi(\tau)=\sum_{i\in\mathbb{Z}}\Big{(}L\rho(i+1)L^{*}+B\rho(i)B^{*}+R\rho(i-1)R^{*}\Big{)}\otimes|i\rangle\langle i|,\;\;\;\tau\in\mathcal{S}(\mathcal{H}\otimes\mathcal{K}),$ (1)
where $L,B,R\in\mathbb{M}_{d}({\mathbb{C}})$ satisfy $L^{*}L+B^{*}B+R^{*}R=I_{d}.$
We say that the triple of matrices $(L,B,R)$ is a coin if eq. (1) holds. Therefore, we can assume that $\mathfrak{h}_{i}=\mathfrak{h}\;\forall i\in\mathbb{Z},$ and if $\mathfrak{h}=\mathbb{C}^{d}$ for some $d$, the dimension of the coin is said to be $d$. In this article, we consider only the case $d<\infty.$
If $B\neq 0,$ the quantum walk exhibits a lazy behavior, meaning that in the context of quantum walks there is a non-zero probability for the walker to stay in the same vertex after a step. We refer to the HOQW above as a HOQW induced by the coin $(L,B,R)$ in which the walk is inherently lazy.
Therefore, this coin represents a homogeneous, nearest-neighbor OQW on the integer line with the left and right transitions given by $L$ and $R$, respectively, and $B$ represents the rate of no jump at some instant.
This lazy behavior introduces distinct properties compared to its non-lazy counterpart. Notably, the presence of the matrix $B$ influences the walk’s propensity to remain in its current vertex, thereby affecting the overall dynamics. If $B=0,$ then $\Phi$ will be called a HOQW induced by a coin $(L,R)$ or a non-lazy HOQW induced by a coin $(L,R).$ The dynamic described above is represented by the graph in Figure 2.
[…]
Figure 2: Lazy Quantum Walk on $\mathbb{Z}$.
The quantum trajectory of this kind of OQW on the line, induced by a coin $(L,B,R)$ and starting from a state $\tau$ of the form $\tau=\sum_{i\in\mathbb{Z}}\tau_{i}\otimes|i\rangle\langle i|,$ is any path generated by the Markov chain $(X_{n},\tau_{n})_{n\geq 0}$, where $X_{n}$ denotes the position of the particle at time $n$ and $\tau_{n}$ its internal degree. The transition probabilities are given by
$\begin{split}\mathbb{P}\left(\,(X_{n+1},\tau_{n+1})=\left(i-1,\frac{L\sigma L^{*}}{\mathrm{Tr}(L\sigma L^{*})}\right)\;\Bigg{|}\;(x_{n},\tau_{n})=(i,\sigma)\,\right)&=\mathrm{Tr}(L\sigma L^{*}),\\ \mathbb{P}\left(\,(X_{n+1},\tau_{n+1})=\left(i,\frac{B\sigma B^{*}}{\mathrm{Tr}(B\sigma B^{*})}\right)\;\Bigg{|}\;(x_{n},\tau_{n})=(i,\sigma)\,\right)&=\mathrm{Tr}(B\sigma B^{*}),\\ \mathbb{P}\left(\,(X_{n+1},\tau_{n+1})=\left(i+1,\frac{R\sigma R^{*}}{\mathrm{Tr}(R\sigma R^{*})}\right)\;\Bigg{|}\;(x_{n},\tau_{n})=(i,\sigma)\,\right)&=\mathrm{Tr}(R\sigma R^{*}),\end{split}$
for every $i\in\mathbb{Z}$, $\sigma\in\mathcal{D}(\mathfrak{h})$, and initial law
$\mathbb{P}\left(\,(X_{0},\tau_{0})=\left(i,\frac{\tau_{i}}{\mathrm{Tr}\,\tau_{i}}\right)\;\right)=\mathrm{Tr}(\tau_{i}),$
and all other transition probabilities are null. Therefore, at time $n$, if the system is located at site $\lvert i\rangle$ with an initial density $\rho\otimes\lvert i\rangle\langle i\rvert$, then at the next time step $n+1$, one of the following occurs:
  • The particle moves to site $\lvert i-1\rangle$ (left) with probability $\operatorname{Tr}(L_{i}\rho L_{i}^{*})$, and the density evolves to
    $\frac{L_{i}\rho L_{i}^{*}}{\operatorname{Tr}(L_{i}\rho L_{i}^{*})}\otimes\lvert i-1\rangle\langle i-1\rvert;$
  • The particle remains at site $\lvert i\rangle$ with probability $\operatorname{Tr}(B_{i}\rho B_{i}^{*})$, and the new density matrix is
    $\frac{B_{i}\rho B_{i}^{*}}{\operatorname{Tr}(B_{i}\rho B_{i}^{*})}\otimes\lvert i\rangle\langle i\rvert;$
  • The particle moves to site $\lvert i+1\rangle$ (right) with probability $\operatorname{Tr}(R_{i}\rho R_{i}^{*})$, leading to the new density matrix
    $\frac{R_{i}\rho R_{i}^{*}}{\operatorname{Tr}(R_{i}\rho R_{i}^{*})}\otimes\lvert i+1\rangle\langle i+1\rvert.$
3 Recurrence criteria for Lazy Open Quantum Walks
3.1 Faithfulness of Density Operators
Density operators play a fundamental role in open quantum systems. In this work, they provide statistical descriptions of HOQWs. Faithful and non-faithful density operators are characterized by their positive definiteness and positive semi-definiteness, respectively, when they evolve in finite-dimensional systems.
Faithful density operators are defined as being positive definite. This means that the eigenvalues of a faithful density operator are strictly positive, and the operator is invertible. Essentially, faithfulness ensures a complete representation of the quantum system, capturing the full richness of its physical properties.
On the other hand, non-faithful density operators are positive semi-definite, thus they have $\lambda=0$ as one of their eigenvalues. The presence of zero eigenvalues implies that the non-faithful density operator is not invertible and may represent specific mixed states or situations where information loss occurs. The positive semi-definiteness criterion relaxes the strictness of positive definiteness, allowing for a broader range of quantum state representations, while still maintaining the one-to-one correspondence between quantum states and density operators.
An important subset of the set of non-faithful density operators is composed of the pure density operators. A pure density is a density $\rho$ that is a projection of rank 1, i.e., there exists some vector $\lvert u\rangle\in\mathbb{C}^{d}$ such that $\rho=\lvert u\rangle\langle u\rvert$.
Let us consider an OQW starting the walk on vertex $\lvert i\rangle$ with initial density operator $\rho.$ For evolution on time $n=0,1,2,\ldots,$ we define
$p_{n}(j)=\mathbb{P}(\mbox{the quantum particle, at time }n,\mbox{ is in site}\;\lvert j\rangle)=\mathrm{Tr}(\rho_{n}(j)),$
where
$\Phi^{n}(\rho\otimes\lvert i\rangle\langle i\rvert)=\sum_{s\in\mathbb{Z}}\rho_{n}(s)\otimes\lvert s\rangle\langle s\rvert.$
With this notation,
$p_{ji;\rho}(n)=p_{n}(\rho\otimes|i\rangle\to|j\rangle)=\mathrm{\mathrm{Tr}}(\rho_{n}(j)\otimes\lvert j\rangle\langle j\rvert)=\mathrm{\mathrm{Tr}}\left(\Phi^{n}(\rho\otimes\lvert i\rangle\langle i\rvert)(I\otimes\lvert j\rangle\langle j\rvert)\right)$
denote the probability of being at site $\lvert j\rangle$ at step $n,$ given that we started at site $\lvert i\rangle,$ with initial density $\rho$ concentrated at $\lvert i\rangle.$ Therefore, the dynamics starts with a density operator $\rho$ concentrated at some vertex $\lvert i\rangle,$ takes the evolution up to time ($n$) through the linear operator $\Phi^{n},$ producing a new density operator
$\rho_{n}=\sum_{s}\rho_{n}(s)\otimes\lvert s\rangle\langle s\rvert=\Phi^{n}(\rho\otimes\lvert i\rangle\langle i\rvert),\quad\mathrm{Tr}\left(\sum_{s}\rho_{n}(s)\right)=1,$
and then we project $\rho_{n}$ onto the subspace generated by vertex $\lvert j\rangle.$ Note that
$\Phi^{n}(\rho\otimes\lvert i\rangle\langle i\rvert)(I\otimes\lvert j\rangle\langle j\rvert)=\rho_{n}(j)\otimes\lvert j\rangle\langle j\rvert,$
which represents the data concentrated at vertex $\lvert j\rangle$ at time $n.$
Consider an OQW on some set of vertices $\mathbb{Z}$, let $i\in\mathbb{Z}$ and $\rho\in\mathcal{D}(\mathfrak{h})$. We say that a vertex $\lvert i\rangle$ is
  1. 1.
    $\rho$-recurrent if
    $\sum_{n=0}^{\infty}p_{ii;\rho}(n)=\infty.$ (2)
    Otherwise, $\lvert i\rangle$ is said to be $\rho$-transient;
  2. 2.
    recurrent, if $\lvert i\rangle$ is $\rho-$recurrent for all $\rho\in\mathcal{D}(\mathfrak{h});$
  3. 3.
    transient, if $\lvert i\rangle$ is $\rho-$transient for all $\rho\in\rho\in\mathcal{D}(\mathfrak{h}).$
Concerning all vertices, an OQW is called: recurrent if every vertex is recurrent; transient if every vertex is transient. Note that equation (2) computes the mean number of returns to site $\lvert i\rangle.$
The following Proposition describes $\rho$-recurrence in a concentrated vertex $i\in\mathbb{Z}$ when assumed that $\lvert i\rangle$ is recurrent for some density and $\mathfrak{h}$ has a finite internal degree of freedom. An immediate consequence will also be given for the case where $\lvert i\rangle$ is $\sigma$-transient for some non-faithful density.
Theorem 3.1.
Consider an OQW on some set of vertices $V,$ let $i\in V$ and assume that $\mathfrak{h}$ is finite-dimensional. If $\lvert i\rangle$ is $\tau$-recurrent for some $\tau\in\mathcal{D}(\mathfrak{h}),$ then
  1. 1.
    vertex $\lvert i\rangle$ is $\rho$-recurrent for all faithful $\rho\in\mathcal{D}(\mathfrak{h})$;
  2. 2.
    there exists a pure density operator $\lvert w\rangle\langle w\rvert$ such that the OQW is $\lvert w\rangle\langle w\rvert$-recurrent, where $\lvert w\rangle$ is an unit vector of $\mathfrak{h}.$
Proof.
Let $\rho\in\mathcal{D}(\mathfrak{h})$ be faithful. Since the state space $\mathcal{D}(\mathfrak{h})$ is compact in the trace-norm topology, the function
$f(\tau)=\inf\{c>0\mid\rho>c\tau\}$
is continuous and thus attains its minimum and maximum. In particular, there exists $c>0$ such that
$\rho>c\tau,\quad\forall\rho\in\mathcal{D}(\mathfrak{h}).$
thus
$\begin{split}\sum_{n=0}^{\infty}p_{ii;\rho}(n)=&\sum_{n=0}^{\infty}\textmd{Tr}\left[\Phi^{n}\left(\rho\otimes\lvert i\rangle\langle i\rvert\right)\right(I\otimes\lvert i\rangle\langle i\rvert)]\\ >&c\sum_{n=0}^{\infty}\textmd{Tr}\left[\Phi^{n}\left(\tau\otimes\lvert i\rangle\langle i\rvert\right)(I\otimes\lvert i\rangle\langle i\rvert)\right]\\ =&c\sum_{n=0}^{\infty}p_{ii;\tau}(n)\\ =&\infty,\end{split}$
giving the first item.
For the second one, let
$\tau=\sum_{j}\lambda_{j}\lvert e_{j}\rangle\langle e_{j}\rvert,$
where each $\lambda_{j}>0$ is an eigenvalue of $\tau$ with corresponding eigenvector $\lvert e_{j}\rangle.$ Then
$\infty=\sum_{n=0}^{\infty}\textmd{Tr}\left[\Phi^{n}\left(\tau\otimes\lvert i\rangle\langle i\rvert\right)\right(I\otimes\lvert i\rangle\langle i\rvert)]=\sum_{j}\left(\lambda_{j}\sum_{n=0}^{\infty}\textmd{Tr}\left[\Phi^{n}\left(\lvert e_{j}\rangle\langle e_{j}\rvert\otimes\lvert i\rangle\langle i\rvert\right)\right(I\otimes\lvert i\rangle\langle i\rvert)]\right),$
where at least one of the members of the sum must be infinity. This completes the proof.
$\square$
By a contra-position argument, we obtain the following direct consequence of Theorem 3.1.
Corollary 3.2.
Let $\Phi$ be an OQW on some set of vertices $V,$ $i\in V,$ and $\dim(\mathfrak{h})<\infty.$ If $\lvert i\rangle$ is $\tau$-transient for some faithful $\tau\in\mathcal{D}(\mathfrak{h}),$ then $\lvert i\rangle$ is transient.
Another direct consequence of Theorem 3.1 is related to the homogeneous case.
Corollary 3.3.
Let $\Phi$ be a HOQW on $\mathbb{Z}$ with $\dim(\mathfrak{h})=d<\infty.$ If $\Phi$ is $\tau$-recurrent for some $\tau\in\mathcal{D}(\mathfrak{h}),$ then
  1. 1.
    $\Phi$ is $\rho$-recurrent for all faithful $\rho\in\mathcal{D}(\mathfrak{h})$;
  2. 2.
    there exists a pure density operator $\lvert w\rangle\langle w\rvert$ such that $\Phi$ is $\lvert w\rangle\langle w\rvert$-recurrent, where $\lvert w\rangle$ is an unit vector of $\mathbb{C}^{d}.$
3.2 Enclosures and Invariant States
Firstly, a subspace $\mathcal{Y}$ of $\mathfrak{h}$ is called an enclosure for a quantum channel $\mathcal{L}$ if for every $\sigma\in\mathcal{D}\mathfrak{h}),$ we have
$\mathrm{supp}(\sigma)\subset\mathcal{Y}\Rightarrow\mathrm{supp}\left(\mathcal{L}(\sigma)\right)\subset\mathcal{Y}.$
If an enclosure contains no other non-trivial enclosures, then it is called a minimal enclosure.
A minimal enclosure with finite dimension is the support of a unique invariant state of the channel, that is, a density operator $\sigma$ such that $\mathcal{L}(\sigma)=\sigma$. For more on enclosures of OQWs, we refer the reader to [6, 7, 3].
Following this property and using the homogeneity of HOQWs induced by a coin, we split $\mathfrak{h}$ into two subsets $\mathcal{Y}$ and $\mathcal{X}$ in such a way that the walk is transient for $\mathcal{D}(\mathcal{Y})$ and recurrent everywhere else, with $\mathcal{Y}$ being an enclosure of the quantum channel with Kraus operators $L,B,R.$
Proposition 3.4.
Consider a HOQW induced by a coin $(L,B,R)$ of dimension $d$. We are in one (and only one) of the following situations:
  • $(L,B,R)$ is recurrent;
  • $(L,B,R)$ is transient;
  • $(L,B,R)$ is recurrent with respect to all densities except for a convex set $\mathcal{D}(\mathcal{Y}).$ In this case, $\mathcal{Y}$ is an enclosure of $\mathcal{L},$
    $\mathfrak{h}=\mathcal{Y}\oplus\mathcal{X}$ (3)
    for some subspace $\mathcal{X}$ of $\mathfrak{h},$ with $\mathcal{Y},\mathcal{X}\neq\{0\}.$
Proof.
We claim $(L,B,R)$ is $\rho$-transient if and only if it is $\rho_{n}$-transient for every $n,$ where $\rho_{n}$ is the density at the $n$-th step of the walk.
Indeed, let us suppose the walk starts at vertex $\lvert 0\rangle$ with initial density $\rho.$ For instance, if the first jump is to $\lvert-1\rangle,$ the density, after this first jump, is
$\rho_{1}=\frac{L\rho L^{*}}{\mathrm{Tr}(L\rho L^{*})}.$
Since the walk is homogeneous, it is $\rho$-transient if and only if it is $\rho_{1}$-transient. After $n$ jumps, the density is
$\rho_{n}=\frac{J_{\pi}\rho J_{\pi}^{*}}{\mathrm{Tr}(J_{\pi}\rho J_{\pi}^{*})},$ (4)
where $J_{\pi}$ is a concatenation of $n$ operators, each one belonging to $\{L,B,R\},$ proving the claim, once this is valid for any $n=1,2,3,\ldots$
Moreover, by equation (4) we have that $\cup_{n}\;\text{supp}(\rho_{n})$ is a subspace of $\mathfrak{h}.$ Joining all the possible initial states in which $(L,B,R)$ is transient, we obtain a decomposition (3), where $\mathcal{Y}$ or $\mathcal{X}$ may be null, and by definition $\mathcal{Y}$ is an enclosure.
Let us suppose that $(L,B,R)$ is neither recurrent nor transient. By Theorem 3.1, there exist unit vectors $\lvert u\rangle,\lvert v\rangle\in\mathfrak{h}$ such that the walk is $\lvert u\rangle\langle u\rvert$-recurrent and $\lvert v\rangle\langle v\rvert$-transient. Let $\mathcal{Y}$ be the subspace of $\mathfrak{h}$ such that $\mathcal{D}(\mathcal{Y})$ is the set of densities of $\mathcal{B}(\mathfrak{h})$ in which $(L,B,R)$ is transient, thus we pick $\rho\in\mathcal{D}(\mathcal{Y})$ and we can affirm that $(L,B,R)$ is $\rho_{n}$-transient for every $n.$ Finally, take $\mathcal{X}=\mathcal{Y}^{\perp}$ and the trichotomy holds.
$\square$
In our subsequent discussion, we consider a coin $(L,B,R)$ and its auxiliary map $\mathcal{L}:\mathcal{B}(\mathfrak{h})\rightarrow\mathcal{B}(\mathfrak{h})$ given by
$\mathcal{L}(\rho)=L\rho L^{*}+B\rho B^{*}+R\rho R^{*},\quad\rho\in\mathcal{B}(\mathfrak{h}).$
Note that $\mathcal{L}$ is a quantum channel, thus it has at least one invariant state, that is, there exists some density operator $\rho_{\infty}\in\mathcal{B}(\mathfrak{h})$ satisfying $\mathcal{L}(\rho_{\infty})=\rho_{\infty}.$ When the invariant state of $\mathcal{L}$ is unique, we will say that the map $\mathcal{L}$ is ergodic.
We will use the following result obtained straightforwardly from [10].
Theorem 3.5.
Consider a Lazy HOQW induced by a coin $(L,B,R)$ of dimension $d.$ If the auxiliary map $\mathcal{L}$ admits only one invariant state $\rho_{\infty},$ then
$\frac{X_{n}}{n}\stackrel{{\scriptstyle n\rightarrow\infty}}{{\longrightarrow}}m,\qquad\text{where }\qquad m=\mathrm{Tr}\left(R\rho_{\infty}R^{*}\right)-\mathrm{Tr}\left(L\rho_{\infty}L^{*}\right),$
in probability.
Proof.
This result can be obtained by using a Doob decomposition as described in [10, Equation 27] and [1, Theorem 5.2] for a Martingale $M_{n}$ satisfying
$X_{n}-n.m=X_{0}+M_{n}+\beta_{n},$
where $M_{n}/n\rightarrow 0$ and $\beta_{n}\rightarrow 0$ when $n\rightarrow 0.$
$\square$
Theorem 3.6 (Chung-Fuchs Theorem for HOQWs with ergodic auxiliary map).
Consider a HOQW induced by a coin $(L,B,R),\;\mathcal{L}$ its auxiliary map and $\mathcal{Y}$ an enclosure. Suppose that the initial density operator is $\rho\in\mathcal{D}(\mathcal{Y})$ with quantum trajectories $(X_{n},\rho_{n})_{n\geq 0}$ and $X_{n}/n\rightarrow 0$ in probability, then $(L,B,R)$ is $\sigma$-recurrent for some $\sigma\in\mathcal{D}(\mathcal{Y}).$
Proof.
The proof is analogous to [11, Theorem 7] and we will proceed by remarking on how it can be generalized. Firstly, the maximum appearing in [11, Lemma 6] can be taken in $\mathcal{D}=\mathcal{D}(\mathcal{Y}).$ Indeed, if the initial state is $\rho\in\mathcal{D}(\mathcal{Y}),$ then $\rho_{n}\in\mathcal{D}(\mathcal{Y})$ for every $n,$ thereby the maximum taken in the proof of [11, Lemma 6] can be reduced to $\mathcal{D}(\mathcal{Y}).$ Therefore, by the same reason, the density operator considered in [11, Theorem 7] belongs to $\mathcal{D}(\mathcal{Y}),$ proving the Theorem for the non-lazy case. Moreover, the lazy case holds as an extension of this proof, since all the results work for homogeneous OQWs on $\mathbb{Z}$.
$\square$
The following theorem presents the first recurrence criterion for non-irreducible HOQWs of arbitrary finite dimension. This result extends [11, Corollary 8 and Proposition 8] for reducible CTOQWs and can also be applied to Lazy HOQWs.
Theorem 3.7 (Recurrence Criteria).
Let us consider a HOQW on $\mathbb{Z}$ induced by a coin $(L,B,R)$ of dimension $d$ such that the auxiliary map $\mathcal{L}$ is ergodic. Then $\mathcal{L}$ has a unique invariant state $\rho_{\infty}$ and
  • $\textmd{Tr}\left(L\rho_{\infty}L^{*}\right)=\textmd{Tr}\left(R\rho_{\infty}R^{*}\right)\Rightarrow$ the walk is recurrent;
  • $\textmd{Tr}\left(L\rho_{\infty}L^{*}\right)\neq\textmd{Tr}\left(R\rho_{\infty}R^{*}\right)\Rightarrow$ the walk is transient.
Proof.
We suppose that the walk is neither recurrent nor transient. In this case, the walk is recurrent for some density and transient for all densities in some $\mathcal{D}(\mathcal{Y}),$ where $\mathcal{Y}$ is an enclosure (Proposition 3.4). Moreover, the unique invariant state $\rho_{\infty}$ of $\mathcal{L}$ must belong to $\mathcal{D}(\mathcal{Y}).$
If $\;\textmd{Tr}\left(L\rho_{\infty}L^{*}\right)=\textmd{Tr}\left(R\rho_{\infty}R^{*}\right),$ then $X_{n}/n\rightarrow 0$ by Theorem 3.5, thus $(L,B,R)$ is $\rho$-recurrent for some $\rho\in\mathcal{D}(\mathcal{Y})$ by Theorem 3.6, which is a contradiction.
[…]
Therefore, if $\textmd{Tr}\left(L\rho_{\infty}L^{*}\right)=\textmd{Tr}\left(R\rho_{\infty}R^{*}\right),$ then the walk is $\rho$-recurrent for some density and thus it is recurrent. If $\textmd{Tr}\left(L\rho_{\infty}L^{*}\right)\neq\textmd{Tr}\left(R\rho_{\infty}R^{*}\right),$ then the walks is transient for every density.
$\square$
We remark that the operator $B$ does influence in the recurrence of $(L,B,R).$ Indeed, even it does not appear on the bullet points that describe the recurrence criterion, the invariant state $\rho_{\infty}$ may change when we pick distinct $B$’s. See example 5.3 below.
3.3 General Recurrence Criteria for HOQWs induced by a coin $(L,B,R)$ of dimension 2
The following theorem is due to [4].
Theorem 3.8.
For a finite quantum channel $\mathcal{L}:\mathcal{B}(\mathfrak{h})\rightarrow\mathcal{B}(\mathfrak{h})$, the following are equivalent:
i.
$\mathcal{L}$ is ergodic;
ii.
$\mathcal{L}$ does not have two invariant subspaces $S_{1}\neq\{0\}$ and $S_{2}\neq\{0\}$ such that $S_{1}\cap S_{2}=\{0\};$
iii.
$\mathcal{L}$ has a minimal invariant subspace, that is, a subspace $S\neq\{0\}$ such that $S\subset S^{\prime}$ for every invariant subspace $S^{\prime}\neq\{0\}.$
Building upon the established preceding proposition, we will advance to the following theorem, delineating a comprehensive criterion for 2-dimensional coins, extending [11, Theorem 17] to encompass the Lazy case. This criterion will be structured into two distinct parts: the first adhering to the logic of the previous theorem, and the second articulated in terms of the eigenvalues of the operators of the coin, independent of the invariant densities of the auxiliary map.
Theorem 3.9.
Consider a HOQW on $\mathbb{Z}$ induced by a coin $(L,B,R)$ of dimension 2.
(1) If $L,B,R$ have at most one common eigenvector, let $\rho_{\infty}$ be the unique invariant density of the auxiliary map. Then, we have
$(1.1)\quad\textmd{Tr}(L\rho_{\infty}L^{*})\neq\textmd{Tr}(R\rho_{\infty}R^{*})\Rightarrow(L,B,R)\mbox{ is transient},$
$(1.2)\quad\textmd{Tr}(L\rho_{\infty}L^{*})=\textmd{Tr}(R\rho_{\infty}R^{*})\Rightarrow(L,B,R)\mbox{ is recurrent}.$
(2) If $L,B,R$ have two linearly independent eigenvectors in common, $\lvert u_{1}\rangle,\lvert u_{2}\rangle$ unit, then they satisfy $\lvert u_{1}\rangle\bot\lvert u_{2}\rangle,$ we can put
$L\lvert u_{j}\rangle=l_{j}\lvert u_{j}\rangle,\qquad B\lvert u_{j}\rangle=b_{j}\lvert u_{j}\rangle,\qquad R\lvert u_{j}\rangle=r_{j}\lvert u_{j}\rangle,\qquad j=1,2,$
and
$(2.1)\quad|l_{1}|=|r_{1}|\mbox{ and }|l_{2}|=|r_{2}|$$\Rightarrow$ $(L,B,R)\mbox{ is recurrent,}$,
$(2.2)\quad|l_{1}|\neq|r_{1}|\mbox{ and }|l_{2}|\neq|r_{2}|$$\Rightarrow$ $(L,B,R)\mbox{ is transient,}$,
$(2.3)\quad|l_{i}|=|r_{i}|\mbox{ and }|l_{j}|\neq|r_{j}|$$\Rightarrow$ $(L,B,R)\mbox{ is transient with respect to }\sigma_{j}=\lvert u_{j}\rangle\langle u_{j}\rvert\mbox{ and it is recurrent}$
[…]
Proof.
If the elements of the coin do not share at least two orthogonal eigenvectors, then the auxiliary map is ergodic, thus item (1) is a particular case of Theorem 3.7.
Now we suppose that $L,B,R$ share two linearly independent eigenvectors, then we can apply Theorem 3.8 to obtain $\lvert u_{1}\rangle,\lvert u_{2}\rangle$ common orthonormal eigenvectors of $L,B,R.$ We assume that $|l_{1}|=|r_{1}|$ and $|l_{2}|=|r_{2}|,$ then for $\rho=\lvert u\rangle\langle u\rvert,$ where $\lvert u\rangle$ is some common eigenvector of $L,B,R,$ we have $p_{j,j+1;\rho}(n)=p_{j,j-1;\rho}(n)=w$ for all $j\in\mathbb{Z}$ and all $n\geq 0$ for some constant $w$ and therefore, in this case, we have a classical homogeneous walk which is recurrent. If $(L,B,R)$ was not transient, then we would be in third situation of proposition 3.4, with
$\mathfrak{h}=\mathcal{Y}\oplus\mathcal{X},\quad\text{where}\quad\text{dim}(\mathcal{Y})\geq 1,\;\;\text{ and }\;\;\text{dim}(\mathcal{X})=2,$
a contradiction. Therefore, we are in the first situation of proposition 3.4 and the walk is recurrent.
Analogously, if $|l_{1}|\neq|r_{1}|$ and $|l_{2}|\neq|r_{2}|,$ then for $\rho_{1}=\lvert u_{1}\rangle\langle u_{1}\rvert,$ and $\rho_{2}=\lvert u_{2}\rangle\langle u_{2}\rvert,$ where $\lvert u_{1}\rangle,\lvert u_{2}\rangle$ are orthonormal common eigenvectors of $L,B,R,$ we have $p_{j,j+1;\rho}(n)<p_{j,j-1;\rho}(n)$ (or $>$) for all $j\in\mathbb{Z}$ and all $n\geq 0$ and therefore in this case we have a classical homogeneous walk which is transient for both $\rho_{1},\rho_{2}$. Since the walk is transient for two distinct eigenvectors, the walk is transient by corollary 3.2.
For the last case, $|l_{i}|=|r_{i}|$ and $|l_{j}|\neq|r_{j}|,$ $(L,B,R)$ is $\rho_{i}$-recurrent and $\rho_{j}$-transient, where $\rho_{i}=\lvert u_{i}\rangle\langle u_{i}\rvert$ and $\rho_{j}=\lvert u_{j}\rangle\langle u_{j}\rvert.$ This situation is a particular case of the last item in Proposition 3.4.
$\square$
4 A General Recurrence Criteria for Open Quantum Walks
Following the approach described in [6, Section 4], as well as the references therein, we can obtain for a coin $(L,R)$ a decomposition
$\mathfrak{h}=\mathfrak{R}\oplus\mathfrak{T},\qquad\mathfrak{R}=\sup\{\mbox{supp}(\omega)\;|\;\omega\mbox{ is an invariant state for }\mathcal{L}\}.$
Since $\mathfrak{h}$ is finite-dimensional, any minimal enclosure is included in $\mathfrak{R},$ and is the support of a unique extremal invariant state. Moreover, we have a unique decomposition of $\mathfrak{R}$ of the form
$\mathfrak{R}=\bigoplus_{\alpha\in A}\mathfrak{X}_{\alpha},$
where $\{\mathfrak{X}_{\alpha}\}_{\alpha}\in A$ is a finite set of mutually orthogonal enclosures and every $\mathfrak{X}_{\alpha}$ is minimal in the set of enclosures satisfying
[…]
Further, every $\mathfrak{X}_{\alpha}$ either is a minimal enclosure or assumes a decomposition (not unique) as the sum of mutually orthogonal isomorphic minimal enclosures $\{\mathfrak{V}_{\alpha,\beta},\;\beta\in I_{\alpha}\}:$
$\mathfrak{X}_{\alpha}=\bigoplus_{\beta\in I_{\alpha}}\mathfrak{V}_{\alpha,\beta}.$
Given a minimal enclosure $\mathfrak{V}_{\beta},$ let $\tau_{\beta}$ be the unique invariant state with support on $\mathfrak{V}_{\beta}.$ Introducing the parameter
$m_{\beta}=\mathrm{Tr}\left(L\tau_{\beta}L^{*}\right)-\mathrm{Tr}\left(R\tau_{\beta}R^{*}\right),$
we have by [6, Lemma 4.2] that $m_{\beta}$ is independent of the particular minimal enclosure $\mathfrak{V}_{\beta}$ in $\mathfrak{X}_{\alpha}.$ Therefore, we can combine Proposition 3.4, Theorem 3.7 and the discussion above to state the property
Proposition 4.1.
Consider a coin $(L,R).$ We have a decomposition
$\mathfrak{h}=\left(\bigoplus_{\alpha\in A}\mathcal{Y}_{\alpha}\right)\oplus\mathcal{X},$ (5)
where $A$ is a finite set, $\mathcal{Y}_{\alpha}$ is a subspace of $\mathfrak{h}$ being the support of exactly one invariant state $\tau_{\alpha}$ for $\mathcal{L},$ $\mathcal{X}$ contains no enclosure, and
  • $\textmd{Tr}\left(L\tau_{\alpha}L^{*}\right)=\dfrac{1}{2}\Rightarrow$ the walk is $\tau_{\alpha}$-recurrent for all $\tau\in\mathcal{D}(\mathcal{Y}_{\alpha})$;
  • $\textmd{Tr}\left(L\tau_{\alpha}L^{*}\right)\neq\dfrac{1}{2}\Rightarrow$ the walk is $\tau_{\alpha}$-transient for all $\tau\in\mathcal{D}(\mathcal{Y}_{\alpha}).$
Once the decomposition on Equation (5) is obtained, we can say that $\mathcal{Y}_{\alpha}$ is either recurrent or transient.
The following result is due to [6] and will be crucial to our last criterion. To state it, consider a coin $(L,R),$ and let $\mathcal{Y}$ be an enclosure for $\mathcal{L}.$ Denoting $p_{\mathcal{Y}}$ the orthogonal projection onto $\mathcal{Y},$ we can define the associated absorption operator (see [5]):
$A(\mathcal{Y}):=\lim_{n\rightarrow\infty}\mathcal{L}^{*n}(p_{\mathcal{Y}}).$
Corollary 4.2.
Consider a coin $(L,R)$ with the decomposition of $\mathfrak{h}$ given on Equation (5), and let $a_{\alpha}(\rho)=\mathbb{E}_{\rho}\left[\mathrm{Tr}\left(A(\mathcal{Y}_{\alpha})\rho_{0}\right)\right].$ Then we have the convergence in law
$\frac{X_{n}-X_{0}}{n}\stackrel{{\scriptstyle n\rightarrow\infty}}{{\longrightarrow}}\sum_{\alpha\in A}a_{\alpha}(\rho)\delta_{m_{\alpha}},$ (6)
where $\delta_{m_{\alpha}}$ is the Dirac measure concentrated in $m_{\alpha}.$
We remark that the convergence in law brought by corollary 4.2 can be upgraded to a convergence a.s., as demonstrated in [9, Theorem 3.5.2].
Let us start a HOQW induced by a coin $(L,R).$ There are two possibilities concerning recurrence when $\text{supp}(\rho_{0})\not\perp\mathcal{X}:$
1. $(L,R)$ is recurrent for $\mathcal{Y}_{\alpha}$ for some $\alpha$ and supp$\left(\mathcal{L}^{n}(\rho_{0})\right)\not\perp\mathcal{Y}_{\alpha}$ for some $n\in\mathbb{N}$: in this case, the walk is $\rho_{0}$-recurrent. Indeed, concerning the quantum trajectories $(\rho_{n})_{n}$, we could refresh the walk on some instant $n$ at some vertex $\lvert X_{n}\rangle,$ which is $\rho_{n}$-recurrent, since supp$\left(\mathcal{L}^{n}(\rho_{0})\right)\not\perp\mathcal{Y}_{\alpha}$ for a sufficiently large $n$. If there was some density $\sigma$ with supp$(\sigma)\not\perp\mathcal{X}$ and $\sigma$-transient, there would be some enclosure not orthogonal to $\mathcal{X},$ which is a contradiction by Proposition 3.4. The conclusion is held by the homogeneity of the coin, since, in this case, the walk either is $\rho_{s}$-recurrent for every $s=0,1,2,\ldots$ or it is $\rho_{s}$-transient for every $s=0,1,2,\ldots$, analogously to the proof of Proposition 3.4.
2. $(L,R)$ is transient for all $\mathcal{Y}_{\alpha}$ satisfying supp$\left(\mathcal{L}^{n}(\rho_{0})\right)\not\perp\mathcal{Y}_{\alpha}$ for every $n\in\mathbb{N}$: the walk is $\sigma$-transient for all $\sigma\in\mathcal{D}(\mathcal{X})$. Indeed, in this case, starting the walk on vertex $X_{0}=0,$ $X_{n}/n$ converges (when $n\rightarrow\infty$) to
$m=\sum_{\alpha\in A}a_{\alpha}(\rho)\delta_{m_{\alpha}}=\sum_{\alpha\in A}a_{\alpha}(\rho),$
as given by Equation (6), where the second equality follows from $m_{\alpha}\neq 0,$ since $(L,R)$ is $\mathcal{Y}_{\alpha}$ transient. Thus, $X_{n}$ converges either to $\infty$- ∞ or $\infty$,∞ , in probability, imposing that the walk is $\rho$-transient and thus $(L,R)$ is transient on $\mathcal{D}(\mathcal{X}).$
Therefore, we have shown the following recurrence criterion:
Theorem 4.3 (A Generalized Criteria).
Consider a coin $(L,R)$ of dimension $d.$ Consider the decomposition of $\mathfrak{h}$ as given in Equation (5), and $\tau_{1},\ldots,\tau_{m}$ the respectively invariant states of $\mathcal{L}$ in $\mathcal{Y}_{\alpha},\alpha=1,\ldots,m.$
$(1)\quad\mathrm{Tr}(L\tau_{j}L^{*})\neq\dfrac{1}{2}\;\forall j\in\{1,\ldots m\}\Rightarrow(L,R)\mbox{ is transient;}$;
$(2)\quad\mathrm{Tr}(L\tau_{j_{k}}L^{*})=\dfrac{1}{2}\;$ for a set of indices $j_{k}\in J\subset\{1,\ldots,m\}$ and $\mathrm{Tr}(L\tau_{j_{l}}L^{*})\neq\dfrac{1}{2}\;$ for the remaining indices $j_{l}\in J^{C},$ and $\sigma$ is a density such that supp$(\sigma)\not\perp\mathcal{X}.$
(2.1) supp$\left(\mathcal{L}^{n}(\sigma)\right)\not\perp\mathcal{Y}_{\alpha}$ for some $\alpha\in J$ and some $n\in\mathbb{N},$ then $(L,R)$ is
$\begin{cases}\rho\text{-transient for }\rho\in\mathcal{D}_{1}:=\mathcal{D}(\oplus_{k\in J^{C}}\mathcal{Y}_{k})\\ \rho\text{-recurrent for }\rho\in\mathcal{D}_{1}^{C}.\end{cases}$
(2.2) supp$\left(\mathcal{L}^{n}(\rho)\right)\perp\mathcal{Y}_{\alpha}$ for all $\alpha\in J$ and all $n\in\mathbb{N},$ then $(L,R)$ is
$\begin{cases}\rho\text{-transient for }\rho\in\mathcal{D}_{2}:=\mathcal{D}(\oplus_{k\in J^{C}}\mathcal{Y}_{k}\oplus\mathcal{X})\\ \rho\text{-recurrent for }\mathcal{D}_{2}^{C}\end{cases}$
$(3)\quad\mathrm{Tr}(L\tau_{\alpha}L^{*})=\dfrac{1}{2}\;$ for every $\alpha\Rightarrow$ the walk is recurrent.
Remark 4.4.
It is worth observing that $\mathcal{D}_{2}$ in Theorem 4.3 actually arises from an enclosure due to the orthogonality between supp$\left(\mathcal{L}^{n}(\rho)\right)$ and $\mathcal{Y}_{\alpha}$; otherwise, it would contradict Proposition 3.4.
5 Examples
Example 5.1.
Our first examples will consider lazy HOQWs of dimension 2.
  1. 1.
    Let $m,n>0$ and assume that $b:=1-m^{2}-n^{2}\in(0,1).$ Then the coin $(L,B,R)$ induced by
    $L=\begin{bmatrix}0&0\\ 0&-n\end{bmatrix},\qquad B=\begin{bmatrix}-\sqrt{b}&m\\ m&\sqrt{b}\end{bmatrix},\qquad R=\begin{bmatrix}n&0\\ 0&0\end{bmatrix}$
    is well-defined.
    The only invariant state for the auxiliary map is $\rho_{\infty}=I/2.$ Therefore,
    $\mathrm{Tr}(L\rho_{\infty}L^{*})=\frac{n^{2}}{2}=\mathrm{Tr}(R\rho_{\infty}R^{*})\Rightarrow\;(L,B,R)\mbox{ is recurrent,}$,
    by Item (1) of Theorem 3.9.
    It is worth noting that if we assume $m=0,$ then the walk is transient by Item (2) of Theorem 3.9. More explicitly, in this case, for the item (2.2), $|l_{1}|=0,\;|l_{2}|=n,\;|r_{1}|=n,\;|r_{2}|=0.$
    $\Diamond$
  2. 2.
    Let $x_{1},x_{2}\in\mathbb{C},$
    $|x_{1}|\leq\sqrt{\frac{2}{3}},\quad|x_{2}|\leq\sqrt{\frac{3}{4}},$
    then consider the coin $(L,B,R),$ where
    $L=\begin{bmatrix}\frac{1}{\sqrt{3}}&0\\ 0&\frac{1}{2}\end{bmatrix},\qquad B=\begin{bmatrix}\sqrt{\frac{2}{3}-x_{1}^{2}}&0\\ 0&\sqrt{\frac{3}{4}-x_{2}^{2}}\end{bmatrix},\qquad R=\begin{bmatrix}x_{1}&0\\ 0&x_{2}\end{bmatrix}.$
    We are in Item (2) of Theorem 3.9, thus we conclude that
    $(1)\quad|x_{1}|=\frac{1}{\sqrt{3}}\mbox{ and }|x_{2}|=\frac{1}{2}$$\Rightarrow$ $(L,B,R)\mbox{ is recurrent,}$,
    $(2)\quad|x_{1}|\neq\frac{1}{\sqrt{3}}\mbox{ and }|x_{2}|\neq\frac{1}{2}$$\Rightarrow$ $(L,B,R)\mbox{ is transient,}$,
    $(3)\quad|x_{1}|=\frac{1}{\sqrt{3}}\mbox{ and }|x_{2}|\neq\frac{1}{2}$$\Rightarrow$ $(L,B,R)\mbox{ is transient with respect to }\sigma=\begin{bmatrix}0&0\\ 0&1\end{bmatrix}\mbox{ and it is}$
    recurrent with respect to all densities but $\mbox{recurrent with respect to all densities but }\sigma,$
    $(4)\quad|x_{1}|\neq\frac{1}{\sqrt{3}}\mbox{ and }|x_{2}|=\frac{1}{2}$$\Rightarrow$ $(L,B,R)\mbox{ is transient with respect to }\sigma=\begin{bmatrix}1&0\\ 0&0\end{bmatrix}\mbox{ and it is}$
    recurrent with respect to all densities but $\mbox{recurrent with respect to all densities but }\sigma.$
$\Diamond$
Example 5.2.
Consider the Lazy HOQW induced by the following coin of dimension 3: $(L,B,R),$ where
$L=\frac{1}{8}\begin{bmatrix}2(1+\sqrt{2})&0&2(1-\sqrt{2})\\ 0&\sqrt{31}&0\\ 2(1-\sqrt{2})&0&2(1+\sqrt{2})\end{bmatrix},\quad B=\frac{1}{16}\begin{bmatrix}\sqrt{30}&2i&\sqrt{30}\\ 2&0&2\\ \sqrt{30}&-2i&\sqrt{30}\end{bmatrix},\quad R=\frac{1}{8}\begin{bmatrix}2(1-\sqrt{2})&0&2(1+\sqrt{2})\\ 0&\sqrt{31}&0\\ 2(1+\sqrt{2})&0&2(1-\sqrt{2})\end{bmatrix}.$
There is a unique invariant state $\rho_{\infty}$ given by
$\rho_{\infty}=\frac{1}{2}\begin{bmatrix}1&0&-1\\ 0&0&0\\ -1&0&1\end{bmatrix},$
and $\mathrm{Tr}\left(L\rho_{\infty}L^{*}\right)=1/2=\mathrm{Tr}\left(R\rho_{\infty}R^{*}\right).$ Therefore, this coin is recurrent by Theorem 3.7.
$\Diamond$
Since the invariant state also depends on the choice of $B,$ it is natural to ask if for a given transient coin $(L,B,R),$ there is another $B^{\prime}$ that makes $(L,B^{\prime},R)$ to be recurrent.
Example 5.3.
Firstly, let
$L=\frac{1}{7}\begin{bmatrix}1&0\\ 0&2\end{bmatrix},\qquad B=\frac{1}{7}\begin{bmatrix}-1&1\\ 2&0\end{bmatrix},\qquad R=\frac{1}{7}\begin{bmatrix}1&1\\ 0&1\end{bmatrix}.$
Then
$\rho_{\infty}=\frac{1}{3}\begin{bmatrix}1&0\\ 0&2\end{bmatrix},$
$\mathrm{Tr}\left(L\rho_{\infty}L^{*}\right)=3/7$ and $\mathrm{Tr}\left(R\rho_{\infty}R^{*}\right)=5/21.$ The coin $(L,B,R)$ is transient. However, if we take
$B^{\prime}=\frac{1}{\sqrt{35}}\begin{bmatrix}5&-1\\ 0&2\end{bmatrix},$
then the invariant state is
$\rho^{\prime}_{\infty}=\begin{bmatrix}1&0\\ 0&0\end{bmatrix},$
and $\mathrm{Tr}\left(L\rho_{\infty}L^{*}\right)=1/\sqrt{7}=\mathrm{Tr}\left(R\rho_{\infty}R^{*}\right).$ Thus the coin $(L,B^{\prime},R)$ is recurrent.
$\Diamond$
We remark that the choices of the coins above were taken in such a way that $L,B^{\prime}$ and $R$ had a common invariant eigenvector, namely $[1,0]^{T},$ which was not an eigenvector of $B.$ This is due to the hard job of, given non-trivial $R$ and $L,$ to find $B^{\prime}$ in such a way that $(L,B^{\prime},R)$ is recurrent. Attempt were conduct using mathematic software with many $L$’s and $R$’s having simple entries, but even though the entries of $B$’s that made $(L,B,R)$ to be transient were very long to write in this paper.
Example 5.4.
Let us consider the coin given in [6, Section 6.2]:
$R=\begin{bmatrix}\sqrt{\frac{3}{8}}&0&0&0\\ -\sqrt{\frac{p_{1}}{2}}&\frac{1}{\sqrt{2}}&0&0\\ -\sqrt{\frac{p_{2}}{2}}&0&\frac{1}{\sqrt{2}}&0\\ \sqrt{\frac{2p_{3}}{3}}&0&0&\frac{1}{\sqrt{3}}\end{bmatrix},\qquad L=\begin{bmatrix}\frac{1}{2\sqrt{2}}&0&0&0\\ \sqrt{\frac{p_{1}}{2}}&\frac{1}{\sqrt{2}}&0&0\\ \sqrt{\frac{p_{2}}{2}}&0&\frac{1}{\sqrt{2}}&0\\ -\sqrt{\frac{p_{3}}{3}}&0&0&\frac{\sqrt{2}}{\sqrt{3}}\end{bmatrix},$
where $p_{1},\;p_{2},p_{3}\geq 0$ and $p_{1}/2+p_{2}/2+p_{3}/2=1/2.$
The invariant states of the auxiliary map are of the form $\rho_{\infty}=\sigma_{1}+\sigma_{2},$ where
$\sigma_{1}=\omega_{1}\begin{bmatrix}0&0&0&0\\ 0&a&b&0\\ 0&b^{*}&1-a&0\\ 0&0&0&0\end{bmatrix},\quad\sigma_{2}=\omega_{2}\begin{bmatrix}0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&0&1\end{bmatrix},\quad\omega_{i},\omega_{j}\geq 0,\;\;\omega_{1}+\omega_{2}=1.$
We have
$m_{1}=\mathrm{Tr}(L\sigma_{1}L^{*})=\frac{1}{2},\qquad m_{2}=\mathrm{Tr}(L\sigma_{2}L^{*})=\frac{2}{3}\neq\frac{1}{2}.$
By item (2) of Theorem 4.3, the walk is transient for
$\rho=\begin{bmatrix}0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&0&1\end{bmatrix}$
and recurrent for all the remaining density operators on $\mathbb{C}^{4}$, for any appropriate choices of $p_{k},\;k=1,2,3.$
Example 5.5.
Consider a HOQW induced by the coin $(L,R),$ where
$L=\begin{bmatrix}\frac{\sqrt{5}}{5}&0&-\frac{\sqrt{5}}{5}\\ 0&\frac{2\sqrt{5}}{5}&\frac{\sqrt{5}}{10}\\ 0&0&\frac{1}{2}\end{bmatrix},\qquad R=\begin{bmatrix}\frac{2\sqrt{5}}{5}&0&\frac{\sqrt{5}}{10}\\ 0&\frac{\sqrt{5}}{5}&-\frac{\sqrt{5}}{5}\\ 0&0&\frac{1}{2}\end{bmatrix}.$
The invariant states of the auxiliary map are of the form $\rho_{\infty}=\sigma_{1}+\sigma_{2},$ where
$\sigma_{1}=\omega_{1}\begin{bmatrix}1&0&0\\ 0&0&0\\ 0&0&0\end{bmatrix},\quad\sigma_{2}=\omega_{2}\begin{bmatrix}0&0&0\\ 0&1&0\\ 0&0&0\end{bmatrix},\quad\omega_{i},\omega_{j}\geq 0,\;\;\omega_{1}+\omega_{2}=1.$
We have
$m_{1}=\mathrm{Tr}(L\sigma_{1}L^{*})=\frac{\sqrt{5}}{5}\neq\frac{1}{2},\qquad m_{2}=\mathrm{Tr}(L\sigma_{2}L^{*})=\frac{2\sqrt{5}}{5}\neq\frac{1}{2}.$
By item (1) of Theorem 4.3, the walk is transient.
Example 5.6.
Let us consider the coin $(L,R)$ of dimension 4 given by
$R=\begin{bmatrix}\frac{\sqrt{2}}{2}&-\frac{\sqrt{5}}{4}&0&\frac{1}{4}\\ 0&\frac{\sqrt{2}}{4}&0&0\\ 0&0&\frac{\sqrt{2}}{2}&0\\ 0&0&0&\frac{\sqrt{6}}{4}\end{bmatrix},\qquad L=\begin{bmatrix}\frac{\sqrt{2}}{2}&\frac{\sqrt{5}}{4}&0&-\frac{1}{4}\\ 0&\frac{1}{2}&0&\frac{\sqrt{5}}{4}\\ 0&0&-\frac{\sqrt{2}}{2}&0\\ 0&0&0&\frac{\sqrt{3}}{4}\end{bmatrix},$
The invariant states of the auxiliary map are of the form $\rho_{\infty}=\sigma_{1}+\sigma_{2},$ where
$\sigma_{1}=\omega_{1}\begin{bmatrix}1&0&0&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\end{bmatrix},\quad\sigma_{2}=\omega_{2}\begin{bmatrix}0&0&0&0\\ 0&0&0&0\\ 0&0&1&0\\ 0&0&0&0\end{bmatrix},\quad\omega_{i},\omega_{j}\geq 0,\;\;\omega_{1}+\omega_{2}=1.$
We have
$m_{1}=\mathrm{Tr}(L\sigma_{1}L^{*})=\frac{1}{2},\qquad m_{2}=\mathrm{Tr}(L\sigma_{2}L^{*})=\frac{1}{2}.$
By item (3) of Theorem 4.3, the walk is recurrent.
6 Conclusion
We present new general recurrence criteria for classes of HOQWs, whose dynamics are governed by Kraus operators that are induced by a finite-dimensional coin.
In the context of these HOQWs, we presented three distinct criteria for analyzing recurrence. Firstly, Theorem 3.7 introduces a criterion for ergodic Lazy coins, where the recurrence is determined by the operators $L$ and $R$ and by the only invariant state $\rho_{\infty}$ associated with the corresponding auxiliary map.
Next, Theorem 3.9 focuses on Lazy HOQWs of dimension 2, offering a complete and easy-to-apply criterion. The simplicity of this criterion comes from the fact that it depends on the knowledge of the eigenvectors of the operators $L$, $B$, and $R$, allowing a practical analysis of the recurrence of these processes that are not necessarily ergodic.
Finally, Theorem 4.3 extends the analysis to non-lazy coins, using a decomposition of the vector space where the operators $L$, $B$, and $R$ act. By checking recurrence in vector subspaces that contain unique invariant states, this general criterion is extended to any initial density, significantly expanding the scope of recurrence analysis for HOQWs in more complex configurations, since it is valid for any finite-dimensional coin.
[…]
Acknowledgements This work was supported by Universidade Federal do Pampa (UNIPAMPA).
[…]
Conflict of interest The author declare that there is no conflict of interest.
[…]
Data Availability The datasets generated during the current study are available from the corresponding author on reasonable request.