### v3

# Recurrence Criteria for Reducible Homogeneous Open Quantum Walks on the Line

Abstract
In this paper, we study the recurrence of Open Quantum Walks induced by finite-dimensional coins on the line ($\mathbb{Z}$) and on the grid ($\mathbb{Z}^{2}$). Two versions are considered: discrete-time open quantum walks (OQW) and continuous-time open quantum walks (CTOQW). We present three distinct recurrence criteria for OQWs on $\mathbb{Z}$, each adapted to different types of coins. The first criterion applies to coins whose auxiliary map has a unique invariant state, resulting in the first recurrence criterion for Lazy OQWs. The second one applies to Lazy OQWs of dimension 2, where we provide a complete characterization of the recurrence for this low-dimensional case. The third one presents a general criterion for finite-dimensional coins in the non-lazy case, which generalizes many of the previously known results for OQWs on $\mathbb{Z}$. Also, we present a general recurrence criterion for OQWs on $\mathbb{Z}^{2}$ via the open quantum jump chain, obtained from a recurrence criterion for CTOQWs on $\mathbb{Z}^{2}$.
keywords:
Homogeneous Open Quantum Walks, Recurrence, Lazy Walks, Coin, Reducibility
1 Introduction
Open Quantum Walks (OQWs) naturally extend classical Markov chains within the framework of open quantum systems. Originally proposed by [2], OQWs describe open quantum systems in which repeated measurements of the particle’s position, together with the particle’s state space, give rise to a Markov chain.
In this context, we first consider homogeneous Open Quantum Walks (HOQWs) defined on the set of integers, that is, the graph of this type of walk has the set of sites $V=\mathbb{Z}$, where three linear operators $L,B,R$ correspond to transitions between sites in the quantum walk, remaining constant, with each step moving only to the nearest neighbors $(R,L)$ or staying at the same site $(B)$.
Recurrence is an important probabilistic property of any walk, often characterized by the expected number of returns to the initial site, bringing a central concern in the quantum context, and can be directly related to classical recurrence criteria for walks in $\mathbb{Z}$. For example, consider the classical homogeneous walk on $\mathbb{Z}$ with graph as shown in Figure 1.
It is well known that this lazy random walk with $r,l\geq 0,\;b>0,\;l+b+r=1,$ is recurrent if and only if the transition parameters $r$ (rate of jumping to the right) and $l$ (rate of jumping to the left) are equal. Also, if $b=0,$ then the walk is non-lazy and it is recurrent if and only if $r=1/2=l$ [21].
[…]
Figure 1: Classical Homogeneous Lazy Random Walk on $\mathbb{Z}$.
In the context of discrete-time Markov chains, a state is called recurrent if, once the chain enters that state, the probability of returning to it is 1. This is equivalent to saying that the average number of returns to that state is infinite, indicating that the chain will revisit the state indefinitely [21].
With this description, it is natural to seek criteria to determine under what conditions an HOQW is recurrent, given $L,B,R.$ This problem has been addressed in the literature with the development of specific criteria, which are essential to understanding the dynamics and long-term behavior of these systems. See, for instance, [12, 17, 16]. These works define recurrence in terms of probability of return, where a state is said to be recurrent if the probability of the walk returning to it is equal to 1. However, it has been shown that for OQWs, the two definitions of recurrence based on returning with probability 1 are not equivalent to the one based on the average number of returns being infinite [4]. This property highlights a distinct characteristic of OQWs compared to classical Markov chains.
Once this difference is understood, we discuss the recurrence criteria for HOQWs, adopting the definition proposed in [4]. Specifically, we analyze the recurrence of HOQWs based on the average number of returns.
Furthermore, recurrence criteria concerning the mean number of returns for higher-order HOQWs have been presented in [15]. The authors of that work extend the analysis of recurrence to a more general class of open quantum walks, such as irreducible coins.
In this article, we present more general criteria for the recurrence of lazy and non-lazy HOQWs, based on the transition operators $L,B,R$ when acting on finite-dimensional vector spaces, that is, we allow the HOQW to be reducible. Furthermore, we present the first recurrence criterion for lazy HOQWs, using results obtained from recent studies on Central Limit Theorems for this kind of HOQWs.
A first Central Limit Theorem (CLT) for Open Quantum Walks (OQWs) was established in [1]. In this work, we rely on more recent developments, including the CLT for Lazy HOQWs presented in [14], as well as the general CLT for HOQWs on lattices with finite internal degrees of freedom given in [10].
By making use of the ideas above and generalizing results from the classical version to recurrence, we also obtain a recurrence criterion for the HOQW on the grid, which is the case where the set of vertices is $\mathbb{Z}^{2}$; however, this is achieved through continuous-time OQWs (CTOQWs). For this purpose, we use recent results by [20], where the introduction of the open quantum jump chain allowed for the association between recurrence in discrete-time and continuous-time versions of OQWs.
This work is structured as follows. In Section 2, we define the open quantum walks induced by a coin and their transition probabilities. Section 3 is devoted to presenting a recurrence criterion for a class of Lazy HOQWs, and one restricted to Lazy HOQWs of dimension 2. In Section 4, we focus on the case where the coin is non-lazy, and we present a general recurrence criterion for finite-dimensional coins. Section 5 illustrates a first pack of comprehensive examples. Section 6 discusses OQWs on the grid, and Section 7 brings examples of HOQWs on those walks.
2 General Settings
Let us consider the set of vertices $\mathbb{Z},$ and $\mathcal{K}=\mathbb{C}^{\mathbb{Z}}$ be the state space of a quantum system with as many degrees of freedom as the number of vertices. To emphasize the position of the particle of our quantum walk, we consider a quantum system described by the separable complex Hilbert space
$\mathcal{H}=\bigoplus_{i\in\mathbb{Z}}\mathfrak{h}_{i},$
where the $\mathfrak{h}_{i}$ are separable Hilbert spaces. The label $i\in\mathbb{Z}$ represents the position of the particle and, when the particle is located at the vertex $i\in\mathbb{Z}$, its internal state is encoded in the space $\mathfrak{h}_{i},$ as it will be detailed below.
We consider the set of density operators over a finite Hilbert space $\mathcal{W},$
$\mathcal{D}(\mathcal{W})=\{\rho:\mathcal{W}\rightarrow\mathcal{W}:\;\rho\geq 0,\;\mathrm{Tr}(\rho)=1\},$
and introduce
$\mathcal{S}(\mathcal{H}\otimes\mathcal{K})=\left\{\tau\in\mathcal{I}(\mathcal{H}\otimes\mathcal{K}):\textstyle\tau=\displaystyle\sum_{j\in\mathbb{Z}}\rho(j)\otimes\lvert j\rangle\langle j\rvert,\quad\rho(j)\geq 0,\quad\sum_{j\in\mathbb{Z}}\mathrm{Tr}\left(\rho(j)\right)=1\right\},$
the set of density operators acting on $\mathcal{H}\otimes\mathcal{K}.$ We are interested in HOQWs of the form
$\Phi(\tau)=\sum_{i\in\mathbb{Z}}\Big(L\rho(i+1)L^{*}+B\rho(i)B^{*}+R\rho(i-1)R^{*}\Big)\otimes|i\rangle\langle i|,\;\;\;\tau\in\mathcal{S}(\mathcal{H}\otimes\mathcal{K}),$ (1)
where $L,B,R\in\mathbb{M}_{d}({\mathbb{C}})$ satisfy $L^{*}L+B^{*}B+R^{*}R=I_{d}.$
We say that the triple of matrices $(L,B,R)$ is a coin if eq. (1) holds. Therefore, we can assume that $\mathfrak{h}_{i}=\mathfrak{h}\;\forall i\in\mathbb{Z},$ and if $\mathfrak{h}=\mathbb{C}^{d}$ for some $d$, the dimension of the coin is said to be $d$. In this article, we consider only the case $d<\infty.$
If $B\neq 0,$ the quantum walk exhibits a lazy behavior, meaning that in the context of quantum walks, there is a non-zero probability for the walker to stay in the same vertex after a step. We refer to the HOQW above as a HOQW induced by the coin $(L,B,R)$ in which the walk is inherently lazy.
Therefore, this coin represents a homogeneous, nearest-neighbor OQW on the integer line with the left and right transitions given by $L$ and $R$, respectively, while $B$ describes the rate of no jump at some instant.
This lazy behavior introduces distinct properties compared to its non-lazy counterpart. Notably, the presence of the matrix $B$ influences the walk’s propensity to remain in its current vertex, thereby affecting the overall dynamics. If $B=0,$ then $\Phi$ will be called a HOQW induced by a coin $(L,R),$ or a non-lazy HOQW induced by a coin $(L,R).$ The dynamic described above is represented by the graph in Figure 2.
[…]
Figure 2: Lazy Quantum Walk on $\mathbb{Z}$.
The quantum trajectory of this kind of HOQW on the line, induced by a coin $(L,B,R)$ and starting from a state $\tau$ of the form $\tau=\sum_{i\in\mathbb{Z}}\tau_{i}\otimes|i\rangle\langle i|,$ is any path generated by the Markov chain $(X_{n},\tau_{n})_{n\geq 0}$, where $X_{n}$ denotes the position of the particle at time $n$ and $\tau_{n}$ its internal degree. The transition probabilities are given by
$\begin{split}\mathbb{P}\left(\,(X_{n+1},\tau_{n+1})=\left(i-1,\frac{L\sigma L^{*}}{\mathrm{Tr}(L\sigma L^{*})}\right)\;\Bigg|\;(x_{n},\tau_{n})=(i,\sigma)\,\right)&=\mathrm{Tr}(L\sigma L^{*}),\\ \mathbb{P}\left(\,(X_{n+1},\tau_{n+1})=\left(i,\frac{B\sigma B^{*}}{\mathrm{Tr}(B\sigma B^{*})}\right)\;\Bigg|\;(x_{n},\tau_{n})=(i,\sigma)\,\right)&=\mathrm{Tr}(B\sigma B^{*}),\\ \mathbb{P}\left(\,(X_{n+1},\tau_{n+1})=\left(i+1,\frac{R\sigma R^{*}}{\mathrm{Tr}(R\sigma R^{*})}\right)\;\Bigg|\;(x_{n},\tau_{n})=(i,\sigma)\,\right)&=\mathrm{Tr}(R\sigma R^{*}),\end{split}$
for every $i\in\mathbb{Z}$, $\sigma\in\mathcal{D}(\mathfrak{h})$, and initial law
$\mathbb{P}\left(\,(X_{0},\tau_{0})=\left(i,\frac{\tau_{i}}{\mathrm{Tr}\,\tau_{i}}\right)\;\right)=\mathrm{Tr}(\tau_{i}),$
and all other transition probabilities are null. Therefore, at time $n$, if the system is located at site $\lvert i\rangle$ with an initial density $\rho\otimes\lvert i\rangle\langle i\rvert$, then at the next time step $n+1$, one of the following occurs:
  • The particle moves to site $\lvert i-1\rangle$ (left) with probability $\operatorname{Tr}(L_{i}\rho L_{i}^{*})$, and the density evolves to
    $\frac{L_{i}\rho L_{i}^{*}}{\operatorname{Tr}(L_{i}\rho L_{i}^{*})}\otimes\lvert i-1\rangle\langle i-1\rvert$;
  • The particle remains at site $\lvert i\rangle$ with probability $\operatorname{Tr}(B_{i}\rho B_{i}^{*})$, and the new density matrix is
    $\frac{B_{i}\rho B_{i}^{*}}{\operatorname{Tr}(B_{i}\rho B_{i}^{*})}\otimes\lvert i\rangle\langle i\rvert$;
  • The particle moves to site $\lvert i+1\rangle$ (right) with probability $\operatorname{Tr}(R_{i}\rho R_{i}^{*})$, leading to the new density matrix
    $\frac{R_{i}\rho R_{i}^{*}}{\operatorname{Tr}(R_{i}\rho R_{i}^{*})}\otimes\lvert i+1\rangle\langle i+1\rvert$.
3 Recurrence criteria for Lazy Open Quantum Walks
3.1 Faithfulness of Density Operators
Density operators play a fundamental role in open quantum systems. In this work, they provide statistical descriptions of HOQWs. Faithful and non-faithful density operators are characterized by their positive definiteness and positive semi-definiteness, respectively, when they evolve in finite-dimensional systems.
Faithful density operators are defined as being positive definite. This means that the eigenvalues of a faithful density operator are strictly positive, and the operator is invertible. Essentially, faithfulness ensures a complete representation of the quantum system, capturing the full richness of its physical properties.
On the other hand, non-faithful density operators are positive semi-definite, thus they have $\lambda=0$ as one of their eigenvalues. The presence of zero eigenvalues implies that the non-faithful density operator is not invertible and may represent specific mixed states or situations where information loss occurs. The positive semi-definiteness criterion relaxes the strictness of positive definiteness, allowing for a broader range of quantum state representations, while still maintaining the one-to-one correspondence between quantum states and density operators.
An important subset of the set of non-faithful density operators is composed of the pure density operators. A pure density is a density $\rho$ that is a projection of rank 1, i.e., there exists some vector $\lvert u\rangle\in\mathbb{C}^{d}$ such that $\rho=\lvert u\rangle\langle u\rvert$.
Let us consider an HOQW starting the walk on vertex $\lvert i\rangle$ with initial density operator $\rho.$ For evolution on time $n=0,1,2,\ldots,$ we define
[…]
where
$\Phi^{n}(\rho\otimes\lvert i\rangle\langle i\rvert)=\sum_{s\in\mathbb{Z}}\rho_{n}(s)\otimes\lvert s\rangle\langle s\rvert$.
With this notation,
[…]
denote the probability of being at site $\lvert j\rangle$ at step $n,$ given that we started at site $\lvert i\rangle,$ with initial density $\rho$ concentrated at $\lvert i\rangle.$ Therefore, the dynamics starts with a density operator $\rho$ concentrated at some vertex $\lvert i\rangle,$ takes the evolution up to time ($n$) through the linear operator $\Phi^{n},$ producing a new density operator
$\rho_{n}=\sum_{s}\rho_{n}(s)\otimes\lvert s\rangle\langle s\rvert=\Phi^{n}(\rho\otimes\lvert i\rangle\langle i\rvert),\quad\mathrm{Tr}\left(\sum_{s}\rho_{n}(s)\right)=1,$
and then we project $\rho_{n}$ onto the subspace generated by vertex $\lvert j\rangle.$ Note that
$\Phi^{n}(\rho\otimes\lvert i\rangle\langle i\rvert)(I\otimes\lvert j\rangle\langle j\rvert)=\rho_{n}(j)\otimes\lvert j\rangle\langle j\rvert$,
which represents the data concentrated at vertex $\lvert j\rangle$ at time $n.$
Consider an HOQW on some set of vertices $\mathbb{Z}$, let $i\in\mathbb{Z}$ and $\rho\in\mathcal{D}(\mathfrak{h})$. We say that a vertex $\lvert i\rangle$ is
  1. 1.
    $\rho$-recurrent if
    $\sum_{n=0}^{\infty}p_{ii;\rho}(n)=\infty.$ (2)
    Otherwise, $\lvert i\rangle$ is said to be $\rho$-transient;
  2. 2.
    recurrent, if $\lvert i\rangle$ is $\rho$-recurrent for all $\rho\in\mathcal{D}(\mathfrak{h});$
  3. 3.
    transient, if $\lvert i\rangle$ is $\rho$-transient for all $\rho\in\mathcal{D}(\mathfrak{h}).$
Concerning all vertices, an OQW is called: recurrent if every vertex is recurrent; transient if every vertex is transient. Note that equation (2) computes the mean number of returns to site $\lvert i\rangle.$
The following Proposition describes $\rho$-recurrence in a concentrated vertex $i\in\mathbb{Z}$ when assumed that $\lvert i\rangle$ is recurrent for some density and $\mathfrak{h}$ has a finite internal degree of freedom. An immediate consequence will also be given for the case where $\lvert i\rangle$ is $\sigma$-transient for some non-faithful density.
Theorem 3.1.
Consider an OQW on some set of vertices $V,$ let $i\in V$ and assume that $\mathfrak{h}$ is finite-dimensional. If $\lvert i\rangle$ is $\tau$-recurrent for some $\tau\in\mathcal{D}(\mathfrak{h}),$ then
  1. 1.
    vertex $\lvert i\rangle$ is $\rho$-recurrent for all faithful $\rho\in\mathcal{D}(\mathfrak{h})$;
  2. 2.
    there exists a pure density operator $\lvert w\rangle\langle w\rvert$ such that the OQW is $\lvert w\rangle\langle w\rvert-$recurrent, where $\lvert w\rangle$ is an unit vector of $\mathfrak{h}.$
Proof.
Let $\rho\in\mathcal{D}(\mathfrak{h})$ be faithful. Since the state space $\mathcal{D}(\mathfrak{h})$ is compact in the trace-norm topology, the function
$f(\tau)=\inf\{c>0\mid\rho>c\tau\}$
is continuous and thus attains its minimum and maximum. In particular, there exists $c>0$ such that
$\rho>c\tau,\quad\forall\rho\in\mathcal{D}(\mathfrak{h}).$
thus
$\begin{split}\sum_{n=0}^{\infty}p_{ii;\rho}(n)=&\sum_{n=0}^{\infty}\textmd{Tr}\left[\Phi^{n}\left(\rho\otimes\lvert i\rangle\langle i\rvert\right)\right(I\otimes\lvert i\rangle\langle i\rvert)]\\ >&c\sum_{n=0}^{\infty}\textmd{Tr}\left[\Phi^{n}\left(\tau\otimes\lvert i\rangle\langle i\rvert\right)(I\otimes\lvert i\rangle\langle i\rvert)\right]\\ =&c\sum_{n=0}^{\infty}p_{ii;\tau}(n)\\ =&\infty,\end{split}$
giving the first item.
For the second one, let
$\tau=\sum_{j}\lambda_{j}\lvert e_{j}\rangle\langle e_{j}\rvert$,
where each $\lambda_{j}>0$ is an eigenvalue of $\tau$ with corresponding eigenvector $\lvert e_{j}\rangle.$ Then
[…]
where at least one of the members of the sum must be infinity. This completes the proof.
$\square$
By a contra-position argument, we obtain the following direct consequence of Theorem 3.1.
Corollary 3.2.
Let $\Phi$ be an OQW on some set of vertices $V,$ $i\in V,$ and $\dim(\mathfrak{h})<\infty.$ If $\lvert i\rangle$ is $\tau$-transient for some faithful $\tau\in\mathcal{D}(\mathfrak{h}),$ then $\lvert i\rangle$ is transient.
Another direct consequence of Theorem 3.1 is related to the homogeneous case.
Corollary 3.3.
Let $\Phi$ be an HOQW on $\mathbb{Z}$ with $\dim(\mathfrak{h})=d<\infty.$ If $\Phi$ is $\tau$-recurrent for some $\tau\in\mathcal{D}(\mathfrak{h}),$ then
  1. 1.
    $\Phi$ is $\rho$-recurrent for all faithful $\rho\in\mathcal{D}(\mathfrak{h})$;
  2. 2.
    there exists a pure density operator $\lvert w\rangle\langle w\rvert$ such that $\Phi$ is $\lvert w\rangle\langle w\rvert-$recurrent, where $\lvert w\rangle$ is an unit vector of $\mathbb{C}^{d}.$
3.2 Enclosures and Invariant States
Firstly, a subspace $\mathcal{Y}$ of $\mathfrak{h}$ is called an enclosure for a quantum channel $\mathcal{L}$ if for every $\sigma\in\mathcal{D}\mathfrak{h}),$ we have
$\mathrm{supp}(\sigma)\subset\mathcal{Y}\Rightarrow\mathrm{supp}\left(\mathcal{L}(\sigma)\right)\subset\mathcal{Y}.$
If an enclosure contains no other non-trivial enclosures, then it is called a minimal enclosure.
A minimal enclosure with finite dimension is the support of a unique invariant state of the channel, that is, a density operator $\sigma$ such that $\mathcal{L}(\sigma)=\sigma$. For more on enclosures of HOQWs, we refer the reader to [10, 11, 4].
Following this property and using the homogeneity of HOQWs induced by a coin, we split $\mathfrak{h}$ into two subsets $\mathcal{Y}$ and $\mathcal{X}$ in such a way that the walk is transient for $\mathcal{D}(\mathcal{Y})$ and recurrent everywhere else, with $\mathcal{Y}$ being an enclosure of the quantum channel with Kraus operators $L,B,R.$
Remark 3.4.
From now on we disregard the trivial cases where at least one of the operators $L,B,R$ has some eigenvalue $\lambda$ of modulus 1. To clarify, let $T\lvert v\rangle=\lambda\lvert v\rangle,$ where $\lvert v\rangle$ is an unitary vector and $T\in\{L,B,R\}.$ In this case, starting the walk with density operator $\rho=\lvert v\rangle\langle v\rvert$, the walker is either absorbed by a vertex $(T=B)$ or the particle always jumps to the same side $(T=L$ or $T=R).$
Proposition 3.5.
Consider an HOQW induced by a coin $(L,B,R)$ of dimension $d$. We are in one (and only one) of the following situations:
  • $(L,B,R)$ is recurrent;
  • $(L,B,R)$ is transient;
  • $(L,B,R)$ is recurrent with respect to all densities except for a convex set $\mathcal{D}(\mathcal{Y}).$ In this case, $\mathcal{Y}$ is an enclosure of $\mathcal{L},$
    $\mathfrak{h}=\mathcal{Y}\oplus\mathcal{X}$ (3)
    for some subspace $\mathcal{X}$ of $\mathfrak{h},$ with $\mathcal{Y},\mathcal{X}\neq\{0\}.$
Proof.
We claim $(L,B,R)$ is $\rho$-transient if and only if it is $\rho_{n}$-transient for every $n,$ where $\rho_{n}$ is the density at the $n$-th step of the walk.
Indeed, let us suppose the walk starts at vertex $\lvert 0\rangle$ with initial density $\rho.$ For instance, if the first jump is to $\lvert-1\rangle,$ the density, after this first jump, is
$\rho_{1}=\frac{L\rho L^{*}}{\mathrm{Tr}(L\rho L^{*})}.$
Since the walk is homogeneous, it is $\rho$-transient if and only if it is $\rho_{1}$-transient. After $n$ jumps, the density is
$\rho_{n}=\frac{J_{\pi}\rho J_{\pi}^{*}}{\mathrm{Tr}(J_{\pi}\rho J_{\pi}^{*})},$ (4)
where $J_{\pi}$ is a concatenation of $n$ operators, each one belonging to $\{L,B,R\},$ proving the claim, once this is valid for any $n=1,2,3,\ldots$
Moreover, by equation (4) we have that $\cup_{n}\;\text{supp}(\rho_{n})$ is a subspace of $\mathfrak{h}.$ Joining all the possible initial states in which $(L,B,R)$ is transient, we obtain a decomposition (3), where $\mathcal{Y}$ or $\mathcal{X}$ may be $\{0\}$, and by definition $\mathcal{Y}$ is an enclosure.
Let us suppose that $(L,B,R)$ is neither recurrent nor transient. By Theorem 3.1, there exist unit vectors $\lvert u\rangle,\lvert v\rangle\in\mathfrak{h}$ such that the walk is $\lvert u\rangle\langle u\rvert-$recurrent and $\lvert v\rangle\langle v\rvert-$transient. Let $\mathcal{Y}$ be the subspace of $\mathfrak{h}$ such that $\mathcal{D}(\mathcal{Y})$ is the set of densities of $\mathcal{B}(\mathfrak{h})$ in which $(L,B,R)$ is transient, thus we pick $\rho\in\mathcal{D}(\mathcal{Y})$ and we can affirm that $(L,B,R)$ is $\rho_{n}$-transient for every $n.$ Finally, take $\mathcal{X}=\mathcal{Y}^{\perp}$ and the trichotomy holds.
$\square$
In our subsequent discussion, we consider a coin $(L,B,R)$ and its auxiliary map $\mathcal{L}:\mathcal{B}(\mathfrak{h})\rightarrow\mathcal{B}(\mathfrak{h})$ given by
$\mathcal{L}(\rho)=L\rho L^{*}+B\rho B^{*}+R\rho R^{*},\quad\rho\in\mathcal{B}(\mathfrak{h}).$
Note that $\mathcal{L}$ is a quantum channel, thus it has at least one invariant state, that is, there exists some density operator $\rho_{\infty}\in\mathcal{B}(\mathfrak{h})$ satisfying $\mathcal{L}(\rho_{\infty})=\rho_{\infty}.$ When the invariant state of $\mathcal{L}$ is unique, we will say that the map $\mathcal{L}$ is ergodic.
We will use the following result obtained straightforwardly from [14].
Theorem 3.6.
Consider a Lazy HOQW induced by a coin $(L,B,R)$ of dimension $d.$ If the auxiliary map $\mathcal{L}$ admits only one invariant state $\rho_{\infty},$ then
$\frac{X_{n}}{n}\stackrel{{\scriptstyle n\rightarrow\infty}}{{\longrightarrow}}m,\qquad\text{where }\qquad m=\mathrm{Tr}\left(R\rho_{\infty}R^{*}\right)-\mathrm{Tr}\left(L\rho_{\infty}L^{*}\right),$
in probability.
Proof.
This result can be obtained by using a Doob decomposition as described in [14, Equation 27] and [1, Theorem 5.2] for a Martingale $M_{n}$ satisfying
$X_{n}-n.m=X_{0}+M_{n}+\beta_{n},$
where $M_{n}/n\rightarrow 0$ and $\beta_{n}\rightarrow 0$ when $n\rightarrow 0.$
$\square$
Theorem 3.7 (Chung-Fuchs Theorem for HOQWs with ergodic auxiliary map).
Consider an HOQW induced by a coin $(L,B,R),\;\mathcal{L}$ its auxiliary map and $\mathcal{Y}$ an enclosure. Suppose that the initial density operator is $\rho\in\mathcal{D}(\mathcal{Y})$ with quantum trajectories $(X_{n},\rho_{n})_{n\geq 0}$ and $X_{n}/n\rightarrow 0$ in probability, then $(L,B,R)$ is $\sigma$-recurrent for some $\sigma\in\mathcal{D}(\mathcal{Y}).$
Proof.
The proof is analogous to [15, Theorem 7] and we will proceed by remarking on how it can be generalized. Firstly, the maximum appearing in [15, Lemma 6] can be taken in $\mathcal{D}=\mathcal{D}(\mathcal{Y}).$ Indeed, if the initial state is $\rho\in\mathcal{D}(\mathcal{Y}),$ then $\rho_{n}\in\mathcal{D}(\mathcal{Y})$ for every $n,$ thereby the maximum taken in the proof of [15, Lemma 6] can be reduced to $\mathcal{D}(\mathcal{Y}).$ Therefore, by the same reason, the density operator considered in [15, Theorem 7] belongs to $\mathcal{D}(\mathcal{Y}),$ proving the Theorem for the non-lazy case. Moreover, the lazy case holds as an extension of this proof, since all the results work for HOQWs on $\mathbb{Z}$.
$\square$
The following theorem presents the first recurrence criterion for non-irreducible HOQWs of arbitrary finite dimension. This result extends [15, Corollary 8 and Proposition 8] for reducible CTOQWs and can also be applied to Lazy HOQWs.
Theorem 3.8 (Recurrence Criteria).
Let us consider an HOQW on $\mathbb{Z}$ induced by a coin $(L,B,R)$ of dimension $d$ such that the auxiliary map $\mathcal{L}$ is ergodic. Then $\mathcal{L}$ has a unique invariant state $\rho_{\infty}$ and
  • $\textmd{Tr}\left(L\rho_{\infty}L^{*}\right)=\textmd{Tr}\left(R\rho_{\infty}R^{*}\right)\Rightarrow$ the walk is recurrent;
  • $\textmd{Tr}\left(L\rho_{\infty}L^{*}\right)\neq\textmd{Tr}\left(R\rho_{\infty}R^{*}\right)\Rightarrow$ the walk is transient.
Proof.
We suppose that the walk is neither recurrent nor transient. In this case, the walk is recurrent for some density and transient for all densities in some $\mathcal{D}(\mathcal{Y}),$ where $\mathcal{Y}$ is an enclosure (Proposition 3.5). Moreover, the unique invariant state $\rho_{\infty}$ of $\mathcal{L}$ must belong to $\mathcal{D}(\mathcal{Y}).$
If $\textmd{Tr}\left(L\rho_{\infty}L^{*}\right)=\textmd{Tr}\left(R\rho_{\infty}R^{*}\right)$ then $X_{n}/n\rightarrow 0$ by Theorem 3.6, thus $(L,B,R)$ is $\rho$-recurrent for some $\rho\in\mathcal{D}(\mathcal{Y})$ by Theorem 3.7, which is a contradiction.
On the other hand, if $\textmd{Tr}\left(L\rho_{\infty}L^{*}\right)\neq\textmd{Tr}\left(R\rho_{\infty}R^{*}\right),$ then $X_{n}/n\rightarrow m\neq 0.$ Therefore $\mathbb{P}_{0,\rho}(X_{n}=0\mbox{ i.o.})=0,$ showing that the mean number of returns to the initial site is finite for any $\rho$, thus the walk is transient, which is also a contradiction. We have shown that either the walk is recurrent or transient.
Therefore, if $\textmd{Tr}\left(L\rho_{\infty}L^{*}\right)=\textmd{Tr}\left(R\rho_{\infty}R^{*}\right),$ then the walk is $\rho$-recurrent for some density and thus it is recurrent. If $\textmd{Tr}\left(L\rho_{\infty}L^{*}\right)\neq\textmd{Tr}\left(R\rho_{\infty}R^{*}\right),$ then the walks is transient for every density.
$\square$
We remark that the operator $B$ does influence in the recurrence of $(L,B,R).$ Indeed, even it does not appear on the bullet points that describe the recurrence criterion, the invariant state $\rho_{\infty}$ may change when we pick distinct $B$’s. See example 5.3 below.
3.3 General Recurrence Criteria for HOQWs induced by a coin $(L,B,R)$ of dimension 2
The following theorem is due to [8].
Theorem 3.9.
For a finite quantum channel $\mathcal{L}:\mathcal{B}(\mathfrak{h})\rightarrow\mathcal{B}(\mathfrak{h})$, the following are equivalent:
i.
$\mathcal{L}$ is ergodic;
ii.
$\mathcal{L}$ does not have two invariant subspaces $S_{1}\neq\{0\}$ and $S_{2}\neq\{0\}$ such that $S_{1}\cap S_{2}=\{0\};$
iii.
$\mathcal{L}$ has a minimal invariant subspace, that is, a subspace $S\neq\{0\}$ such that $S\subset S^{\prime}$ for every invariant subspace $S^{\prime}\neq\{0\}.$
Building upon the established preceding proposition, we will advance to the following theorem, delineating a comprehensive criterion for 2-dimensional coins, extending [15, Theorem 17] to encompass the Lazy case. This criterion will be structured into two distinct parts: the first adhering to the logic of the previous theorem, and the second articulated in terms of the eigenvalues of the operators of the coin, independent of the invariant densities of the auxiliary map.
Theorem 3.10.
Consider an HOQW on $\mathbb{Z}$ induced by a coin $(L,B,R)$ of dimension 2.
(1) If $L,B,R$ have at most one common eigenvector, let $\rho_{\infty}$ be the unique invariant density of the auxiliary map. Then, we have
$(1.1)\quad\textmd{Tr}(L\rho_{\infty}L^{*})\neq\textmd{Tr}(R\rho_{\infty}R^{*})\Rightarrow(L,B,R)\mbox{ is transient},$
$(1.2)\quad\textmd{Tr}(L\rho_{\infty}L^{*})=\textmd{Tr}(R\rho_{\infty}R^{*})\Rightarrow(L,B,R)\mbox{ is recurrent}.$
(2) If $L,B,R$ have two linearly independent eigenvectors in common, $\lvert u_{1}\rangle,\lvert u_{2}\rangle$ unit, then they satisfy $\lvert u_{1}\rangle\bot\lvert u_{2}\rangle,$ we can put
$L\lvert u_{j}\rangle=l_{j}\lvert u_{j}\rangle,\qquad B\lvert u_{j}\rangle=b_{j}\lvert u_{j}\rangle,\qquad R\lvert u_{j}\rangle=r_{j}\lvert u_{j}\rangle,\qquad j=1,2,$
and
$(2.1)\quad|l_{1}|=|r_{1}|\mbox{ and }|l_{2}|=|r_{2}|$ $\Rightarrow$ $(L,B,R)\mbox{ is recurrent,}$
$(2.2)\quad|l_{1}|\neq|r_{1}|\mbox{ and }|l_{2}|\neq|r_{2}|$ $\Rightarrow$ $(L,B,R)\mbox{ is transient,}$
[…]
Proof.
If the elements of the coin do not share at least two orthogonal eigenvectors, then the auxiliary map is ergodic, thus item (1) is a particular case of Theorem 3.8.
Now we suppose that $L,B,R$ share two linearly independent eigenvectors, then we can apply Theorem 3.9 to obtain $\lvert u_{1}\rangle,\lvert u_{2}\rangle$ common orthonormal eigenvectors of $L,B,R.$ We assume that $|l_{1}|=|r_{1}|$ and $|l_{2}|=|r_{2}|,$ then for $\rho=\lvert u\rangle\langle u\rvert$, where $\lvert u\rangle$ is some common eigenvector of $L,B,R,$ we have $p_{j,j+1;\rho}(n)=p_{j,j-1;\rho}(n)=w$ for all $j\in\mathbb{Z}$ and all $n\geq 0$ for some constant $w$ and therefore, in this case, we have a classical homogeneous walk which is recurrent. If $(L,B,R)$ was not transient, then we would be in third situation of proposition 3.5, with
$\mathfrak{h}=\mathcal{Y}\oplus\mathcal{X},\quad\text{where}\quad\text{dim}(\mathcal{Y})\geq 1,\;\;\text{ and }\;\;\text{dim}(\mathcal{X})=2,$
a contradiction. Therefore, we are in the first situation of proposition 3.5 and the walk is recurrent.
Analogously, if $|l_{1}|\neq|r_{1}|$ and $|l_{2}|\neq|r_{2}|,$ then for $\rho_{1}=\lvert u_{1}\rangle\langle u_{1}\rvert$, and $\rho_{2}=\lvert u_{2}\rangle\langle u_{2}\rvert$, where $\lvert u_{1}\rangle,\lvert u_{2}\rangle$ are orthonormal common eigenvectors of $L,B,R,$ we have $p_{j,j+1;\rho}(n)<p_{j,j-1;\rho}(n)$ (or >>) for all $j\in\mathbb{Z}$ and all $n\geq 0$ and therefore in this case we have a classical homogeneous walk which is transient for both $\rho_{1},\rho_{2}$. Since the walk is transient for two distinct eigenvectors, the walk is transient by corollary 3.2.
For the last case, $|l_{i}|=|r_{i}|$ and $|l_{j}|\neq|r_{j}|,$ $(L,B,R)$ is $\rho_{i}$-recurrent and $\rho_{j}$-transient, where $\rho_{i}=\lvert u_{i}\rangle\langle u_{i}\rvert$ and $\rho_{j}=\lvert u_{j}\rangle\langle u_{j}\rvert$. This situation is a particular case of the last item in Proposition 3.5.
$\square$
4 A General Recurrence Criteria for Open Quantum Walks
Following the approach described in [10, Section 4], as well as the references therein, we can obtain for a coin $(L,R)$ a decomposition
[…]
Since $\mathfrak{h}$ is finite-dimensional, any minimal enclosure is included in $\mathfrak{R},$ and is the support of a unique extremal invariant state. Moreover, we have a unique decomposition of $\mathfrak{R}$ of the form
$\mathfrak{R}=\bigoplus_{\alpha\in A}\mathfrak{X}_{\alpha},$
where $\{\mathfrak{X}_{\alpha}\}_{\alpha}\in A$ is a finite set of mutually orthogonal enclosures and every $\mathfrak{X}_{\alpha}$ is minimal in the set of enclosures satisfying
[…]
Further, every $\mathfrak{X}_{\alpha}$ either is a minimal enclosure or assumes a decomposition (not unique) as the sum of mutually orthogonal isomorphic minimal enclosures $\{\mathcal{Y}_{\alpha,\beta},\;\beta\in I_{\alpha}\}:$
$\mathfrak{X}_{\alpha}=\bigoplus_{\beta\in I_{\alpha}}\mathcal{Y}_{\alpha,\beta}.$
Given a minimal enclosure $\mathcal{Y}_{\beta},$ let $\tau_{\beta}$ be the unique invariant state with support on $\mathcal{Y}_{\beta}.$ Introducing the parameter
$m_{\beta}=\mathrm{Tr}\left(L\tau_{\beta}L^{*}\right)-\mathrm{Tr}\left(R\tau_{\beta}R^{*}\right),$
we have by [10, Lemma 4.2] that $m_{\beta}$ is independent of the particular minimal enclosure $\mathcal{Y}_{\beta}$ in $\mathfrak{X}_{\alpha}.$ Therefore, we can combine Proposition 3.5, Theorem 3.8 and the discussion above to state the property
Proposition 4.1.
Consider a coin $(L,R).$ We have a decomposition
$\mathfrak{h}=\left(\bigoplus_{\alpha\in A}\mathcal{Y}_{\alpha}\right)\oplus\mathcal{X},$ (5)
where $A$ is a finite set, $\mathcal{Y}_{\alpha}$ is a subspace of $\mathfrak{h}$ being the support of exactly one invariant state $\tau_{\alpha}$ for $\mathcal{L},$ $\mathcal{X}$ contains no enclosure, and
  • $\textmd{Tr}\left(L\tau_{\alpha}L^{*}\right)=\dfrac{1}{2}\Rightarrow$ the walk is $\tau_{\alpha}$-recurrent for all $\tau\in\mathcal{D}(\mathcal{Y}_{\alpha})$;
  • $\textmd{Tr}\left(L\tau_{\alpha}L^{*}\right)\neq\dfrac{1}{2}\Rightarrow$ the walk is $\tau_{\alpha}$-transient for all $\tau\in\mathcal{D}(\mathcal{Y}_{\alpha}).$
Once the decomposition on Equation (5) is obtained, we can say that $\mathcal{Y}_{\alpha}$ is either recurrent or transient.
The following result is due to [10] and will be crucial to our last criterion. To state it, consider a coin $(L,R),$ and let $\mathcal{Y}$ be an enclosure for $\mathcal{L}.$ Denoting $p_{\mathcal{Y}}$ the orthogonal projection onto $\mathcal{Y},$ we can define the associated absorption operator (see [9]):
$A(\mathcal{Y}):=\lim_{n\rightarrow\infty}\mathcal{L}^{*n}(p_{\mathcal{Y}}).$
Corollary 4.2.
Consider a coin $(L,R)$ with the decomposition of $\mathfrak{h}$ given on Equation (5), and let $a_{\alpha}(\rho)=\mathbb{E}_{\rho}\left[\mathrm{Tr}\left(A(\mathcal{Y}_{\alpha})\rho_{0}\right)\right].$ Then we have the convergence in law
$\frac{X_{n}-X_{0}}{n}\stackrel{{\scriptstyle n\rightarrow\infty}}{{\longrightarrow}}\sum_{\alpha\in A}a_{\alpha}(\rho)\delta_{m_{\alpha}},$ (6)
where $\delta_{m_{\alpha}}$ is the Dirac measure concentrated in $m_{\alpha}.$
We remark that the convergence in law brought by corollary 4.2 can be upgraded to a convergence a.s., as demonstrated in [13, Theorem 3.5.2].
Let us start an HOQW induced by a coin $(L,R).$ There are two possibilities concerning recurrence when $\text{supp}(\rho_{0})\not\perp\mathcal{X}:$
[…]
$m=\sum_{\alpha\in A}a_{\alpha}(\rho)\delta_{m_{\alpha}}=\sum_{\alpha\in A}a_{\alpha}(\rho),$
as given by Equation (6), where the second equality follows from $m_{\alpha}\neq 0,$ since $(L,R)$ is $\mathcal{Y}_{\alpha}$ transient. Thus, $X_{n}$ converges either to -$\infty$ or $\infty,$ in probability, imposing that the walk is $\rho$-transient and thus $(L,R)$ is transient on $\mathcal{D}(\mathcal{X}).$
Therefore, we have shown the following recurrence criterion:
Theorem 4.3 (A Generalized Criteria).
Consider a coin $(L,R)$ of dimension $d.$ Consider the decomposition of $\mathfrak{h}$ as given in Equation (5), and $\tau_{1},\ldots,\tau_{m}$ the respectively invariant states of $\mathcal{L}$ in $\mathcal{Y}_{\alpha},\alpha=1,\ldots,m.$
$(1)\quad\mathrm{Tr}(L\tau_{j}L^{*})\neq\dfrac{1}{2}\;\forall j\in\{1,\ldots m\}\Rightarrow(L,R)\mbox{ is transient;}$
[…]
$\begin{cases}\rho\text{-transient for }\rho\in\mathcal{D}_{1}:=\mathcal{D}(\oplus_{k\in J^{C}}\mathcal{Y}_{k})\\ \rho\text{-recurrent for }\rho\in\mathcal{D}_{1}^{C}.\end{cases}$
[…]
$\begin{cases}\rho\text{-transient for }\rho\in\mathcal{D}_{2}:=\mathcal{D}(\oplus_{k\in J^{C}}\mathcal{Y}_{k}\oplus\mathcal{X})\\ \rho\text{-recurrent for }\mathcal{D}_{2}^{C}\end{cases}$
$(3)\quad\mathrm{Tr}(L\tau_{\alpha}L^{*})=\dfrac{1}{2}\;$ for every $\alpha\Rightarrow$ the walk is recurrent.
Remark $4.$
[…]
5 Examples - 1D
Example 5.1.
Our first examples will consider lazy HOQWs of dimension 2.
  1. 1.
    Let $m,n>0$ and assume that $b:=1-m^{2}-n^{2}\in(0,1).$ Then the coin $(L,B,R)$ induced by
    $L=\begin{bmatrix}0&0\\ 0&-n\end{bmatrix},\qquad B=\begin{bmatrix}-\sqrt{b}&m\\ m&\sqrt{b}\end{bmatrix},\qquad R=\begin{bmatrix}n&0\\ 0&0\end{bmatrix}$
    is well-defined.
    The only invariant state for the auxiliary map is $\rho_{\infty}=I/2.$ Therefore,
    $\mathrm{Tr}(L\rho_{\infty}L^{*})=\frac{n^{2}}{2}=\mathrm{Tr}(R\rho_{\infty}R^{*})\Rightarrow\;(L,B,R)\mbox{ is recurrent,}$
    by Item (1) of Theorem 3.10.
    It is worth noting that if we assume $m=0,$ then the walk is transient by Item (2) of Theorem 3.10. More explicitly, in this case, for the item (2.2), $|l_{1}|=0,\;|l_{2}|=n,\;|r_{1}|=n,\;|r_{2}|=0.$
    $\Diamond$
  2. 2.
    Let $x_{1},x_{2}\in\mathbb{C},$
    $|x_{1}|\leq\sqrt{\frac{2}{3}},\quad|x_{2}|\leq\sqrt{\frac{3}{4}},$
    then consider the coin $(L,B,R),$ where
    $L=\begin{bmatrix}\frac{1}{\sqrt{3}}&0\\ 0&\frac{1}{2}\end{bmatrix},\qquad B=\begin{bmatrix}\sqrt{\frac{2}{3}-x_{1}^{2}}&0\\ 0&\sqrt{\frac{3}{4}-x_{2}^{2}}\end{bmatrix},\qquad R=\begin{bmatrix}x_{1}&0\\ 0&x_{2}\end{bmatrix}$.
    We are in Item (2) of Theorem 3.10, thus we conclude that
    $(1)\quad|x_{1}|=\frac{1}{\sqrt{3}}\mbox{ and }|x_{2}|=\frac{1}{2}$ $\Rightarrow$ $(L,B,R)\mbox{ is recurrent,}$
    $(2)\quad|x_{1}|\neq\frac{1}{\sqrt{3}}\mbox{ and }|x_{2}|\neq\frac{1}{2}$ $\Rightarrow$ $(L,B,R)\mbox{ is transient,}$
[…]
    $\mbox{recurrent with respect to all densities but }\sigma,$
[…]
    $\mbox{recurrent with respect to all densities but }\sigma.$
$\Diamond$
Example 5.2.
Consider the Lazy HOQW induced by the following coin of dimension 3: $(L,B,R),$ where
$L=\frac{1}{8}\begin{bmatrix}2(1+\sqrt{2})&0&2(1-\sqrt{2})\\ 0&\sqrt{31}&0\\ 2(1-\sqrt{2})&0&2(1+\sqrt{2})\end{bmatrix},\quad B=\frac{1}{16}\begin{bmatrix}\sqrt{30}&2i&\sqrt{30}\\ 2&0&2\\ \sqrt{30}&-2i&\sqrt{30}\end{bmatrix},\quad R=\frac{1}{8}\begin{bmatrix}2(1-\sqrt{2})&0&2(1+\sqrt{2})\\ 0&\sqrt{31}&0\\ 2(1+\sqrt{2})&0&2(1-\sqrt{2})\end{bmatrix}$.
There is a unique invariant state $\rho_{\infty}$ given by
$\rho_{\infty}=\frac{1}{2}\begin{bmatrix}1&0&-1\\ 0&0&0\\ -1&0&1\end{bmatrix}$,
and $\mathrm{Tr}\left(L\rho_{\infty}L^{*}\right)=1/2=\mathrm{Tr}\left(R\rho_{\infty}R^{*}\right).$ Therefore, this coin is recurrent by Theorem 3.8.
$\Diamond$
Since the invariant state also depends on the choice of $B,$ it is natural to ask if for a given transient coin $(L,B,R),$ there is another $B^{\prime}$ that makes $(L,B^{\prime},R)$ to be recurrent.
Example 5.3.
Firstly, let
$L=\frac{1}{7}\begin{bmatrix}1&0\\ 0&2\end{bmatrix},\qquad B=\frac{1}{7}\begin{bmatrix}-1&1\\ 2&0\end{bmatrix},\qquad R=\frac{1}{7}\begin{bmatrix}1&1\\ 0&1\end{bmatrix}$.
Then
$\rho_{\infty}=\frac{1}{3}\begin{bmatrix}1&0\\ 0&2\end{bmatrix}$,
$\mathrm{Tr}\left(L\rho_{\infty}L^{*}\right)=3/7$ and $\mathrm{Tr}\left(R\rho_{\infty}R^{*}\right)=5/21.$ The coin $(L,B,R)$ is transient. However, if we take
$B^{\prime}=\frac{1}{\sqrt{35}}\begin{bmatrix}5&-1\\ 0&2\end{bmatrix}$,
then the invariant state is
$\rho^{\prime}_{\infty}=\begin{bmatrix}1&0\\ 0&0\end{bmatrix}$,
and $\mathrm{Tr}\left(L\rho_{\infty}L^{*}\right)=1/\sqrt{7}=\mathrm{Tr}\left(R\rho_{\infty}R^{*}\right).$ Thus the coin $(L,B^{\prime},R)$ is recurrent.
$\Diamond$
We remark that the choices of the coins above were taken in such a way that $L,B^{\prime}$ and $R$ had a common invariant eigenvector, namely $1,0]^{T},$ which was not an eigenvector of $B.$ This is due to the hard job of, given non-trivial $R$ and $L,$ to find $B^{\prime}$ in such a way that $(L,B^{\prime},R)$ is recurrent. Attempt were conduct using mathematic software with many $L$’s and $R$’s having simple entries, but even though the entries of $B$’s that made $(L,B,R)$ to be transient were very long to write in this paper.
Example 5.4.
Let us consider the coin given in [10, Section 6.2]:
$R=\begin{bmatrix}\sqrt{\frac{3}{8}}&0&0&0\\ -\sqrt{\frac{p_{1}}{2}}&\frac{1}{\sqrt{2}}&0&0\\ -\sqrt{\frac{p_{2}}{2}}&0&\frac{1}{\sqrt{2}}&0\\ \sqrt{\frac{2p_{3}}{3}}&0&0&\frac{1}{\sqrt{3}}\end{bmatrix},\qquad L=\begin{bmatrix}\frac{1}{2\sqrt{2}}&0&0&0\\ \sqrt{\frac{p_{1}}{2}}&\frac{1}{\sqrt{2}}&0&0\\ \sqrt{\frac{p_{2}}{2}}&0&\frac{1}{\sqrt{2}}&0\\ -\sqrt{\frac{p_{3}}{3}}&0&0&\frac{\sqrt{2}}{\sqrt{3}}\end{bmatrix}$,
where $p_{1},\;p_{2},p_{3}\geq 0$ and $p_{1}/2+p_{2}/2+p_{3}/2=1/2.$
The invariant states of the auxiliary map are of the form $\rho_{\infty}=\sigma_{1}+\sigma_{2},$ where
$\sigma_{1}=\omega_{1}\begin{bmatrix}0&0&0&0\\ 0&a&b&0\\ 0&b^{*}&1-a&0\\ 0&0&0&0\end{bmatrix},\quad\sigma_{2}=\omega_{2}\begin{bmatrix}0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&0&1\end{bmatrix},\quad\omega_{i},\omega_{j}\geq 0,\;\;\omega_{1}+\omega_{2}=1.$
We have
$m_{1}=\mathrm{Tr}(L\sigma_{1}L^{*})=\frac{1}{2},\qquad m_{2}=\mathrm{Tr}(L\sigma_{2}L^{*})=\frac{2}{3}\neq\frac{1}{2}.$
By item (2) of Theorem 4.3, the walk is transient for
$\rho=\begin{bmatrix}0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&0&1\end{bmatrix}$
and recurrent for all the remaining density operators on $\mathbb{C}^{4}$, for any appropriate choices of $p_{k},\;k=1,2,3.$
Example 5.5.
Consider an HOQW induced by the coin $(L,R),$ where
$L=\begin{bmatrix}\frac{\sqrt{5}}{5}&0&-\frac{\sqrt{5}}{5}\\ 0&\frac{2\sqrt{5}}{5}&\frac{\sqrt{5}}{10}\\ 0&0&\frac{1}{2}\end{bmatrix},\qquad R=\begin{bmatrix}\frac{2\sqrt{5}}{5}&0&\frac{\sqrt{5}}{10}\\ 0&\frac{\sqrt{5}}{5}&-\frac{\sqrt{5}}{5}\\ 0&0&\frac{1}{2}\end{bmatrix}$.
The invariant states of the auxiliary map are of the form $\rho_{\infty}=\sigma_{1}+\sigma_{2},$ where
$\sigma_{1}=\omega_{1}\begin{bmatrix}1&0&0\\ 0&0&0\\ 0&0&0\end{bmatrix},\quad\sigma_{2}=\omega_{2}\begin{bmatrix}0&0&0\\ 0&1&0\\ 0&0&0\end{bmatrix},\quad\omega_{i},\omega_{j}\geq 0,\;\;\omega_{1}+\omega_{2}=1.$
We have
$m_{1}=\mathrm{Tr}(L\sigma_{1}L^{*})=\frac{\sqrt{5}}{5}\neq\frac{1}{2},\qquad m_{2}=\mathrm{Tr}(L\sigma_{2}L^{*})=\frac{2\sqrt{5}}{5}\neq\frac{1}{2}.$
By item (1) of Theorem 4.3, the walk is transient.
Example 5.6.
Let us consider the coin $(L,R)$ of dimension 4 given by
$R=\begin{bmatrix}\frac{\sqrt{2}}{2}&-\frac{\sqrt{5}}{4}&0&\frac{1}{4}\\ 0&\frac{\sqrt{2}}{4}&0&0\\ 0&0&\frac{\sqrt{2}}{2}&0\\ 0&0&0&\frac{\sqrt{6}}{4}\end{bmatrix},\qquad L=\begin{bmatrix}\frac{\sqrt{2}}{2}&\frac{\sqrt{5}}{4}&0&-\frac{1}{4}\\ 0&\frac{1}{2}&0&\frac{\sqrt{5}}{4}\\ 0&0&-\frac{\sqrt{2}}{2}&0\\ 0&0&0&\frac{\sqrt{3}}{4}\end{bmatrix}$,
The invariant states of the auxiliary map are of the form $\rho_{\infty}=\sigma_{1}+\sigma_{2},$ where
$\sigma_{1}=\omega_{1}\begin{bmatrix}1&0&0&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\end{bmatrix},\quad\sigma_{2}=\omega_{2}\begin{bmatrix}0&0&0&0\\ 0&0&0&0\\ 0&0&1&0\\ 0&0&0&0\end{bmatrix},\quad\omega_{i},\omega_{j}\geq 0,\;\;\omega_{1}+\omega_{2}=1.$
We have
$m_{1}=\mathrm{Tr}(L\sigma_{1}L^{*})=\frac{1}{2},\qquad m_{2}=\mathrm{Tr}(L\sigma_{2}L^{*})=\frac{1}{2}.$
By item (3) of Theorem 4.3, the walk is recurrent.
6 2D Open Quantum Walks
We start this section with the continuous-time OQW.
6.1 Continuous-Time Open Quantum Walks
In order to define the CTOQWs, we recall that an operator semigroup $\mathcal{T}$ on a Hilbert space $\mathcal{K}$ is a family of bounded linear operators $(T_{t})$ acting on $\mathcal{K}$, $t\geq 0,$ such that
$T_{t}T_{s}=T_{t+s},\;\forall s,t\in\mathbb{R}^{+}\;\mbox{ and }\;T_{0}=I_{\mathcal{K}}.$
If $t\mapsto T_{t}$ is continuous for the operator norm of $\mathcal{K}$, then $\mathcal{T}$ is said to be uniformly continuous, which is equivalent to the following assertion [7]: There exists a bounded linear operator $L$ on $\mathcal{K}$ such that $T_{t}=e^{tL},\;\forall t\geq 0.$ In this case,
$L=\lim_{t\rightarrow\infty}\frac{1}{t}(T_{t}-I_{\mathcal{K}})$
and the operator $L$ is called the generator of $\mathcal{T}.$
[…]
Definition 6.1.
Let $V$ be a finite or countable infinite set and $\mathcal{H}$ be a Hilbert space of the form (2). A Continuous-time Open Quantum Walk in $V$ is an uniformly continuous QMS on $\mathcal{I}_{1}(\mathcal{H})$ with Lindblad operator of the form
$\begin{split}\mathcal{\mathcal{L}}:\mathcal{I}_{1}(\mathcal{H})&\rightarrow\mathcal{I}_{1}(\mathcal{H})\\ \rho&\mapsto-i[\textbf{H},\rho]+\sum_{i,j\in V}\left(S_{i}^{j}\rho S_{i}^{j^{*}}-\frac{1}{2}\{S_{i}^{j*}S_{i}^{j},\rho\}\right),\end{split}$ (7)
where $A,B]\equiv AB-BA$ and $\{A,B\}\equiv AB+BA.$
To rearrange Equation 7 We can write $S_{i}^{j}=R_{i}^{j}\otimes\lvert j\rangle\langle i\rvert$ for bounded operators $R_{i}^{j}\in\mathcal{B}(\mathfrak{h}_{i},\mathfrak{h}_{j}),$ the bounded operators H and $S_{i}^{j}$ on $\mathcal{H}$ are of the form $\textbf{H}=\sum_{i\in V}H_{i}\otimes\lvert i\rangle\langle i\rvert$, $H_{i}$ is self-adjoint on $\mathfrak{h}_{i},$ and $\sum_{i,j\in V}S_{i}^{j*}S_{i}^{j}$ converges in the strong sense.
Letting $\rho=\sum_{i\in V}\rho(i)\otimes\lvert i\rangle\langle i\rvert\in\mathcal{D},$ denoting
$e^{t\mathcal{L}}(\rho)=\mathcal{T}_{t}(\rho)=\sum_{i\in V}\rho_{t}(i)\otimes\lvert i\rangle\langle i\rvert,\forall t\geq 0,$
we have
$\frac{d}{dt}\rho_{t}(i)=-i[H_{i},\rho_{t}(i)]+\sum_{j\in V}\left(R_{j}^{i}\rho_{t}(j)R_{j}^{i^{*}}-\frac{1}{2}\{R_{i}^{j*}R_{i}^{j},\rho_{t}(i)\}\right).$
Equation (7) can be written as given in [3, Equation 18.7]:
$\mathcal{L}(\rho)=\sum_{i\in V}\left(G_{i}\rho(i)+\rho(i)G_{i}^{*}+\sum_{j\in V}R_{j}^{i}\rho(j)R_{j}^{i*}\right)\otimes\lvert i\rangle\langle i\rvert,\qquad G_{i}=-iH_{i}-\frac{1}{2}\sum_{j\in V}R_{i}^{j*}R_{i}^{j}.$ (8)
In this paper we deal with the quantum trajectories of CTOQWs, which are more complex than the trajectories od discrete-time HOQWs. Thus, we recall that for a givenfinite or countably infinite set of vertices $V,$ a CTOQW is a stochastic process evolving on a Hilbert space of the form (2), the label $i\in V$ represents the position of the walker and when the walker is located at $i\in V,$ its internal state is encoded in $\mathfrak{h}_{i},$ that is, $\mathfrak{h}_{i}$ describes the internal degrees of freedom of the walker when it is at site $i\in V.$
Given $i\in V,$ we denote the set of density operators concentrated at vertex $i$ by
$\mathcal{D}_{i}=\{\rho\in\mathcal{I}(\mathfrak{h}_{i}),\;\rho\geq 0,\;\mathrm{Tr}(\rho)=1\}.$
Thus, for $\rho\in\mathcal{D}_{i},$ we have $\rho\otimes\lvert i\rangle\langle i\rvert\in\mathcal{D}.$
Starting the walk on site $\lvert i\rangle$ with initial density operator $\rho\in\mathcal{D}_{i},$ the quantum measurement of the “position” gives rise to a probability distribution $p_{0}$ on $V,$ such that
$p_{0}(j)=\mathbb{P}(\mbox{the quantum particle is in site}\;\lvert j\rangle)=\mathrm{Tr}(\rho(j)),$
and for evolution on time $t\geq 0,$
[…]
where
$e^{t\mathcal{L}}(\rho)=\sum_{k\in V}\rho_{t}(i)\otimes\lvert k\rangle\langle k\rvert$.
6.2 Quantum Trajectories
Let $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{P})$ be a probability space where independent Poisson point processes $N^{ij},i,j\in V,\;i\neq j$ ($N^{ii}=0$ by convention) on $\mathbb{R}^{2}$ are defined. The jump from site $i$ to site $j$ on the graph $V$ will be governed by these Poisson point processes [3]:
Definition 6.2.
Consider a CTOQW with generator of the form (8) and an initial density operator $\mu=\sum_{i\in V}\rho(i)\otimes\lvert i\rangle\langle i\rvert\in\mathcal{D}.$ The quantum trajectory describing the indirect measurement of the position of the CTOQW is the Markov chain represented by the density operators $(\mu_{t})_{t\geq 0}$ such that $\mu_{0}=\rho_{0}\otimes\lvert X_{0}\rangle\langle X_{0}\rvert$, where $X_{0}$ and $\rho_{0}$ are random variables with distribution
$\mathbb{P}\left((X_{0},\rho)=\left(i,\frac{\rho(i)}{\mathrm{Tr}(\rho(i))}\right)\right)=\mathrm{Tr}\left(\rho(i)\right)\mbox{ for all }i\in V,$
and $\mu_{t}=:\rho_{t}\otimes\lvert X_{t}\rangle\langle X_{t}\rvert$ satisfies the stochastic differential equation
$\begin{split}\mu_{t}=&\mu_{0}+\int_{0}^{t}M(\mu_{s^{-}})ds\\ +&\sum_{ij}\int_{0}^{t}\int_{\mathbb{R}}\left(\frac{S_{i}^{j}\mu_{s^{-}}S_{i}^{j*}}{\mathrm{Tr}(SS_{i}^{j*}\mu_{s^{-}}S_{i}^{j*})}-\mu_{s^{-}}\right)1_{0<y<\mathrm{Tr}(S_{i}^{j}\mu_{s^{-}}S_{i}^{j*})}N^{ij}(dy,ds)\end{split}$ (10)
for all $t\geq 0,$ where
$M(u)=\mathcal{L}(u)-\sum_{ij}\left(S_{i}^{j}\mu S_{i}^{j*}-\mu\mathrm{Tr}(S_{i}^{j}\mu S_{i}^{j*})\right).$
Hence, for a fixed $\mu=\sum_{i}\rho(i)\otimes\lvert i\rangle\langle i\rvert\in\mathcal{D},$
$M(\mu)=\sum_{i}\left(G_{i}\rho(i)+\rho(i)G_{i}^{*}-\rho(i)\mathrm{Tr}\left(G_{i}\rho(i)+\rho(i)G_{i}^{*}\right)\right)\otimes\lvert i\rangle\langle i\rvert$.
For simplicity, suppose $X_{0}=i_{0}$ for some $i_{0}\in V$ and $\rho_{0}\in V(\mathfrak{h}_{i_{0}}).$ For all $t\geq 0,$ consider the solution of Equation (10)
$\eta_{t}=\rho_{0}+\int_{0}^{t}\left(G_{i_{0}}\eta_{s}+\eta_{s}G_{i_{0}}^{*}-\eta_{s}\mathrm{Tr}\left(G_{i_{0}}\eta_{s}+\eta_{s}G_{i_{0}}^{*}\right)\right)ds,$
which is a density operator on $\mathfrak{h}_{i_{0}}.$ For $j\neq i_{0},$ define
$T_{1}^{j}=\mathrm{inf}\{t\geq 0;N^{i_{0},j}\left({u,y|0\leq u\leq t,0\leq y\leq\mathrm{Tr}(R_{i_{0}}^{j}\eta_{u}R_{i_{0}}^{j*})}\right)\geq 1\}.$
Since the random variables $T_{1}^{j}$ are mutually independent and nonatomic, we can define $T_{1}=\inf_{j\neq i_{0}}\{T_{1}^{j}\}$ once there exists a unique $j\in V$ such that $T_{1}^{j}=T_{1}.$ The random variable $T_{1}$ is said to be the first jump time of the CTOQW conditional on $X_{0}=i_{0}.$
The first jump time to site $\lvert j\rangle$ is then denoted by $T_{1}^{j}$ and has distribution
$\mathbb{P}(T_{1}^{j}>\varepsilon)=e^{-\int_{0}^{\varepsilon}\mathrm{Tr}(R_{i_{0}}^{j}\eta_{u}R_{i_{0}}^{j*})du},$
thus
$\mathbb{P}(T_{1}\leq\varepsilon)\leq\varepsilon\sum_{j\neq i_{0}}\|R_{i_{0}}^{j*}R_{i_{0}}^{j}\|.$
The strongly convergence of $\sum_{ij}S_{i}^{j*}S_{i}^{j}$ implies that $\mathbb{P}(T_{1}>0)=1.$ Thereby, on $0,T_{1}],$ we can define the solution $(X_{t},\rho_{t})_{t\geq 0}$ as
$(X_{t},\rho_{t})$ $=$ $(i_{0},\eta_{t})\mbox{ for }t\in[0,T_{1})\mbox{ and }$
$(X_{T_{1}},\rho_{T_{1}})$ $=$ $\left(j,\frac{R_{i}^{j}\eta_{T_{1}-}R_{i}^{j*}}{\mathrm{Tr}(R_{i}^{j}\eta_{T_{1}-}R_{i}^{j*})}\right)\;\mbox{ if }T_{1}=T_{1}^{j}.$
Now we solve
$\eta_{t}=\rho_{T_{1}}+\int_{0}^{t}\left(G_{j}\eta_{s}+\eta_{s}G_{j}^{*}-\eta_{s}\mathrm{Tr}\left(G_{j}\eta_{s}+\eta_{s}G_{j}^{*}\right)\right)ds$
and then obtain the second jump time $T_{2}.$ So on we obtain an increasing sequence of jumps $(T_{n})_{n}$ with $\lim_{n\rightarrow\infty}T_{n}=\infty$ almost surely (see section 18.2.3 of [3] for more details). This means that the walk does not explode, thus the walker has a finite number of jumps in a finite interval. For details concerning explosions of classical Markov chains, see [21, Section 2.2].
6.3 Recurrence of CTOQWs
Movin forward, we deal with the following class of 2-dimensional CTOQWs, that is, we will consider the nearest-neighbor quantum random walk on $V=\mathbb{Z}^{2}.$
Definition 6.3.
Consider a CTOQW on $V=\mathbb{Z}^{2}$. If there exists $A_{1},A_{2},A_{3},A_{4},H\in\mathcal{B}(\mathbb{C}^{n})$ such that
$H_{k}=H\;\forall k\in\mathbb{Z}^{2},\quad A_{1}=R_{(i,j)}^{(i+1,j)},\;A_{2}=R_{(i,j)}^{(i,j+1)},\;A_{3}=R_{(i,j)}^{(i-1,j+1)},\;A_{4}=R_{(i,j)}^{(i,j-1)},$
and $R_{x}^{y}=0_{n}$ for the remaining $x,y\in\mathbb{Z}^{2},$ then the CTOQW is said to be induced by a coin $(A,H)$ of dimension $n.$
For this kind of CTOQW we can let $\mathfrak{h}_{i}=\mathfrak{h}$ for all $i\in\mathbb{Z}^{2}.$ If $\dim(\mathfrak{h})=n,$ then $A_{1},A_{2},A_{3},A_{4}$ and $H$ can be represented by square matrices of order $n.$
The auxiliary map of a coin $(A,H)$ is defined as
$\mathbb{L}(\rho)=G\rho+\rho G^{*}+\sum_{j=1}^{4}A_{j}\rho A_{j}^{*},\qquad G=-iH-\frac{1}{2}\sum_{k=1}^{4}A_{k}^{*}A_{k}.$
The map $\mathbb{L}$ has at least one stationary state $\rho_{\infty},$ that is, $\mathbb{L}(\rho_{\infty})=0$ [5]. When the stationary state is unique, we define
$m=\begin{bmatrix}m_{1}\\ m_{2}\end{bmatrix},\quad m_{1}=\mathrm{Tr}\left(A_{1}\rho_{\infty}A_{1}^{*}\right)-\mathrm{Tr}\left(A_{3}\rho_{\infty}A_{3}^{*}\right),\quad m_{2}=\mathrm{Tr}\left(A_{2}\rho_{\infty}A_{2}^{*}\right)-\mathrm{Tr}\left(A_{4}\rho_{\infty}A_{4}^{*}\right).$
To demonstrate the recurrence criteria of this work for CTOQWs, we recall the following statistical results for those random walks.
Theorem 6.4 ([6]).
Consider a coin $(A,H)$, where its auxiliary map has a unique invariant state $\rho_{\infty}.$ Then
$\frac{X_{t}-tm}{\sqrt{t}}\stackrel{{\scriptstyle t\rightarrow\infty}}{{\longrightarrow}}N(0,\sigma^{2})$ (11)
for some $\sigma>0.$
Proposition 6.5 (Law of Large Numbers for CTOQWs ).
[19, Theorem 15] If the auxiliary map of a coin $(D)$ has a unique stationary state, then
$\frac{X_{t}}{t}\rightarrow m\quad\mbox{a.s}.$
Let us recall the classical definition of recurrence of CTOQWs. For this, $p_{ji;\rho}(t)$ will denote the probability of being at site $j$ at time $t,$ given that we started at site $i,$ with initial density $\rho$ concentrated at $i$:
[…]
Therefore, the walk starts with a density operator $\rho$ concentrated at some vertex $\lvert i\rangle,$ takes the evolution up to time $t>0$ through the quantum Markov semigroup generated by $\mathcal{L},$ producing a new density operator
$\rho_{t}=\sum_{k}\rho_{t}(k)\otimes\lvert k\rangle\langle k\rvert=e^{t\mathcal{L}}(\rho\otimes\lvert i\rangle\langle i\rvert),\quad\mathrm{Tr}\left(\sum_{k}\rho_{t}(k)\right)=1.$
Now we project $\rho_{t}$ onto the subspace generated by vertex $\lvert j\rangle$ giving
$e^{t\mathcal{L}}(\rho\otimes\lvert i\rangle\langle i\rvert)(I\otimes\lvert j\rangle\langle j\rvert)=\rho_{t}(j)\otimes\lvert j\rangle\langle j\rvert$,
which represents the data concentrated at vertex $\lvert j\rangle$ at time $t.$
Let $i\in V,\rho\in\mathcal{D}_{i}.$ We say that vertex $i$ is
  • $\rho$-recurrent if
    $\int_{0}^{\infty}p_{ii;\rho}(t)dt=\infty.$
    Otherwise, $i$ is said to be $\rho$-transient
  • recurrent, if $i$ is $\rho$-recurrent for all $\rho\in\mathcal{D}_{i};$
  • transient, if $i$ is $\rho$-transient for all $\rho\in\mathcal{D}_{i}.$
By definition, $i$ is $\rho$-recurrent when the mean number of returns to $i$ is infinite.
Definition 6.6.
A CTOQW is said to be:
  • recurrent if every vertex is recurrent;
  • transient if every vertex is transient.
Proposition 6.7.
A CTOQW induced by a coin is $\rho$-recurrent if and only if it is $\rho_{T_{n}}$-recurrent for any possible jump time $T_{n}.$
Proof.
Note that if we start the walk with $(X_{0},\rho_{0})=(i,\rho),$ then at some time $t>0,$ we have $(X_{t},\rho_{t})=(j,\rho_{T}).$ If we refresh the walk at this instant, then the walk is still recurrent with respect to the density $\rho_{t},$ by the homogeneity in space. The other hand follows the same idea, assuming by contraposition that the walk is $\rho$-transient.
$\square$
Proposition 6.8.
If $(A,H)$ is $\tau$-recurrent for some density $\tau\in\mathcal{D}$ and $\mathfrak{h}$ has at most one non-trivial invariant subspace for $\Phi:\mathfrak{h}\rightarrow\mathfrak{h}$ given by
$\Phi(.)=\sum_{k}A_{k}(.)A_{k}^{*},$
then $(A,H)$ is recurrent.
Proof.
Let us start supposing that $(A,H)$ is recurrent with respect to some density operator $\tau$ with support on $\mathcal{V}.$ Let $\rho\in\mathcal{D}$ be another density operator. Any $\rho_{T_{n}}$ is of the form
$\rho_{T_{n}}=\frac{A_{k}\eta A_{k}^{*}}{\mathrm{Tr}\left(A_{k}\eta A_{k}^{*}\right)}\text{ for some }k\in\{1,2,3,4\},$ (12)
[…]
$\sigma:=\rho_{T_{n}}-\frac{\tau}{a}\geq 0.$
Therefore,
$\frac{1}{\mathrm{Tr}(\sigma)}\int_{0}^{\infty}p_{ii;\rho_{T_{n}}}(t)\;\text{d}t=\int_{0}^{\infty}p_{ii;\frac{\sigma}{\mathrm{Tr}(\sigma)}}(t)\;\text{d}t+\frac{1}{a\mathrm{Tr}(\sigma)}\int_{0}^{\infty}p_{ii;\tau}(t)\;\text{d}t=\infty.$
By Proposition 6.7, $(A,H)$ is $\rho$-recurrent.
$\square$
6.4 Recurrence Criterion
This section is devoted to exhibit a recurrence criterion for finite dimensional coins $(A,H),$ where the auxiliary map has at most one common invariant subspace. The result uses a Central Limit Theorem for CTOQWs and the following Lemma.
Lemma 6.9.
Let us consider a homogeneous CTOQW on $\mathbb{Z}^{d}.$ Let $\varepsilon>0$ and $m\geq 1$ be an integer. Then, for every $\rho\in\mathcal{D}$,
$\int_{0}^{\infty}\mathbb{P}_{0,\rho}(|X_{t}|<m\varepsilon)dt\leq(2m)^{d}\cdot\max_{\sigma\in\mathcal{D}}\int_{0}^{\infty}\mathbb{P}_{0,\sigma}(|X_{t}|<\varepsilon)dt.$
Proof. Firstly, note that
$\int_{0}^{\infty}\mathbb{P}_{0,\rho}\left(|X_{t}|<m.\varepsilon\right)dt\leq\int_{0}^{\infty}\sum_{k}\mathbb{P}_{0,\rho}\left(X_{t}\in k.\varepsilon+[0,\varepsilon)^{d}\right)dt,$ (13)
where the sum over $k$ is on the interval $\{-m,-m+1,\ldots,m-2,m-1\}^{d}$.
Let
$T_{k}=\inf\left\{z\geq 0:X_{z}\in k.\varepsilon+\left[0,\varepsilon\right)^{d}\right\}\;\text{ and }\;\zeta_{\rho}^{k,\varepsilon}=\int_{0}^{\infty}\mathbb{P}_{0,\rho}\left(X_{t}\in k.\varepsilon+[0,\varepsilon)^{d}\right)dt.$
. By the Law of Total Probability, and Fubini’s Theorem, we can rewrite $\zeta_{\rho}^{k,\varepsilon}$ as
$\begin{split}\zeta_{\rho}^{k,\varepsilon}=&\int_{0}^{\infty}\int_{0}^{t}\mathbb{P}_{0,\rho}\left(X_{t}\in k.\varepsilon+[0,\varepsilon)^{d},\;T_{k}=y\right)dy\;dt\\ =&\int_{0}^{\infty}\int_{0}^{\infty}\mathbb{P}_{0,\rho}(|X_{t}-X_{y}|<\varepsilon,T_{k}=y)dy\;dt\\ =&\int_{0}^{\infty}\int_{y}^{\infty}\mathbb{P}_{0,\rho}(|X_{t}-X_{y}|<\varepsilon,T_{k}=y)dt\;dz\\ =&\int_{0}^{\infty}\mathbb{P}_{0,\rho}(T_{k}=z)\int_{y}^{\infty}\mathbb{P}_{0,\rho}(|X_{t}-X_{y}|<\varepsilon)dt\;dy,\end{split}$ (14)
[…]
Denote $\mathbb{P}_{0,\rho}^{k,y}=\mathbb{P}_{0,\rho}\left(T_{k}=y\right).$ An application of Fubini’s Theorem gives
$\begin{split}\zeta_{\varepsilon,k}^{\rho}=&\int_{0}^{\infty}\mathbb{P}_{0,\rho}^{k,y}\int_{y}^{\infty}\mathbb{P}_{0,\rho}\left(|X_{t}-X_{y}|<\varepsilon\right)dtdy\\ =&\int_{0}^{\infty}\mathbb{P}_{0,\rho}^{k,y}\int_{y}^{\infty}\sum_{\sigma\in\mathcal{D},j\in\mathbb{Z}}\mathbb{P}_{0,\rho}\left(|X_{t}-X_{y}|<\varepsilon,(X_{y},\rho_{y})=(j,\sigma)\right)dtdy\\ =&\int_{0}^{\infty}\mathbb{P}_{0,\rho}^{k,y}\sum_{\sigma\in\mathcal{D},j\in\mathbb{Z}}\int_{y}^{\infty}\mathbb{P}_{0,\rho}\left(|X_{t}-X_{y}|<\varepsilon|(X_{y},\rho_{y})=(j,\sigma)\right)\mathbb{P}_{0,\rho}\left((X_{y},\rho_{y})=(j,\sigma)\right)dtdy\\ \zeta_{\varepsilon,k}^{\rho}=&\int_{0}^{\infty}\mathbb{P}_{0,\rho}^{k,y}\sum_{\sigma\in\mathcal{D},j\in\mathbb{Z}}\int_{y}^{\infty}\mathbb{P}_{j,\sigma}\left(|X_{t-y}-X_{0}|<\varepsilon\right)dt\mathbb{P}_{0,\rho}\left((X_{y},\rho_{y})=(j,\sigma)\right)dy\\ =&\int_{0}^{\infty}\mathbb{P}_{0,\rho}^{k,y}\sum_{\sigma\in\mathcal{D},j\in\mathbb{Z}}\int_{0}^{\infty}\mathbb{P}_{j,\sigma}\left(|X_{s}-j|<\varepsilon\right)ds\mathbb{P}_{0,\rho}\left((X_{y},\rho_{y})=(j,\sigma)\right)dy\\ =&\int_{0}^{\infty}\mathbb{P}_{0,\rho}^{k,y}\sum_{\sigma\in\mathcal{D}}\int_{0}^{\infty}\mathbb{P}_{0,\sigma}\left(|X_{s}|<\varepsilon\right)ds\sum_{j\in\mathbb{Z}}\mathbb{P}_{0,\rho}\left((X_{y},\rho_{y})=(j,\sigma)\right)dy\\ \leq&\int_{0}^{\infty}\mathbb{P}_{0,\rho}^{k,y}\max_{\tau\in\mathcal{D}}\int_{0}^{\infty}\mathbb{P}_{0,\tau}\left(|X_{s}|<\varepsilon\right)ds\sum_{\sigma\in\mathcal{D},j\in\mathbb{Z}}\mathbb{P}_{0,\rho}\left((X_{y},\rho_{y})=(j,\sigma)\right)dy\\ =&\int_{0}^{\infty}\mathbb{P}_{0,\rho}^{k,y}dy\max_{\tau\in\mathcal{D}}\int_{0}^{\infty}\mathbb{P}_{0,\tau}\left(|X_{s}|<\varepsilon\right)ds\\ =&\max_{\tau\in\mathcal{D}}\int_{0}^{\infty}\mathbb{P}_{0,\tau}\left(|X_{s}|<\varepsilon\right)ds.\end{split}$
Therefore, Equation (13) and the inequality above give
$\int_{0}^{\infty}\mathbb{P}_{0,\rho}(|X_{t}|<m\varepsilon)dt\leq\sum_{k}\zeta_{\rho}^{k,\varepsilon}\leq\sum_{k}\max_{\tau\in\mathcal{D}}\int_{0}^{\infty}\mathbb{P}_{0,\tau}(|X_{s}|<\varepsilon)ds=(2m)^{d}\cdot\max_{\tau\in\mathcal{D}}\int_{0}^{\infty}\mathbb{P}_{0,\tau}(|X_{s}|<\varepsilon)ds.$
Theorem 6.10.
If $X_{t}$ denotes a CTOQW on $\mathbb{Z}^{2}$ and $X_{t}/t^{1/2}\Rightarrow$ a Gaussian measure, then the CTOQW is recurrent for some density.
Proof.
Let us denote $u(t,m)=\mathbb{P}(|X_{t}|<m).$ By Lemma 6.9,
$\max_{\tau\in D}\int_{0}^{\infty}\mathbb{P}(|X_{t}|<1)dt=\max_{\tau\in D}\int_{0}^{\infty}u(t,1)dt\geq(4m^{2})^{-1}\int_{0}^{\infty}u(t,m)dt.$
Assume that $m/\sqrt{t}\rightarrow c,$ then
$u(t,m)\rightarrow\int_{[-c,c]^{2}}f(x)dx,$ (15)
where $f(x)$ represents the density of the limiting normal distribution. We denote the right-hand side of Equation (15) by $g(c)$ and let $t=\theta m^{2}$ to obtain
$u(\theta m^{2},m)\rightarrow g(\theta^{-1/2}).$
Now we use the chenge of variables
$m^{-2}\int_{0}^{\infty}u(t,m)dt=\int_{0}^{\infty}u(\theta m^{2},m)d\theta,$
let $m\rightarrow\infty,$ and use Fatou’s lemma to get
$\liminf_{m\rightarrow\infty}(4m^{2})^{-1}\int_{0}^{\infty}u(t,m)dt\geq 4^{-1}\int_{0}^{\infty}g(\theta^{-1/2})d\theta.$ (16)
Since the normal density is positive and continuous at $0,$
$g(c)=\int_{[-c,c]^{2}}n(x)dx\sim n(0)(2c)^{2}$
as $c\rightarrow 0.$ So $g(\theta^{-1/2})\sim 4n(0)/\theta$ as $\theta\rightarrow\infty,$ the integral in Equation (16) diverges. As a result, $\max_{\tau\in D}\int_{0}^{\infty}\mathbb{P}(|X_{t}|<1)dt=\infty,$ leading us to conclude that the walk is $\tau$-recurrent.
$\square$
Note that the maximum in Lemma 6.9 can be assumed to be taken over the jump times $(T_{n})_{n}.$ Therefore, by Proposition 6.7 and Theorem 6.10 we obtain the following result.
Theorem 6.11.
If $X_{t}$ denotes a CTOQW induced by a coin $(A,H),$ there is at most one minimal enclosure for $\Phi$ and $X_{t}/t^{1/2}\Rightarrow$ a Gaussian measure, then the CTOQW is recurrent.
We already have a sufficient condition to recurrence given by the Theorem 6.11. For transience, we also have a sufficient condition:
Theorem 6.12.
Let us consider a CTOQW induced by a coin $(A,H),$ having at most one minimal enclosure, and let $\rho_{\infty}$ be the unique stationary state of the auxiliary map. If $m\neq 0,$ then the CTOQW is transient.
Proof.
Let us start the CTOQW at some vertex $\lvert i\rangle\in\mathbb{Z}^{d}$ with initial density operator $\rho.$ By the Law of Large Numbers, Proposition 6.5,
$\frac{X_{t}}{t}\rightarrow m\neq 0^{d}.$
This means that at least one coordinate of $X_{t}$ is unbounded, from where we conclude that
$\mathbb{P}\left(X_{t}=X_{0}\;\text{i.o.}\right)=0.$
Hence, the CTOQW is transient.
$\square$
Finally, we can present a recurrence criteria for CTOQWs on $\mathbb{Z}^{2}$ induced by a coin $(A,H).$
Theorem 6.13 (Recurrence criteria for CTOQWs on $\mathbb{Z}^{2}$).
Let us consider a CTOQW induced by a coin $(A,H),$ having at most one minimal enclosure, and let $\rho_{\infty}$ be the unique stationary state of the auxiliary map. Then
  • $m=0\;\Rightarrow\;$ the CTOQW is recurrent;
  • $m\neq 0\;\Rightarrow\;$ the CTOQW is transient.
Proof.
For $m\neq 0,$ the result is given in Theorem 6.12. It remains to proof that $m=0$ implies that the CTOQW is recurrent. Firstly, if $m=0,$ then by Theorem 6.4
$\frac{X_{t}}{\sqrt{t}}=\frac{X_{t}-t.m}{\sqrt{t}}\Rightarrow\text{ a non-degenerate normal distribution,}$
thus we use Theorem 6.10 to obtain a density operator $\tau$ such that the CTOQW is $\tau$-recurrent. Therefore, $(A,H)$ is recurrent by Theorem 6.11.
$\square$
6.5 Recurrence criteria for HOQWs
6.5.1 Criterion for auxiliary map having at most one minimal enclosure
To present a complete criterion for the recurrence of finite-dimensional coins $(A)$, we will start this section with an easily verifiable criterion based on the quantum jump chains associated with CTOQWs, as presented in [20]. The complete recurrence criterion will be developed from a decomposition of the Hilbert space $\mathfrak{h}$ into subspaces that support invariant states. The central idea is identifying and characterizing these subspaces, thus we will be able to determine the recurrence of quantum walks by analyzing the evolution within each invariant component. Furthermore, these invariant states will play an indispensable role in formulating the criterion.
Let $\Pi$ denote a discrete HOQW given by
$\Pi(\rho)=\sum_{i,j\in V}M_{j}^{i}\rho(j)M_{j}^{i*}\otimes\lvert i\rangle\langle i\rvert$.
This HOQW is the open quantum jump chain of the CTOQW with Lindblad generator
$\mathbb{L}=\Psi-I,\;\;\Psi(\rho)=\sum_{i,j\in V}B_{j}^{i}\rho(j)B_{j}^{i*},\;B_{j}^{i}=M_{j}^{i}\otimes\lvert i\rangle\langle j\rvert$, (17)
where $I$ is the identity operator acting on $\mathcal{H}$ [20].
Theorem 6.14.
[20, Theorem 5.6] A vertex $i\in V$ is $\rho$-recurrent for the HOQW $\Pi$ if and only if it is $\rho$-recurrent for the CTOQW generated by $\mathbb{L}$ given by Equation (17).
We consider an HOQW with a set of vertices $V=\mathbb{Z}^{2}.$ Analogously to the continuous-time version, if there exists $D_{1},D_{2},D_{3},D_{4}\in\mathcal{B}(\mathbb{C}^{n})$ such that
$D_{1}=M_{(i,j)}^{(i+1,j)},\;D_{2}=M_{(i,j)}^{(i,j+1)},\;D_{3}=M_{(i,j)}^{(i-1,j+1)},\;D_{4}=M_{(i,j)}^{(i,j-1)},$
and $M_{x}^{y}=0_{n}$ for the remaining $x,y\in\mathbb{Z}^{2},$ then the HOQW $\Pi$ is said to be induced by a coin $(D)$ of dimension $n.$
The HOQW induced by a coin $(D)$ is the homogeneous, nearest-neighbor HOQW on the grid, with the right, up, left and down transitions given by $D_{1},D_{2},D_{3},D_{4},$ respectively. Therefore, the quantum trajectory of those walks, starting from a state $\tau$ of the form $\tau=\sum_{i\in\mathbb{Z}}\rho(i)\otimes|i\rangle\langle i|,$ is any path generated by the Markov chain $(X_{n},\tau_{n})_{n\geq 0}$, where $X_{n}$ denotes the position of the particle at time $n$ and $\tau_{n}$ its internal degree. The transition probabilities are given by
$\begin{split}\mathbb{P}\left(\,(X_{n+1},\tau_{n+1})=\left((i+1,j),\frac{D_{1}\sigma D_{1}^{*}}{\mathrm{Tr}(D_{1}\sigma D_{1}^{*})}\right)\;\Bigg|\;(X_{n},\rho_{n})=((i,j),\sigma)\,\right)&=\mathrm{Tr}(D_{1}\sigma D_{1}^{*}),\\ \mathbb{P}\left(\,(X_{n+1},\tau_{n+1})=\left((i,j+1),\frac{D_{2}\sigma D_{2}^{*}}{\mathrm{Tr}(D_{2}\sigma D_{2}^{*})}\right)\;\Bigg|\;(X_{n},\rho_{n})=((i,j),\sigma)\,\right)&=\mathrm{Tr}(D_{2}\sigma D_{2}^{*}),\\ \mathbb{P}\left(\,(X_{n+1},\tau_{n+1})=\left((i-1,j),\frac{D_{3}\sigma D_{3}^{*}}{\mathrm{Tr}(D_{3}\sigma D_{3}^{*})}\right)\;\Bigg|\;(X_{n},\rho_{n})=((i,j),\sigma)\,\right)&=\mathrm{Tr}(D_{3}\sigma D_{3}^{*}),\\ \mathbb{P}\left(\,(X_{n+1},\tau_{n+1})=\left((i,j-1),\frac{D_{4}\sigma D_{4}^{*}}{\mathrm{Tr}(D_{4}\sigma D_{4}^{*})}\right)\;\Bigg|\;(X_{n},\rho_{n})=((i,j),\sigma)\,\right)&=\mathrm{Tr}(D_{4}\sigma D_{4}^{*}),\end{split}$
for every $(i,j)\in\mathbb{Z}^{2}$, $\sigma\in\mathcal{D}$, and initial law
$\mathbb{P}\left(\,(X_{0},\rho_{0})=\left((i,j),\frac{\rho(i,j)}{\mathrm{Tr}\,\rho(i,j)}\right)\;\right)=\mathrm{Tr}(\rho(i,j)),$
and all other transition probabilities are null. The graph of this kind of HOQW is represented in Figure 3.
$D_{2}$$D_{2}$$D_{2}$$D_{2}$$D_{2}$$D_{2}$$D_{2}$$D_{2}$$D_{2}$$D_{2}$$D_{2}$$D_{2}$$D_{1}$$D_{1}$$D_{1}$$D_{1}$$D_{1}$$D_{1}$$D_{1}$$D_{1}$$D_{1}$$D_{1}$$D_{1}$$D_{1}$$D_{4}$$D_{4}$$D_{4}$$D_{4}$$D_{4}$$D_{4}$$D_{4}$$D_{4}$$D_{4}$$D_{4}$$D_{4}$$D_{4}$$D_{3}$$D_{3}$$D_{3}$$D_{3}$$D_{3}$$D_{3}$$D_{3}$$D_{3}$$D_{3}$$D_{3}$$D_{3}$$D_{3}\vdots\vdots\vdots\vdots\ldots\vdots\vdots\vdots\vdots$
Figure 3: Homogeneous Open Quantum walk on $\mathbb{Z}^{2}$.
Now we can give a recurrence criteria for HOQWs on $\mathbb{Z}^{2}.$ To this, note that if $\rho$ is an invariant state for the auxiliary map $\Phi$ of $(D),$ defined as
$\Phi(.)=\sum_{j=1}^{4}D_{j}(.)D_{j}^{*},$
then it is a stationary state for its open quantum jump chain $\mathbb{L}.$ Consequently, if $\Phi$ has at most one minimal enclosure, then $\Phi$ has a unique invariant state $\rho_{\infty}$ and we can let
$m=\begin{bmatrix}m_{1}\\ m_{2}\end{bmatrix},\quad m_{1}=\mathrm{Tr}\left(D_{1}\rho_{\infty}D_{1}^{*}\right)-\mathrm{Tr}\left(D_{3}\rho_{\infty}D_{3}^{*}\right),\quad m_{2}=\mathrm{Tr}\left(D_{2}\rho_{\infty}D_{2}^{*}\right)-\mathrm{Tr}\left(D_{4}\rho_{\infty}D_{4}^{*}\right)$
to obtain, as a consequence of Theorems 6.13 and 6.14, the following criterion.
Theorem 6.15 (Recurrence criterion for HOQWs on $\mathbb{Z}^{2}$).
Let us consider an HOQW induced by a coin $(D),$ where the map
$\Phi(.)=\sum_{j=1}^{4}D_{j}(.)D_{j}^{*}$
has at most one minimal enclosure, and let $\rho_{\infty}$ be the unique invariant state of $\Phi.$ Then
  • $m=0\;\Rightarrow\;$ the HOQW is recurrent;
  • $m\neq 0\;\Rightarrow\;$ the HOQW is transient.
6.5.2 A generalized recurrence criteria for HOQWs on $\mathbb{Z}^{2}$
We have already presented several recurrence criteria for discrete-time and continuous-time Open Quantum Walks on the set of vertices $\mathbb{Z}^{2}$, constrained by a minimal enclosure. These criteria are interesting tools for understanding the long-term behavior of open quantum walks. Building on these consequences, we aim to derive a stronger recurrence criterion for discrete-time HOQWs, allowing for a more comprehensive analysis of their statistics.
At this point we will allow the auxiliary map of a coin $(D)$ to have more than one invariant state, thus we consider the decomposition of $\mathfrak{h}$ given on Section 4. Therefore, given a minimal enclosure $\mathcal{Y}_{\alpha},$ let $\tau_{\alpha}$ be the unique invariant state with support on $\mathcal{Y}_{\alpha}.$ Introducing the parameter
$\tilde{m}_{\alpha}=\begin{bmatrix}m_{1,\alpha}\\ m_{2,\alpha}\end{bmatrix},\quad m_{1,\alpha}=\mathrm{Tr}\left(D_{1}\tau_{\alpha}D_{1}^{*}\right)-\mathrm{Tr}\left(D_{3}\tau_{\alpha}D_{3}^{*}\right),\quad m_{2,\alpha}=\mathrm{Tr}\left(D_{2}\tau_{\alpha}D_{2}^{*}\right)-\mathrm{Tr}\left(D_{4}\tau_{\alpha}D_{4}^{*}\right),$
we have by [10, Lema 4.2] that $\tilde{m}_{\alpha}$ is independent of the particular minimal enclosure $\mathcal{Y}_{\alpha}$ in $\mathfrak{X}_{\alpha}.$
Therefore, we can combine Theorem 6.15 and results concerning CTOQWs above to state the next property, which is a continuous-time version on the grid of Proposition 4.1.
Proposition 6.16.
Consider a coin $(D).$ We have a decomposition
$\mathfrak{h}=\left(\bigoplus_{\alpha\in A}\mathcal{Y}_{\alpha}\right)\oplus\mathcal{X},$ (18)
where $A$ is a finite set, $\mathcal{Y}_{\alpha}$ is a subspace of $\mathfrak{h}$ being the support of exactly one invariant state $\tau_{\alpha}$ for $\Phi,$ $\mathcal{X}$ contains no enclosure, and
  • $\tilde{m}_{\alpha}=0\Rightarrow$ the walk is $\tau_{\alpha}$-recurrent for all $\tau\in\mathcal{S}(\mathcal{Y}_{\alpha})$;
  • $\tilde{m}_{\alpha}\neq 0\Rightarrow$ the walk is $\tau_{\alpha}$-transient for all $\tau\in\mathcal{S}(\mathcal{Y}_{\alpha}).$
Proof.
Fixed a subspace $\mathcal{Y}_{\alpha}$ which is the support of exactly one invariant state $\tau_{\alpha}\in\mathcal{D},$ we assume that $\tilde{m}_{\alpha}=0.$ Then $(D)$ is $\rho$-recurrent for every $\rho$ by Theorem 6.11. If $\tilde{m}_{\alpha}\neq 0,$ then $(D)$ is transient by Theorem 6.12.
$\square$
With the decomposition on Equation (18) established, we can say that $\mathcal{Y}_{\alpha}$ is either recurrent or transient. It remains to understand what happens when a quantum state has support outside of $\oplus_{\alpha}\mathcal{Y}_{\alpha}.$
Analogously to the one dimensional case, our recurrence criterion is:
Theorem 6.17 (A Generalized Criterion for coins $(D)$).
Consider a coin $(D)$ of dimension $d<\infty.$ Consider the decomposition of $\mathfrak{h}$ as given in Equation (18), and $\tau_{1},\ldots,\tau_{w}$ the respectively invariant states of $\Phi$ in $\mathcal{Y}_{\alpha},\;\alpha=1,\ldots,w.$
$(1)\quad\tilde{m}_{\alpha}\neq 0\;\forall j\in\{1,\ldots w\}\Rightarrow(D)\mbox{ is transient;}$
[…]
$\begin{cases}\rho\text{-transient for }\rho\in\mathcal{D}_{1}:=\mathcal{D}(\oplus_{k\in J^{C}}\mathcal{Y}_{k})\\ \rho\text{-recurrent for }\rho\in\mathcal{D}_{1}^{C}.\end{cases}$
[…]
$\begin{cases}\rho\text{-transient for }\rho\in\mathcal{D}_{2}:=\mathcal{D}(\oplus_{k\in J^{C}}\mathcal{Y}_{k}\oplus\mathcal{X})\\ \rho\text{-recurrent for }\mathcal{D}_{2}^{C}\end{cases}$
$(3)\quad\tilde{m}_{\alpha}=0\;$ for every $\alpha\Rightarrow$ the walk is recurrent.
7 Examples
Example 7.1.
We start with a CTOQW induced by a coin $(A,H),$
$A_{1}=\begin{bmatrix}3&-1\\ 0&0\end{bmatrix},\quad A_{2}=\begin{bmatrix}1&-2\\ 2i&0\end{bmatrix},\quad A_{3}=\begin{bmatrix}1&1\\ -2&2\end{bmatrix},\quad A_{4}=\begin{bmatrix}-2i&i\\ 0&2\end{bmatrix},\quad H=\begin{bmatrix}-1&h\\ \overline{h}&2\end{bmatrix},\;h\in\mathbb{C}.$
In this case,
$m=\frac{2\;\text{Re}(h)-10\;\text{Im}(h)+19}{\alpha}\;\begin{bmatrix}4\\ 1\end{bmatrix}$,
where $\alpha$ depends on $h,$ but it is never null. Therefore, by Theorem 6.13, $(A,H)$ is recurrent if and only if
$\text{Re}(h)=5\;\text{Im}(h)-\frac{19}{2}.$
$\Diamond$
Example 7.2.
Consider an HOQW induced by a coin $(D),$ where
$D_{1}=\frac{1}{2\sqrt{30}}\begin{bmatrix}2&0&-2\\ 0&4&1\\ 0&0&\sqrt{5}\end{bmatrix},\;D_{2}=\frac{1}{2\sqrt{6}}\begin{bmatrix}2&0&-2\\ 0&4&1\\ 0&0&\sqrt{5}\end{bmatrix},\;D_{3}=\frac{1}{2\sqrt{30}}\begin{bmatrix}4&0&1\\ 0&2&-2\\ 0&0&\sqrt{5}\end{bmatrix},\;D_{4}=\frac{1}{2\sqrt{6}}\begin{bmatrix}4&0&1\\ 0&2&-2\\ 0&0&\sqrt{5}\end{bmatrix}$.
The invariant states of the auxiliary map are of the form $\rho_{\infty}=\omega_{1}\sigma_{1}+\omega_{2}\sigma_{2},$ where
$\sigma_{1}=\begin{bmatrix}1&0&0\\ 0&0&0\\ 0&0&0\end{bmatrix},\quad\sigma_{2}=\begin{bmatrix}0&0&0\\ 0&1&0\\ 0&0&0\end{bmatrix},\quad\omega_{i},\omega_{j}\geq 0,\;\;\omega_{1}+\omega_{2}=1.$
We have
$\tilde{m}_{1}=\frac{1}{10}\begin{bmatrix}-1\\ -5\end{bmatrix},\qquad\tilde{m}_{2}=\frac{1}{10}\begin{bmatrix}1\\ 5\end{bmatrix}$.
By item (1) of Theorem 6.17, the walk is transient.
Example 7.3.
Consider the coins $(A,H_{1})$ and $(A,H_{2}),$ where $A_{k}=D_{k}$ for $k=1,2,3,4$ as in Example 7.2, and
$H_{1}=\begin{bmatrix}0&0&0\\ 0&0&0\\ 0&0&0\end{bmatrix},\qquad H_{2}=\begin{bmatrix}0&1&0\\ 1&0&0\\ 0&0&0\end{bmatrix}$.
Its straightforward that $(A,H_{1})$ is transient, since we cain join the results of Example 7.2 and Theorem 6.14. However, it is very interesting to see that $(A,H_{2})$ is recurrent. Indeed,
$\rho_{\infty}=\frac{1}{2}\begin{bmatrix}1&0&0\\ 0&1&0\\ 0&0&0\end{bmatrix}$
is the only stationary state and $m=0,$ thus the coin is recurrent by Theorem 6.13.
$\Diamond$
[…]
Acknowledgements This work was supported by Universidade Federal do Rio Grande do Sul (UFRGS).
[…]
Conflict of interest The author declare that there is no conflict of interest.
[…]
Data Availability The datasets generated during the current study are available from the corresponding author on reasonable request.